A Complete Bibliography of Publications in Mathematics of Computation, 1980–1989

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

02 January 2024 Version 2.30

Title word cross-reference

#10999 [Bla82]. #1291 [Pex81, Tem81, vK83a]. #27887a [AD88]. #3506 [Fet81c]. #361 [Sal84, SL85]. #4049 [Fet82a]. #4774 [Sal81]. #484 [Sha82b]. #4914 [Lon83, Smi83]. #5375 [Kni88]. #5542 [GS80]. #5546 [Bri82]. #5952 [Car80, Fet81e, Fet82a]. #6725 [Fet81b]. #7119 [Che83]. #7933 [Kel82b]. #8017 [Bac80]. #8606 [vK83b]. #A256 [Dav82].

(1 + 2x) exp(x^{2}) erfc x [SL81]. (a, b) ← (\frac{a+3b}{4}, \frac{\sqrt{ab}+b}{2}) [BB89]. 0 [LMv86]. 0, 1 [Kal82]. 0 ≤ x < ∞ [SL81]. 1 [GL82, LMv86, LC88, Set80]. 1 + 2^{a} = 3^{b}5^{c} + 2^{d}3^{e}5^{f} [Ale85]. 10^{1}0 [te 86b, te 86a]. 11 [ACHv80]. 19 [SF85, SS86b]. 1 < n < 7 [NB89]. 2 [CL83, GL85, LS89]. 2 000 000 [SF85]. 21 [MPR81]. 25 [NN80]. 25 · 10^{9} [PSW80]. 26 [NN80]. 284 457 [SS86b]. 2^{m} ± and1 [Bri82]. 2^{n−2} ≡ 1(mod n) [Rot84]. 2^{n−k} ≡ 1(mod n) [She86].

2 ∗ 3 ∗ 5 ∗ ⋯ ∗ p + 1 [Tem80]. 2 · 3 · 5 ⋯ p ± 1 [BCP82, Kel83a]. 3 [Bea86a, BGZ85, DF81, Hag83, Kis83]. 32 [Rus89]. 3 · 2^{n} + 1 [Kel82b]. 42 [AD88]. 50 · 10^{9} [KSW86]. 9 [SS86b]. _{3}F_{2}(1) [Lav87]. A [CR82, Sha81]. A(α) [Sha82a]. A5^{n} − 1 [Wil87]. A7^{n} − 1 [Wil87]. A^{4} + B^{4} + C^{4} = D^{4} [Elk88]. A^{4} + B^{4} = C^{4} + D^{4} [Zaj83]. α [Coh84]. Ax = λBx [McC81a]. B [DM83, LA82, LA84]. BN [Pia81]. coth [Lin86]. D < 10^{7} [LQ88b]. D^{1/2} [EL89]. Δu = u − u^{3} [Lun80]. e [Bai88b]. e^{−x} [Und86]. F_{17} [Gos80]. F_{7} [Bac80]. Γ [BB82]. GL(n, Z) [PP80a, PP80b, PP80c]. h [Mos81]. H^{−1} [Nak84]. I_{0}(ξ) [Las82b]. I_{e} [Oku83]. ∫_{0}^{1} t^{−1} log^{n−1} t log^{p}(1 − t)dt [Köl82]. ∫_{0}^{∞} e^{−μt}t^{ν−1} log^{m} tdt [Köl83b]. ∫_{0}^{p} u^{n}e^{−u^{2}}(u + x)^{−1}du [NV82]. ∫_{0}^{π/2} J_{0}^{2}(γ sin x)dx [SF87]. ∫_{0}^{π/2} J_{ν}^{2}(λ cos θ)dθ [Won88]. ∫_{0}^{π/2} log^{n} cos x log^{p} sin xdx [Köl83a]. J_{0}(x) [WBR82]. J_{1}(x) [WBR82]. J_{n}(z) [CM83]. k! + 1 [Tem80]. K [Lue89, Jae83, Jae85]. k · 2^{m} + 1 [CW80, Kel82c]. k · 2^{n} + 1 [BCW81, BCW82, Kel82a, Kel83b, Jae83, Jae85]. Kf = g [Lue89]. Ki_{n}(x) [Bla82]. L [Rot82]. L^{2} [Jia87]. L^{∞} [Jia87, Gol80, BD80]. L_{1} [SH80]. L_{2} [CT87b]. L_{2}[0, 1] [Cho82]. L_{∞} [SW82a, Sch80a, SH80]. L_{p} [CT87b, Las86]. λ [EM87]. log(e^{x}e^{y}) [NT87]. LU [Cha85, Elm86, JT81, Cha84]. M [Col82, AJS83, HP84]. M_{12} [BR88]. M_{22} [Mal88]. ℚ [Mal88]. Q(\sqrt{2 + \sqrt{2}}) [CD86]. Q(\sqrt{3 + \sqrt{2}}) [CD86]. Q(\sqrt[3]{D}) [Wil81b, Wil82b]. Q(\sqrt{D}) [Wil82b]. GL_{6}(2) [Dar86]. SL(3, Z) [GGT87]. sn(x, k) [Fra81b, Wri81a, Wri81b]. n! ± 1 [BCP82, Kel83a]. N [Col82, Mur82, Lag83, NB89]. n − 1 [Lag83]. N = 13 [Tro85]. n = 8 [PP80b]. O(n^{1/10.89}) [AL81]. Aut(M_{22}) [Mal88]. lcm [Mor89, Mor90]. mod p [Hud84, Sch85c].

PSL_{3}(𝔽_{4}) · 2_{2} [Mal88]. H^{∘}^{1} [Wah84, SW82a]. P [Ana87, BS88a, CL83, Dea85, Jam80, SS89, Wag80]. p + 1 [Wil82a]. p^{6} [Jam80]. p_{r}(n) [Cos82]. π [Bai88a, Bai88b, BB86, NS84]. π(x) [LMO85]. π(x) − li(x) [te 87]. ψ(x) − θ(x) [Per85, Per87]. ψ(x; 3, l) [McC84a]. ψ(z) [McC81b]. Q [Rot82, ST87]. Q(\sqrt{3}) [CD88]. Q(\sqrt{dp}) [CL83]. Q(\sqrt{m}) [Vau85a]. R1031 [WD86]. R^{2} [Mae86]. R^{3} [Gir88]. S [DM82a]. Σ(k) [Bra83]. SL_{2}(Z) [FMPT84]. \sqrt{D} [Wil81a]. ∑ X_{i} = ∏ X_{i} [Bro84b]. T [Sid80a, Mar80a]. tanh [Lin86]. θ(x; 3, l) [McC84a]. u^{2} + 5v^{2} = m [Wil80a]. u_{t} = νu_{xx} + u(1 − u) [Puc89]. W_{p}^{1} [CT87b]. x^{3} + y^{3} + z^{3} = 3 [Cas85]. x^{n} + y^{n} + z^{n} = 3 [SB84]. xy + k [Bro85]. y^{..} = Ay + B(t)y^{.}c(t) [Gek82]. y^{2} − k = x^{3} [Ste86a]. Y^{2} = X(X^{2} + p) [BC84]. y^{2} = x^{3} − 7x + 10 [BT83]. Y_{0}(x) [WBR82]. Y_{1}(x) [WBR82].

-abundant [Coh84]. -adic [Dea85]. -boundedness [Jia87]. -clusters [NB89]. -connected [DF81]. -convergence [BD80]. -D [Bea86a, LC88]. -dimensional [AJS83]. -Distribution [Mar80a]. -divisibility [Wag80]. -Factorizations [Cha84, Cha85]. -function [Oku83]. -functions [BB82]. -Galerkin [Nak84]. -invariant [EM87]. -Invariants [CL83]. -norm [EL89]. -Projection [CT87b, SW82a]. -projections [Jia87]. -range [Mos81]. -regular [HP84]. -splines [DM83, LA82, LA84]. -stability [CR82]. -Stable [Sha81, Ana87, DM82a, Sha82a]. -superlinear [ST87]. -transformation [Sid80a]. -version [BS88a, SS89].

01021a [Ano88b].

1 [KP88]. 10 [Fet84b]. 15 [Fet81d, KP88, vK83f]. 16 [vK83c, vK83d, vK83e, vK83g]. 16th [Sal81]. 1933 [Sal80]. 1983 [GW84]. 19th [Sal81].

2 [DNW88, Mah80]. 2-D [DNW88, Mah80]. 23 [Dav82]. 26 [Sal84, SL85]. 27 [Fet82a]. 29 [Bac80, Bri82, GS80, Lon83, Smi83].

3 [LQ88a]. 3-D [Bea86b]. 3-Sylow [LQ88a]. 30 [AD88, Kel82b]. 32 [Bla82, Sha82b]. 33 [Car80, Fet81e, Fet82a, Kel82a, Kro80, Ske86a]. 34 [Fra81a, vK83b]. 35 [Kel82c]. 36 [Fet82b, Had82]. 37 [BCW82, Mei85]. 38 [Kel83a, Pex81, Tem81, vK83a]. 39 [LA84, te 86c].

40 [Fai84, Fet81b, Jae85, Vau84]. 41

[Has86, Rab85, Rab88b, te 86d]. 42 [Cha85]. 43 [Fet81c]. 44 [Gra87a, Hea86, Per87]. 45 [Lew87, Wil88]. 47 [HH88b]. 48 [Zag88].

50 [Kni88]. 50th [Gau94]. 51 [Bac80]. 52 [Bri82, GS80, Mor90]. 53 [Kel82b, Stu81a]. 54 [Fet82a]. 56 [Che83]. 57 [Bla82]. 58 [AD88, Sal81, Sha82b].

6 [Phi87]. 65009 [Fet82b]. 69-0172 [Fet81b].

80f [Kro80]. 80h [Kel82a]. 80j [Ske86a]. 81f [Fra81a]. 81g [Kru83, Sol86, vK82]. 81i [Kel82c]. 81k [II89]. 82a [Fet82b]. 82c [Had82]. 82g [Phi87]. 83b [Mei85]. 83c [Kel83a]. 83i [LA84]. 83m [te 86c]. 84c [Fai84]. 84e [Vau84]. 84i [Rab85, Rab88b]. 84k [Jae85]. 84m [Has86]. 85e [te 86d]. 85i [Cha85]. 86b [Hea86]. 86k [Per87]. 86m [Lew87, Wil88]. 87a [Gra87a]. 87g [HH88b]. 87k [Zag88]. 88m [Ano88b].

A. [II89]. A.-P. [II89]. Abel [Egg89, Lub83, Lub85]. Abelian [Set80, van82a]. absolutely [DW89]. Absorbing [Hal82, Hig86a, Hig87, TH86]. Abstract [Kea87, Las86]. Abundance [BH86]. abundant [Coh84]. Academic [Ano88b, Car80, Che83, Fet81e, Fet82a, Kru83, Sol86, vK82]. Accelerated [Hug86, AH81, Had82, HH88b, Mar80c]. accelerating [AJS83, Lev80]. Acceleration [Bre82b, Kel86, Saa84, GL82]. accelerators [SF82]. Accuracy [JS85, Mon85a, BM82b, Eri85c, GL85, Sch80b]. Accurate [Gar84, HH88a, MNP84, BDK82, BSC81, DK85, HTE81, OS86b, San88, Sch83a]. adapted [Eri85c]. Adaptive [VB81c, EJ88, MT86, OO84]. Addendum [Fra81a, Zag88]. addition [McC86]. Additive [CS80a, CS83a, LO83]. ADI [HV89]. adic [Dea85]. Admitting [För87]. advanced [Gre82]. Advection [Pas80, Hal86]. Aerospace [Fet81b]. ahead [PTL85]. Air [Fet81b]. Aitken [MP85]. Algebraic [Boy85a, Boy85b, BE88b, Smy84, AD88, BE88a, Bru82, BW88, BW89, Eas86, GN88, Ger83a, KLL88, LP86, Rhe84, Sid80b, Smy81, Toi86, Vau83, Vau84, Vau85b]. Algorithm [AL81, Bai88a, Ger88, GM84, HJ81, HW84, PSS82, Sto80, Yan87, BF89a, BC82b, CZ81, Cas80, DL88b, EK89, Fie81, FM85, For81, Gar88a, Ger83c, Gla87, Gre82, Hil82, JF87, Kiw85, Las82a, Leo80, Mar80b, McC83, Mit83a, PTL85, SL84, Sey87, Sim84, Spe86, Tan88, WW87, Wun83, Wun85]. Algorithmic [Wil80a, Mon81]. Algorithms [BP88a, BB86, BP87, MNSS89, Mos81, SS85, van88, BPX88, BM80, Bul80, Kim88, McC86, New85, Stu81b]. aliquot [GS80]. almost [Kis81]. alternating [Phi80]. alternative [McC83]. Always [Bro85]. amalgamated [Sto81]. Amicable [Bor81, BH86, te 84, te 86a, te 86b, BB88]. analogue [WD84]. Analysis [AK81, BOP80, BR82, Ben84a, Ben84b, BR85, Dem88, Gar88b, II89, JP82, KL85, Pai80, PS83, Sid80a, Ste84b, Tha81, BSC81, BHK84, BPX88, Cyb83, Elm86, EN87, FP85, FO82, Gal83, IIMPL81, JP86, OS86a, Sal81, Wei84b, Wun83, Ziv82, Stu81a]. Analytical [Mas83]. Anderson [Kin83]. anharmonic [IIMPL81]. Anisotropic [DMG89b]. Anniversary [Gau94]. antiautomorphisms [WWL81]. any [Lag83]. AOR [Mar81, Pap83]. Apollonian [Boy82a]. Application [CD87, GM85b, Ioa85a, PdW86, de 86, Ell83, GM85a, Ger83a, Gir87, HTN83, Has86, Rei85, SWW83, Spe86, dW85a]. Applications [II89, PP80b, Ril83, BJVZ89, Cyb83, Mah80]. Applied [Tro80, Hac80, LeV86, Sar87]. Approach [Ste84b, VB81c, CMSY83, SN83]. Approximants [NZ82, Und86]. Approximate [Lev83, Sch86, JS81, Kim88, NV88, Vér82]. approximating [Mon88, Tsu86]. Approximation [BF86, CQ82, CH84, Cho82, DLR81, DW83, Egg84, JE82, Kea83, KG85b, MORR81, MW86c, Sha84, SL81, WF82, ADT81, AG83, BO89, BDK82, Bre89b, Dea85, DRS85, DW81, DS80b, Ell83, EGL80, FM86, FCM88, Gar84, Gev84, GR81, KG85a, LV89, Lue89, Mil87, Mit83b, Neu87, Pie88, Rei85, SB89, SS86a, Ste86b, SN83, VB81b, dS80a, dS81a]. approximation-theoretic [VB81b]. Approximations [Bal84a, Ben83, BE88b, DD86a, DD86b, Hea85, Ise82, IN83, Mic87, Mik89, MS87, New84, BD80, BDK83, Bal84b, Bla82, BP88b, BP88c, BE88a, CM80, CL84b, CTW89, Dur87, DNW88, EO80, EO81, FG88, GT81, GT85, GT87, HLR88, Hea86, Hig86a, HS85, JF87, JLTW87, Las86, Mic83, Nas87, OS83, OHW85, OT88, Pen81, RS82, SJI85, Tad83, Tan88, WBR82]. arbitrarily [For88]. Arbitrary [Met86, AG83]. arc [BS88b]. argument [Lew85b, Lew87]. arise [GT84]. Arising [Que81, NS84]. Arithmetic [FP84a, Gor87, GW85a, Man84, Pri83, Wil80b, GW85b, Gra87a, Hud85, McC84b, Pri85, Rot82, Stu81b]. Arithmetic-Harmonic [FP84a]. ARL [Fet81b]. Artificial [Hal86]. Aspects [SW86a]. assisted [BBG86, Vrs86]. Associated [Bau86a, Lac84, Bau86b, DiD82, Lew86, MPR81, SW82b, Vrs86]. Asymptotic [AW83, BLL87, BL88, Che82, GV87, HK87, IL88, SW82b, Sta81, Ste84c, SF87, Won88, Boy81, Eli86, FW85, HC81, Mei85, SW85, Und86, WM81]. Asymptotically [Dix81]. asymptotics [Tem86]. attenuation [Loc81]. attractors [HLR88]. August [Gau94]. author [Bla82, Bri82, Had82]. Automatic [Del81, HTN83, Has86]. automorphism [SS88]. Avdelas [Had82]. average [MZ88]. Averages [Mar82b]. Averaging [Hud85]. away [Lay81]. Axiomatization [Man84].

Babuska [Pit80]. Backtrack [BC82b]. Backward [BE88a, BE88b, BPST89, Eld82a, Eld82b, LP86]. Badii [Kni88]. Baillie [Kel82a]. Band [DMS84]. Banded [BP88a]. Base [Fet81b, Leh81, Leo80]. Based [Sha81, LO83, Nie83, Sha82a, SSO83]. Bases [GS89, LS87, Ste88]. basic [Sch80b]. basin [DLR81]. basis [Por85, VB81a]. Beckers [Shi84]. Behavior [Ste84c, Gal83, Gau83, Tad84a]. Bell [How80]. Below [te 86a, te 86b, Zag82]. bending [Pie88]. Berlin [Fet84b]. Bernoulli [HP84, TW87, Wag80]. Bertrand [Ioa85a]. Berwick [Tza86]. Bessel [Ben83, Cha82a, CM83, CS80b, Gla81, Las82b, LW82, New84, Pie84, Tem86, TW80, WBR82]. Best [Cho82, Sha84, Dea85, Net80]. best-possible [Net80]. between [BF89a, Leh88b, Odl87]. Bi [IN87, KBF87]. Bi-cyclide [KBF87]. Bi-Orthogonality [IN87]. Bickley [Bla82]. bifurcation [BB85a, Mit83a]. biharmonic [Mon88]. Binary [PdW86, de 86, JS87a]. Binomial [Goe88, Hud84]. Birch [BGZ85]. Birkhoff [Bör89, Dyn84]. Bisection [Adl83, Kea87, Sty80]. bits [KLL88]. Bivariate [CH89, DL88a, AEM87]. Björck [Kin83]. Block [BP88a, Cas83a, Cas83b, Pap83, RW84, Mak80, Mak81]. block-by-block [Mak80, Mak81]. Blurred [Swa89b, Swa89a]. bodies [Dja81]. Bois [Dav82]. Book [Tha81, Ano88b]. Borders [Swa89b, Swa89a]. Borweins [Bai88a]. Boston [Ano88b]. Bound [BN82a, BC89a, BC89b, EKS88, Kea87, Par83, Cha82a, Let86, TW89, Zag82]. boundaries [Tre85]. Boundary [AV85a, AV85b, Ben84a, Ben84b, Gar87a, Gir88, GR89b, MW86b, McI86a, McI86b, Mic87, Slo83, Arn83, BS88a, BSC^{+}80, BL88, CT87a, CS87b, DL81, Dut89, ET84, EN87, Fai84, FG88, Gar87b, Gar88c, GT81, GT85, GT87, Gui87, HK87, HH88a, Hal82, Hal86, Hig86a, Hig87, HV89, II89, JN80, Kat80, KFD83, Las86, LeV86, Lun86, MW86a, MR83a, Mat87, McL86, Mic83, Mun81, NV88, OS86a, OO84, Sch86, Shu87a, SS89, Tad83, TH86, Wei84a, Wei86, Wei84b, Win81]. bounded [LS86, TL85]. boundedness [Jia87]. Bounds [Olv88, BBG82, CGT88, HS85, JS85, MNR89, Mie80, Ral83, TW83]. Box [DL88a, Wei84b]. Brasil [II89]. Breakpoints [Cho82]. Breeding [BH86]. Brent [te 86c, New85]. Brillhart [Bac80]. British [Gau94]. Buhler [Kel83a]. Bur [Smi83]. Burgers [Rob89]. Byrd [Fet81c].

C [Bla82, Fai84, Fet81b, Kel82c, Sha82b].

Calculating [MT80, FP85, Wil80c]. Calculation [FW84, HHR^{+}87, Lev83, WCS80, Wri81a, BK89, Dja81, Mas83, SWW83]. Calculations [CGS89, EO80, GT84, Luc83b]. Calculator [Tha81]. Calculus [Sal80]. Calderón [II89]. Cambridge [Stu81a]. capturing [Sze89]. Cardinal [CJW87, Lun80]. Carlo [Bue84, SL84]. Carmichael [Dub89]. Case [Cas83a, Cas83b, DMG89a, DMG89b, PdW86, PP80a, PP80c, de 86, AH81, EL89, Had82, TW89]. Cases [Car89]. Caslin [Fet81b]. Cauchy [Rab85, Rab88b, BS88b, CM87b, FM86, Ioa83, Ioa85a, Ioa85b, Mas89, Oke87, Rab83, Rab88a, RL89]. Cauchy-type [Ioa83, Ioa85a]. cell [Cot87]. Cells [BB85b]. century [Gau94, Sal81]. Certain [Bre80, Hud84, JA82, Wah84, Akr85, BLL87, BBG87, Bru82, BW87, Cha82b, CR82, CR83, Eas86, Gar88a, Gre82, Hil82, Las82b, Let86, LS83, Mas89, TW86, Wag80, Wil80c, WD84]. Chains [Löh89, McC86]. change [FM85, Ste84a]. character [MP87]. characterization [ST87]. Characterizing [Ada87]. Characters [Dar86]. Chebyshev [Bla82, FO86b, Has86, Per87, Boy81, Boy82b, BS84, CM83, DW81, FR89, Fis88, FO86a, För87, GR81, Got81a, GO82, HTN83, Hor80, JF87, Per85, Saa84, Sch80b, Scr88, SL81, Tan88, WBR82, Xu89]. Chebyshev-Type [FO86b, FO86a]. Chemical [BB85b]. Choice [Gfr87b, Gfr87a]. Circle [Mae86]. Circular [CS87a, CsAS87, GO82]. Class [Bue87a, CM87a, Gla81, How80, HW82, IS86, Mol87, Sch83b, SW88a, Ser80, Was87, van82a, BC86, BW88, BW89, BPv89, Bue84, BS84, CMSY83, CW87, CGT88, CR82, DW85b, EN87, HHR^{+}87, Hil82, HS85, Let86, LQ88a, Mas83, MPR81, MP87, Mun81, SWW83, Set80, Sta81, SW88b, TW86, WDS83, Wil88]. class-number [TW86]. Classes [Len84, Mac80, DM82a]. Classgroup [CL83]. classical [MZM89]. Cliffs [vK86]. Closed [Köl82, DLR81, LS85]. Closure [BC82a]. clusters [NB89]. Codes [Gor87, Sha81, SW86b, Sha82a]. coefficient [AD82, Fal83, GOT81b, Ric81]. Coefficients [Bal84a, Fra81b, Goe88, Hud84, NT87, Wri81a, Wri81b, Bal84b, BDS85, Bro84a, Cha87, EGL80, FW80, Fra81a, Hor80, Kar86, Kee89, Kip84]. Cohen [Was86]. Cohomology [Rus89]. Collocation [AW83, AW84, Bru89, LN89, MR83a, dS80a, dS81a, AD82, Bru85, CG88, CS87b, CES88, Fis88, KS87, Kum88, Nak84, SW85, Sch86]. collocation- [Nak84]. collocation-Galerkin [AD82]. collocation-type [KS87, Kum88]. Columbia [Gau94]. column [JS87a]. Combined [LL83b]. Comment [II89]. comments [TW80]. Common [Ben83, KR89, Lac84]. Commutator [BC82a]. Comp [AD88, Bac80, BCW82, Bla82, Bri82, Cha85, Fai84, Fet82b, Fet82a, Fra81a, Gra87a, GS80, Had82, Has86, Hea86, HH88b, Jae85, Kel82a, Kel82b, Kel82c, Kel83a, Kro80, LA84, Lew87, Mei85, Per87, Rab85, Rab88b, Sha82b, Ske86a, Vau84, Wil88, Zag88, te 86c, te 86d]. Compact [Swa80, MP87]. Comparison [BK89, Scr88]. comparisons [SF82]. Complete [DM83, Man84, PP80c, Fet81b, SS88]. Complex [ALS89, GO82, Sch83b, Ser80, BS88b, Buc87, BW87, CH84, CM83, Cor80, CT89, DW85b, Ell83, Fet81b, GR81, Kim88, Rei85, SN83, Tan88, Wil88]. complex-valued [GR81]. Complexity [Wer84b, Wer86a, FP85, Wer80, Wer86b]. Composite [Dyn84, Ise82, YB88, EP86, Jae83, Jae85, KP89, MT86]. Compressible [DR83]. Comput [Mor90]. Computation [AMS84, Bai88a, BP89a, BP89b, BBG82, Cha84, DW85b, Ell83, ELS89, Kar80, Lan81, MP87, PZ82, PWZ82, Sch85c, SW88b, TL81, TZ87, Wil88, te 86a, te 86b, BM80, Bru82, Buc87, BW89, BF89b, BPv89, Cha85, CW87, Gau94, GR89a, Las82b, Mat87, MNP84, Sid80b, Tem86, TW80]. Computational [Wer80, Fai84, Gau94, KFD83, SW88c, Tha81]. Computations [van82a, EG87]. Computer [CD86, Man84, BBG86, Gar88a, Ril83, Vrs86, WWL81]. computer-assisted [BBG86]. Computing [AS86, BC82a, BC82b, CGM86a, CGM86b, Can87, Dar86, GS89, GW85a, GW85b, Gra87a, KR89, LMO85, LS85, New86, Poh87, Sil88, EM87, JF87, Kro80, Lev82a, Mar80b, McC81b, Mos81, Vau85b]. Concerning [Ste88, Wil81b, BCW81, BCW82, SW88c]. Condition [Hei89, Gui87, OHW85, Sch85b, Tad84b, Vil88]. Conditions [Gir88, Pit80, BS88a, BC86, FG88, HK87, HH88a, Hal82, Hal86, Hig86a, Hig87, LeV86, Slo83, Swa80, Tad83, TH86]. conductor [ACHv80]. configurations [DM83]. conformal [Rei85]. Congruence [Rot84, She86]. Congruences [TW87, AEM87]. congruential [AW89, ELT88, Nie89]. Conjecture [Yam86b, BGZ85, Dea85, Yam86a]. Conjectures [Fra81b]. Conjugate [SS85]. connected [DF81, WT82]. Connection [Leh88b]. conservation [CS89, CM80, EO81, GL85, HTE81, JS87b, Luc86a, Luc86b, OS82, OS83, OT88, San88, Shu87a, Shu87b, Spe87, Sze89, Tad87, Vil88]. conservative [Sch85b, Tad84b, Win80a]. consistent [DM82a]. Constant [Bue87b, GW85a, Bai88b, BM80, Cha87, GW85b, Gra87a]. constants [Bai88b, BF89a, BW87, Dea85, Net80, SW82b]. constrained [DF83, Neu87, ST87, Tap88]. constraint [BS88a]. Constraints [MW86c, LP86, Mil87, SB89]. Constructing [PP85a, PP85b]. Construction [Kre86, Ske86b, Ske86c, ATZ83, BPS86a, BPS87, BPS88, BPS89, Fer80, Gla87, Vau85a]. contain [SS86b]. containing [Las82b, Mas83, SS88, Tem86, Tre85]. contains [Mah85]. Continued [Dud87, PW85, Wil81b, Cha80, Gau83, Jag82, McC83, TW83, Wil81a, WD84, WW87, Wun85]. Continuous [AM89, Ben84a, BJVZ89, JF87, Zen86]. Continuum [II89]. Control [Win80b, CT87a]. Convection [DMG89a, DMG89b]. Convection-Diffusion [DMG89a, DMG89b]. Convenient [GT85, GT87]. Convergence [AW83, BDK83, BM82a, Boy80b, Bra84, BP87, Bre82b, CN88, Cot87, Del89, EN87, Gfr87b, Gre86, Hac80, HT81, Hug86, KG85b, LC88, Lor86, LP86, LS86, Luk88, Mak80, Mic87, Pie88, Puc89, Rab88a, Rob89, Sze89, Yse86, van88, AJS83, BD80, BM81, Boy81, Boy82b, CS87b, CES88, CM87b, DRS85, DW81, Dut89, Eri85b, Far86, Gau83, Gev84, Gev89, Gfr87a, HH88b, HV89, Ioa83, Ioa85b, Jag82, JS87b, Kin80, KG85a, Lev80, LR84, Mar80b, Mar81, Mas89, OT88, Par80, Pen81, Phi80, Rab86b, San83, SW85, Sch84c, Shi84, Shi87, Sid80a, SF82, SS89, ST87, Sty80, Tsu87, Vér82, Ypm83, dS81b]. Convergent [Bai88a, Bea86b, Gar88b, KMS^{+}86, Arn83, Bea86a, McC81b, PT85]. Converging [BB86]. Convex [Cos86, Met86, Kiw85, Spe86]. convexity [DEH89, FM86]. convexity-preserving [DEH89]. convolution [Mah80]. coordinate [KBF87]. coordinates [WT82]. Cormack [Kel82c]. Corner [Wah84, Dja81]. corner-shaped [Dja81]. corners [Pit81]. corrected [vK82]. Correcting [Gor87]. Correction [Pai82, Rab85, KBF87]. corrections [Auz87]. corrector [JM85]. Corrigenda [Ano80, BCW82, Bla82, BP88b, Fai84, Fet82a, God84a, Had82, Hea86, HH88b, LA84, Mor90, Rab88b, Vau84, Wil88, te 86c, te 86d]. Corrigendum [Ano88a, Cha85, Fet81a, Fet82b, Gra87a, GS80, Has86, W.80, Jae85, Kel83a, Kro80, Lew87, Mil87, Per87, Sha82b, Shi19, Ske86a, Fra81a]. Cosine [BDS85, BD89a, BD89b]. cosines [Sal84, SL85]. Cost [Kea87, New86]. counterexample [Dea85]. Counting [EKS88, JS87a, Shi86, Shi19]. coupled [IIMPL81]. coupling [JN80]. course [Stu81a]. covers [LS85]. Crandall [Kel83a]. crimes [Gol80]. Criteria [Sch83c, Bri82, GT81, GT85, GT87]. criterion [Iga84]. Critical [FCM88, BvtW82, te 86c, te 86d, vt83, vtW86]. Crosswind [JSW87]. Cryptosystems [Kob87]. Cubatures [Mae86]. Cubic [Car89, ET85a, ET85b, GS89, JS81, LO82, Was87, WCS80, Yan87, BC86, CS87c, DEH89, DW85b, Let86, LQ88b, SWW83, TW86, Wil80c, WDS83, WD84, Wil88]. Cumulative [AMS84]. Curiosity [Ste88]. Curvature [MS87]. Curve [Can87, Kob87, Yan87, van88, BGZ85, Las82a, Mon87, SH80]. curve-fitting [SH80]. Curves [BM82c, Kre86, Rüc87, Sch85c, Sil88, Zag87, ACHv80, CES88, Rei86, Sar87, TZ87, Zag88]. Cusp [FMPT84]. Cyclic [ET85a, EMT85, Gra87b, RW84, Tro80, Tro85, HHR^{+}87, Leh88b, SW86b, Tro89, Vau85a]. cyclide [KBF87]. Cyclotomic [BS89, BSSW87, IS86, LL87, SW88a, BW87, CT89, SWW83]. Cyclotomy [LL83a]. cylindrical [CS80b].

D [Bri82, Fet81c, te 86c, Bea86a, Bea86b, DNW88, GL85, LC88, Mah80]. D.C [Lon83, Smi83]. Data [Fra82, BPST89, CT87c, CTW89, Eri85a, GV87, JLTW87, Las86, Nie83]. Davis [Che83]. Decay [DMS84]. Decimal [Bai88a, Sal84, SL85]. Decomposition [Bun82, Mah85]. Decompositions [JT81, Pai80, Ste84c, MP87]. Defect [Auz87]. Definite [Gla83, Man80]. definiteness [Rab86a]. Definition [GfZ88, BF88]. Deformations [BB85a]. degenerate [Noc89, Soc86]. Degree [Bau86a, Met86, AG83, Bau86b, GO82, Rab80, SW85, Sal84, SL85]. delay [Bru89]. denominator [Mah85]. Density [Cos82]. Dependent [BOP80, Bal84a, HT81, AD82, Bal84b, BDS85, DW83, Kar86, Kee89, Tad83]. derivative [CS80b, CS82, CS83b, Eli86, KG82]. derivatives [Apo85, Ioa85b, McC81b, Net80]. derived [McC86, Toi81]. Description [PP80c]. Design [Gar89]. Determination [Bra83, EL89, FT84, Wil82b, AW89, BW87, CT89, DLR81, Net80, Rei86, Set80]. determined [Mol87]. determining [Fie81, Gar88a, Zul88]. devices [Zlá86]. diagonal [Kat80]. diagonally [Mar80c, Mar81]. diagram [BB85a]. Diaz [Fai84]. Difference [Bue87b, DW83, Eva83, Gar87a, Gar88b, GfZ88, Mic87, Sch85a, Soc86, Sto80, te 87, BSC81, BPST89, Cas80, CM80, EO81, For88, Gar87b, Gar88c, GT81, GT85, GT87, GT84, GOT81b, Gre82, Hac80, HTE81, Hig86a, HS85, Hof85, Kar80, Lew85a, Luc86b, LR80, MZM89, McC83, Mic83, Mun81, Nas87, OS82, OT88, San83, San88, Slo83, Sta81, SJI85, Swa80, Tad83, Tad84b, Tre85, Wei84a, Wei86, Whi83, Ypm83]. Difference/Finite [Gar88b]. Differences [Isa80, Mor89, Sal80, Kro80, MNP84, Mor90, NS85]. differencing [San83]. Different [te 81, Sch85a]. Differential [AV85b, BE88b, Cas83a, Cas83b, DD86a, DD86b, Gek82, Maj86b, Rhe84, AV85a, BW85, BE88a, Bru84, Bru89, Cha82b, Cha87, CS80a, CS83a, CR82, CR83, DHT81, Fat80, FO82, Gre82, HLR88, Hor80, Ise81, Kat80, LO83, LeV86, Lew85a, LP86, LR80, Maj86a, Mak80, Mak81, MZM89, RHD81, Sch83a, Swa80, TZ89, Wer80, dS80a, dS81a, de 87, vt85a, vt85b]. Differential-algebraic [Rhe84]. differential-difference [Gre82, Lew85a, MZM89]. Differential/Algebraic [BE88b, BE88a]. Differentiation [BE88b, AJ81, BE88a, LP86]. Differing [LM81]. Difficult [Dem88]. Diffusion [DMG89a, DMG89b, Cha87, Gal83, Hal86, JS86, JSW87, LC88, OS86b, Ric81, SW83, Sze89]. digit [Sha82b]. Digits [Bai88a, LMv86]. Dimension [CH89, ET84, EN87, Gar88a]. Dimensional [MMM88, PP80a, PP80c, PH84, VB81c, AJS83, BS88b, BK89, CT87a, CN88, Cot87, Dja81, Hal82, Hof85, Lev82a, Mas89, Neu87, OS86b, VB81a, VB81b]. Dimensions [BM82a, PP80b, PP80c, BM82b, Gui87, SW83, Sze89, WZ88]. Diophantine [Ale85, Bro84b, Cas85, Dea85, Gup84, Hil82, Hud84, Mol87, SB84, Wil80a, Zaj83]. Direct [DM82b]. Dirichlet [Auz87, Bra81, BP89a, BP89b, HW82, Wer82]. discontinuity [Lay81]. Discontinuous [DHT81, DD86a, DD86b, CS89, JP86, Ric88]. discrepancy [AW89]. Discrete [AB87b, Bal84a, Ben84b, Tsu87, AB87a, BDK82, Bal84b, CTW89, DK85, FM86, GV87, JLTW87, Sch80a]. Discretization [GR89b, BB85a, BL88, CT87c, DW81, Eld82a, Eld82b, Gev89, Hal82, SSP85]. Discretizations [Maj86b, Kee89, Maj86a, MZ88, Spe87]. discriminant [BF89b, God84a, God84b, LQ88b, Sey87, Vau83, Vau84, Vau85b]. Displacement [DR83]. Dissection [Ana80]. Dissipation [KG85b, KG85a]. dissipative [Lay81]. distance [Ziv82]. Distinct [Lun88, Kis85]. Distribution [Mar80a, Pom81, Ray88a, FW84, Hud85, Odl87, Ray88b, Wri84]. distributions [SW86b]. divergence [LV89]. divergence-free [LV89]. Divided [NS85, Kro80, MNP84]. divisibility [Wag80]. Divisible [Kis83, BB88, Kis85]. Division [Mon85b]. Divisor [Ray88a, Ray88b]. Divisors [Goe88, KR89, Len84, Wag83]. Does [Wer84b]. Domain [GGT87, LC88]. domains [Auz87, Gol81, Gol82, HK87, Pit81, RHD81, Sch80a]. dominant [Mar80c, Mar81]. dominating [Gaá88].

Double [Lev82b, Lev83, IIMPL81, Lev80, Tem86]. Doubled [Löh89]. doubly [WT82]. doubly-connected [WT82]. Dover [Dav82, vK83b]. drives [GS80]. due [BB85a]. Dueck [Wil88]. Dyer [BGZ85]. dynamic [HTE81]. Dynamics [Le 81a, HH88a].

edition [Car80, Fet81c, Fet81e, Fet84b, Pex81, Stu81a, Tem81, vK82]. Edwards [Bla82]. Effect [McC86]. Effective [Eas86, MNR89, PZ82, PWZ82, Wil87]. effects [Hud85]. efficiency [McC86]. Efficient [BP88a, BS80, ELS89, Smi89, Wil80a, Kin80, Mit83a, NV88, Sid80b, Spe86]. eigenfunctions [Gar84]. Eigenvalue [Mar82a, MORR81, Nas87, Pai82, PSS82, Saa84, ER80]. Eigenvalues [BP88a, GK89, BO89, Gar84, dS80a, dS81a]. eigenvector [BBG82]. eigenvectors [BO89]. Eight [Mor89, PP80c, Hag80, Mor90]. Eighteen [Pri83]. Eighth [BP81]. Einstein [GM85a, GM85b]. elasticity [PS83]. Element [AK81, Ben84a, Ben84b, CT87b, Gar88b, Gir88, SW82a, Ste84b, Wah84, Yse86, ADT81, AW82, BS88a, BO89, BD81, BR82, CN86, CMSY83, CT87a, CS89, DW89, DL81, Dur87, DNW88, Eri85a, Eri85b, Eri85c, EJ88, Fai84, Fie81, Fri84, Gol80, Gol81, Gol82, Han82, JN80, JP82, JS87b, JSW87, KFD83, KL85, Lus80, MZ88, MNR89, Mil85b, Mon88, OS86b, Pit81, PS83, PS85, RS82, Sch80a, SW83, SS89, Sze89, TZ89, Wer82, Win80a, Win81, Zlá86]. element-Galerkin [BO89]. element-penalty [CMSY83]. Elementary [Mey84, Smi89]. Elements [BBP81, AM89, BR85, Bre89a, CF89, Gau83, Shi84, Shi87]. eleven [Hag83, Kis85]. Elimination [Ske80]. ellipsoidal [ATZ83]. Elliptic [ACHv80, AKS85, BM82c, Car87a, Car87b, Car88a, Car88b, Car89, Fra81b, Gor89, Kob87, Kre86, MM89, Rüc87, Sch85c, Sil88, Wer86a, Wri81a, Wri81b, Zag87, de 86, BW85, BPS86a, BPS86b, BPS87, BPX88, BPS88, BP88b, BP88c, BPS89, BGZ85, CT87a, DR85, EJ88, Fet81c, Fet81b, HK87, Kip84, Las82a, MZ88, MT86, Mil85b, Mon87, Mun81, Nak84, Rei86, RHD81, Sar87, Sta81, TZ87, Wer82, Wer86b, Zag88]. Enclosure [GR89b]. endpoint [Boy81, LS86, Sid80b]. energy [Gev84]. engineers [Fet81c]. Englewood [vK86]. English [Car80, Dav82, Fet81e]. enlarged [vK82]. entire [GL82]. Entropy [OHW85, EO80, Sch85b, Tad84b, Tad87, Vil88]. enumeration [SS88]. equality [DF83]. Equally [LL82]. Equation [Bra84, BC84, CD88, McI86a, PdW86, Poh87, SS84, Ste86a, Wil80a, de 86, Ale85, AD82, ADT81, AW82, AM89, BDK83, BS88b, Bro84b, Bro84a, Cas85, Coh82, CES88, DK85, FM86, Hal82, Hal86, Hig86a, Hig87, Hof85, JP86, Las82a, Lun80, McL86, Mon88, OHW85, Puc89, Rob89, SB84, SB89, Swa80, Win80a, Zaj83, dS80a, dS81a]. Equations [AB87b, AV85b, Bal84a, BD89b, Ben84a, Ben84b, Cas83a, Cas83b, Coh88, DMG89a, DMG89b, DD86a, DD86b, Dup82, Gek82, Gir88, Gup84, HT81, Kel86, KG85b, Le 81a, Le 81b, Maj86b, McI86b, Mis84, Pas80, Str89, Tho81, Yse86, AB87a, AV85a, BD80, BDK82, Bal84b, BDS85, BD89a, BD81, BR82, BW85, BM81, BJVZ89, BT87, Bör89, BHN82, Bru84, Bru85, Bru89, Cas80, CG88, Cha82b, Cha87, CN88, CS80a, CS83a, CS87b, Cot87, CL84b, CR82, CR83, CT87c, CTW89, Cuy82, DHT81, DL88b, Dic80, DR85, Egg84, Egg89, Eld82a, Eld82b, EN85, Fat80, FT84, FO82, FG88, Gaá88, Gev84, Gev89, Gla82, Gla87, Gol82, Gra82, Gre82, GV87, Hac80, HLR88, HS81, Hil82, Hor80, Ioa83, JS81, JS86, JLTW87, JM85, Kat80, Kee89, KL85, KG85a, KS87]. equations [Kum88, Lay81, LO83, LeV86, LC88, LP86, Lub83, Lub85, LR80, Maj86a, Mak80, Mak81, MZM89, MT86, Mik89, Mol87, Nak84, Neu87, PS83, Por85, Rhe84, RHD81, Sar87, Sch83a, Sch81, Slo81, SS86a, Soc86, TZ89, Toi86, TH86, Ven86, Vér82, Wer80, Wol83, Wri85, Zlá86, de 87, vt85a, vt85b]. Equivalence [BT87, Mon81, SSP85]. Equivalent [Ske86a]. Erdélyi [Fet81d, KP88, vK83c, vK83d, vK83e, vK83f, vK83g]. Errata [Sal81, Ano88b, Bac80, Bri82, Car80, Che83, Dav82, Fet81d, Fet81c, Fet81e, Fet81b, Fet82b, Fet84b, Kel82a, Kel82b, Kel82c, KP88, Kni88, Kru83, Lon83, Pex81, Phi87, Sal84, SL85, Smi83, Stu81a, Tem81, vK82, vK83a, vK83c, vK83d, vK83e, vK83b, vK83f, vK83g]. Erratum [Fet82a, vK86, Sol86]. Error [CTW89, Fal83, Gor87, HS85, JR82, JLTW87, Olv88, Pai80, TZ89, VB81c, Akr85, BW85, BN82b, BL88, Bre89b, Che82, CMSY83, Com89, DNW88, Egg89, FO82, Gol80, LM80, McC84b, Mie80, NV88, Por85, Ral83, RS82, Ric88, Sar87, TW83, VL80, Wer82, Ziv82]. Error-Correcting [Gor87]. errors [JSW87, Kiw85]. escalation [Mah81]. Essentially [CGS89]. Estimate [Mae86, BN82b, Ric88]. Estimates [BP87, GK89, HT81, Lay81, McC84a, Pai82, Per85, Per87, Wah84, BHK84, Che82, CMSY83, Com89, CTW89, DR85, Dur87, DNW88, Egg89, Fal83, Gol80, JR82, JLTW87, McC84b, NV88, RS82, Sar87, Tho80, TZ89]. Estimating [PSS82, Bre89b]. Estimation [Por85, VB81c, GT84]. estimators [BW85]. Euclidean [CD86, PH84]. Euler [Phi87, Bai88b, BM80, CN88, Cot87, JS86, Rot82]. Euler-summation [Phi87]. evaluating [CM87b, WDS83]. Evaluation [AB80, BP88a, LA82, LA84, Lev82b, Lin82, LW82, Lin86, NV82, SF87, FW85, Gra88, Mah81, Mah85, Mas89, Mei85, Sid82a, Smi89, Stu81b, WM81]. evaluations [AJ81, Lav86]. Even [PP80b, SW85]. every [Hag80]. Evolution [Dup82, Mik89, Vér82]. exact [AW89, FCM88, Rab80]. example [Gra83]. Existence [Cha84, CL83, Mah80, Cha85]. Expansion [Fra81b, Pie84, Won88, Wri81a, Wri81b, Eli86, SW82b, Wil81a]. expansions [BLL87, BL88, CM83, IL88, LM80, Mun81, Sch80b, Sta81]. expectation [Bue84]. experiments [Gar89]. Explicit [CD88, McC84a, McC84b, Rei86, Ste86b, Coh82, Luc83b]. Exponential [Isa80, Ise82, IN83, Mil85a, de 86, EGL80, Mas83, MNP84]. exponentially [Arn83]. exponentiation [McC86]. Expressions [Köl82, KBF87, Stu81b]. extended [Lew85a]. extension [CR83, Vau83, Vau84]. Extensions [Fet85, BJVZ89, Ker83, Vau85a, Zen86]. Exterior [Ben84a, Ben84b, Gui87, HH88a]. extrapolation [Cha82b, Kin80, Noc89, Sid82a, Sid88].

F

[Fet81d, Fet81c, Fet84b, KP88, Kni88, Pex81, Tem81, vK83a, vK83c, vK83d, vK83e, vK83f, vK83g, Shi87]. Faber [CsAS87, CS87a, Ell83]. FAC [MT86]. Factor [Gos80, Mah85]. Factorial [Mur82]. Factoring [AD88, BS89, Ger83b, Ser80, Wil82a, Bac80, Bue84, CZ81, SL84, WW87, Wun85].

Factorization [BP81, Dix81, GA81, Man80, vK85, AO83, Bac80, KLL88, Mon87, Sey87]. Factorizations [BMS88a, BMS88b, Cha84, DM82b, KR89, Nau83, Bri82, Cha85, Elm86, SS88]. Factors [GM82, Kel83b, Wil82b, te 81, Bre82a, FM85, Hag80, Hag83, Kis85, Loc81]. false [EP86]. Family [GR88, Loc89, Han82]. Fast [Dix81, Ger88, KSW86, EK89, For81, MT86, New85, PSS89, Tan88]. faster [Sty80]. Fejér [FR89]. Fermat [BP81, Bre82a, GM82, Kel83b, Leh81, Sch83c, SW81, TW89, YB88]. Fermi [GM85a, GM85b]. FFTs [Swa86]. Fibonacci [Ano88b, BMS88a, BMS88b, MP85]. Field [Wil80b, WCS80, Bru82, BW88, BW89, DW85b, Gir87, Pas84, Vau83, Vau84, Vau85b, WDS83, Wil88]. Fields [BP89b, Bue87a, CM87a, CL83, ET85a, EMT85, ET85b, GS89, Gra87b, GA81, IS86, KR89, LS87, LO82, Rüc87, Sch83b, Sch85c, SW88a, Was87, van82a, AD88, BW87, BP89a, BF89b, BPv89, CZ81, CW87, CS87c, Ger83a, God84a, God84b, HHR^{+}87, Let86, LQ88a, LQ88b, Mol87, Rei86, SWW83, Set80, SW88b, TW86, Wil80c, Zlá83, vK85]. Fifteen [BSSW87]. Filament [Gre86]. filtering [GV87]. Filters [ELS89]. find [Iga84]. Finding [Lar80, Wri85, Kal82, Las82a, Leo80, NS85, PT85]. Finite [AK81, BO89, BBP81, Ben84a, Ben84b, CN86, CT87b, DW83, Eri85a, Fri84, Gar88b, Gir88, GA81, LS87, Neu87, PP80a, PP80b, PP80c, Que81, Rüc87, Sal80, SW82a, Sch85c, Ste84b, SJI85, Wah84, Wil80b, Yse86, Zlá86, ADT81, AW82, AM89, BS88a, BD81, BR82, BR85, Bre89a, CZ81, CMSY83, CS89, CF89, DW89, Dur87, DNW88, Eri85b, Eri85c, EJ88, Fai84, Fie81, For88, GR89a, Ger83a, GT84, Gol80, Gol81, Gol82, GOT81b, Han82, HTE81, HS85, Hof85, JN80, JP82, JS87b, JSW87, Kar80, KFD83, KL85, LS85, Lus80, Mah80, MZ88, Mil85b, Mon88, Mun81, Nas87, OS86b, Pit81, PS83, PS85, RS82, San83, Sch80a, SW83, Shi87, SS86a, Sto81, Sze89, TZ89, Tre85, Wer82, Whi83, Win80a, Win81, vK85]. finite-difference [GOT81b, HS85, Hof85, Kar80, Tre85]. Finite-dimensional [Neu87]. Finite-Element [Yse86, OS86b]. finite-section [SS86a]. finite-sheeted [LS85]. Finitely [BB85b]. First [Bre80, Pie84, YP89, AD82, AW84, Cha82a, Egg84, Eli86, Fet81b, Gla82, Gla87, Lus80, TW89, Win81]. first-order [Lus80, Win81]. Fitting [IN83, Lar80, Yan87, SH80]. Five [Kat80]. Five-diagonal [Kat80]. Fixed [Alf82, LM81]. flat [KBF87]. flat-ring [KBF87]. flow [EO80, Fis88, Mos89, OHW85]. flows [CN86, Dja81, Ros83]. following [Spo84]. Force [Fet81b]. Forecasting [LL82]. Form [BCP82, CW80, HJ81, Gaá88, Kel82a, Kel82c, Kel83a, Kel83b, McC83, Tsu86]. Forms [FMPT84, CD87, McD89, Sey87, Ske86a, Wil87]. formula [BN82b, Far86, HC81, Ioa85a, Phi87]. Formulae [CGM86a, CGM86b, GR88, HW82, Min80a, Akr85, Boj81, GN88, Oke87]. Formulas [Apo85, Fet84b, FO86b, FO86a, Lav87, Mur82, Pex81, Sha81, Sha86, Ske86b, Ske86c, Tem81, vK83a, Che82, Dyn84, For88, Lon83, LP86, Mas89, Mon86, Olm86, Sch84c, Sha82a, Ske86a, Smi83, vK83b]. formulation [CN86, CR83, Fri84, KG82, OO84, Pas84]. Four [Bra83, MMM88]. Four-Dimensional [MMM88]. Four-State [Bra83]. Fourier [LA84, BT87, CGS89, For81, Kip84, LA82, Lun83]. Fourth [Eva83, Car80, Fet81e, HTE81, Kat80, Slo83, Stu81a]. fractals [Gar88a]. Fraction [Wil81b, Cha80, Jag82, Mah81, Mah85, McC83, Wil81a, WD84, WW87, Wun85]. Fractional [GfZ88, Isa80, Lub85, Mon81]. Fractions [Dud87, PW85, Gau83, TW83]. framework [CS89]. Fredholm [BT87, HS81]. Free [Bea86b, Bea86a, LV89, MNSS89, NV88, Sto81]. Frequency [IN83]. Fresnel [Hea86, Hea85]. Fricke [Coh88]. Friedman [Fet81c]. Friedrichs [Tad84a]. friendly [Sid88]. Frobenius [Toi81, Yam86a, Yam86b]. full [OHW85]. Fully [Bal84a, BDK82, Bal84b, DK85]. Function [Alf82, Bau86a, Boy80b, Bra83, BI86, CT87b, EKS88, Fet85, Fra81b, IS86, IN83, Laf84, Mac80, Mil85a, SW81, WH83, Wri81a, Wri81b, Apo85, AJ81, Bau86b, BBG87, BF88, BvtW82, CN86, Cha80, CM83, CS80b, CS82, CS83b, Eli86, FW80, Fra81a, FW84, Ker83, Las82b, Leh88a, LW82, Loc81, Lun83, Lun80, LM80, Mah85, MNP84, Odl87, Oku83, Per85, Per87, Tem86, TW80, VL80, Wri84, te 86c, te 86d, vt83, vtW86]. functional [LM80, Wol83, de 87]. functional-differential [de 87]. Functionals [Mee80]. Functions [Ben83, CW83, För87, GM85b, JA82, KCL82, KE84, Lac84, Lar80, Lin86, Pie84, Que81, Ray88a, Xu89, ATZ83, BLL87, BB82, Bla82, BBG87, Boy81, Boy82b, BS84, Cha82a, CH84, DiD82, DS80b, FM86, Fet81d, Fet84b, Fri84, GL82, GM85a, GN88, GR81, Hab83, IL88, Kip84, LV89, Lev82a, Lew85b, Lew87, Lon83, LS86, Mas83, Mon86, MP87, New84, Pex81, Ray88b, SW85, Sch80b, Smi83, Smi89, Tem81, TL85, VB81a, Vrs86, WBR82, vK83a, vK83c, vK83d, vK83e, vK83b]. Fundamental [GGT87, PZ82, PWZ82, BF89a, Buc87, CS87c, Ril83, Zlá86]. Further [Fet85, Ioa83, Laf84, TW80]. fusion [Gar89].

G [Dav82, Fet81d, Had82, Kel82c, KP88, Stu81a, Wil88, vK83c, vK83d, vK83e, vK83f, vK83g]. Galerkin [AD82, Arn83, AB87a, AB87b, BO89, BD80, BDK82, BDK83, CS89, DHT81, DK85, Egg89, ET84, EN87, Gev84, Gra82, Ioa83, JP86, Kee89, KG85a, KG85b, Las86, Lay81, Lun86, McL86, Nak84, Ric88, Sar87, SS89, Tho81, Tho80]. Galerkin/Runge [Kee89]. Gallas [Car80]. Galois [Mal88]. Gamma [BI86, Laf84, Cha80, FW80, Fra81a, FW84, Ker83, Wri84]. Gap [Wei81]. Gaps [Bre80, YP89]. Gas [Le 81a, HH88a]. Gauss [Rab85, Rab88b, Bra84, CGM86a, CGM86b, GN88, GR88, KE84, Rab80, Rab83, Rab86a, Rab86b]. Gaussian [Gra87a, Akr85, Che82, Dyn84, GM85a, GM85b, GW85a, GW85b, Ioa85b, Leh88b, LR84, Ske80, VL80]. Gautschi [Ker83]. General [Min80a, WF82, Auz87, Cyb83, Fri84, McC83, Olm86, SSP85, Tap88, CS89]. generalization [FHS86]. Generalized [BSC^{+}80, Ger88, HP84, HTN83, Has86, Kea87, Lin82, Loc89, Mil85a, PdW86, DM82a, Eli86, ER80, Ioa85a, Lav86, LC88, MP85, Rab80, SJI85, Van82b]. Generated [LS81, EHGL89, Ger83c, Nie89]. Generating [Mar80a, te 84, FW84, Leo80, Wri84]. Generation [For88, Mon85a, Spo84, PSS89, WT82, WW87]. generator [ELT88]. Generators [Tez88, EHGL89]. genuinely [Tad84a]. geodesics [LS85]. Geometric [Mar82b, Jia87, dD85]. German [Fet84b]. GFSR [Tez88]. give [dS81b].

Given [te 84, BS88b, JE82, JS87a, Leo80, Zag82]. Glance [GLS87]. Global [DR85, Far86]. globally [PT85]. Goldberg [NT87]. Goldstine [Sal81]. Golomb [Kel82b]. Goodness [Mae86]. Governing [CL83]. Government [Lon83]. GPST [LC88]. Graded [Gar88c, Bru85]. Graded-mesh [Gar88c]. Gradient [SS85]. Gradient-like [SS85]. Gradshteın [Fet82a, Kru83, Sol86, vK82]. Gradshteyn [Fet81e, vK82]. Gradsteın [Car80, Fet81e, Fet82b]. graphs [DF81, Lon83, SS88, Smi83, vK83b]. Greatest [KR89]. greedy [Ger83c]. Grid [Bea86b, DW83, Sch84a, Sch84b, Bea86a, Hac80, HS81, MT86, Olm86]. Grid-Free [Bea86b, Bea86a]. Grids [KMS^{+}86, For88, OS83]. Group [Dar86, Lan81, BR88, DW85b, Gra83, Leo80, LQ88a, MNR89, Wil88]. Groups [Bue87a, BC82a, CM87a, Jam80, Mal88, Que81, Rus89, Sch83b, Yam86b, BPv89, BC82b, MP87, SS88, Yam86a]. Gunderson [SW81].

H [Bla82, Bri82, Kel82c, Sal81, Tza86]. H. [Sha82b]. half [Egg89, Gau94, SS86a]. half-century [Gau94]. half-line [Egg89, SS86a]. Hall [vK86]. Halley [Cuy82]. Halley-iteration [Cuy82]. Hamilton [CL84b]. Hammerstein [KS87, Kum88]. Hand [Saa87]. Handbook [Fet81c, Lon83, Smi83, vK83b]. Hankel [CS82, CS83b, FW85]. Hansen [vK86]. Harmonic [Ben84a, Ben84b, FP84a, CH84, FM86]. Haselgrove [SM82]. Hausdorff [Gar88a]. Having [Boy80a, Boy89a, Boy89b, McD89]. Heat [McI86a, McI86b, AM89]. Hecke [CD87, GGT87]. height [TZ87]. Heights [Sil88]. helical [Gar89]. Hellan [Com89]. Helmholtz [Gol82]. Henrici [Tha81]. Hensel [von84]. Hermite [Boy80b, DEH89, Dyn84, SSO83]. Herrmann [Com89]. Hessian [GT84]. Hessians [Sha80]. Heuristics [CM87a, Was86]. High [Eri85b, FW80, Fra81a, Gar87a, GR89b, Har86, Vil88, BM80, DK85, Gar87b, LR80, Sch80b, Shu87b, dS81b]. High-Order [Gar87a, Eri85b, Vil88, DK85, Gar87b, LR80, Shu87b, dS81b]. High-precision [FW80, Fra81a, BM80]. Higher [BM82b, BS80, Fet81d, Sch84a, Sch84b, vK83c, vK83d, vK83e, Apo85, BDK82, CN88, Vrs86]. higher-order [CN88]. highly [Fat80, Sch83a]. Hilbert [Coh82, Ger88, Vér82]. Hill [Fet81d, KP88, vK83c, vK83d, vK83e, vK83f, vK83g]. history [Sal81]. homotopy [Zul88]. Howell [Spo84]. Howland [Lin82, LW82, Lin86]. Howland-Type [Lin86]. HP [Tha81]. HP-25 [Tha81]. Hybrid [DD86a, DD86b, LL83b, MORR81, DL88b]. Hybrid-Combined [LL83b]. Hydrocode [Hic81]. Hyperbolic [Bal84a, BD89b, BLM89, KG85b, Mic87, PdW86, AD82, BD80, Bal84b, BDS85, BD89a, Dut89, FG88, Gev84, GT81, GT85, GT87, HTE81, JS85, JP86, JS87b, KG85a, Lay81, LO83, LeV86, Luc86a, Lus80, OS82, Slo83, Spe87, Tad83, Ven86, Vil88, Wer86b, Win81]. Hyperelliptic [Can87, LCC88, LCC89]. Hypergeometric [JA82, Lav86, Lew85b, Lew87, LCC88, LCC89]. Hyperperfect [Hag81, Min80b, te 81].

ideal [BW87, BW88]. identification [Fal83, Ric81]. Identities [BBG88, BBG87]. II [God84a, te 86c, AW84, BM82b, Ben83, Boy84, Boy89a, Boy89b, BPS87, BvtW82, Bue87a, BC82b, Cas83b, CS89, DM83, Eld82b, Fet81d, God84b, GT81, GT87, Kis83, PWZ82, Ray88b, Smy81, VB81b, Yam86a, Yam86b, de 86, vK83d, vK83g]. III [te 86d, BPS88, PP80a, VB81c, vK83e, vt83]. Ill [Gfr87b, EN85, Gfr87a, Mar80b, Neu87]. Ill-Posed [Gfr87b, EN85, Gfr87a, Mar80b, Neu87]. illustrations [dS81a]. imaginary [CW87, HHR^{+}87, Set80]. Immanant [GM84]. Implement [NOPEJ87]. Implementation [CL87a, CL87b, For81, NV88, Ril83]. Implementing [Wun85]. Implicit [Bru84, DW83, JT81, Sha81, HTE81, Hof85, Sha82a, Swa80]. Implies [Ske80]. imposing [FG88]. improperly [Han82]. Improved [Eri85c, FP85, Hei89, EN85, McC86]. improvements [Luc83a]. Improving [Pen81, Wil80c]. inaccurate [DM82a]. including [FP85]. Incomplete [BPST89, LCC89, Man80, RW84, Elm86, Mah85]. Incompressible [Gir88, JS86]. Increased [Wer84b]. Indefinite [Kea83, Wer86a, BPX88, BP88b, BP88c, Dav82, DiD82]. independence [DM83]. Independent [BP89b, LL82, BP89a, GT81]. Index [AMS84, Sto81]. Inductive [Min80a]. Inequalities [BI86, Dud87, JA82, Laf84, Mey84, Ker83, Mit83a, Net80]. Inequality [Tro80, Tro85, Tro89]. Inexact [Alf82]. Infinite [AB80, BBG88, Mar82a, PSS89, Bru89, Lev80, Mat87, Sch86, Sid82a, Sid82b, Sid88, Slo81, SSO83, SN83]. inflow [Tad83]. inflow-dependent [Tad83]. Influence [Wah84]. infrastructure [BW88]. inhomogeneous [Gaá88]. Initial [AV85b, Cas83a, Cas83b, Mic87, Win80b, Alf80, Ana87, AV85a, CT87c, CTW89, Dut89, GT81, GT85, GT87, HV89, JLTW87, Mic83, Par80, Spi85, Win81, dW85a]. initial-boundary [Dut89, GT81, GT85, GT87, HV89, Mic83, Win81]. initial-value [Par80]. Insensitive [Sha81, Sha82a]. instability [Tad83]. Integer [Boy85a, Boy85b, BM82c, BT83, Bue87b, HJ81, LM81, Lun88, Nau83, Col82, Jag82, WD84]. Integers [Dix81, GW85a, LMv86, Smy84, Ste88, Col82, Ger83c, GW85b, Gra87a, Smy81, Wil87, Wró84]. Integral [AB87b, Fet85, GS89, Kea83, Köl83a, Köl83b, McI86a, McI86b, NV82, PP85b, SS86a, Tho81, Zag87, Arn83, AB87a, BB82, BJVZ89, BT87, BS88b, BHN82, Bru85, CG88, CS87b, CES88, Egg84, Egg89, EN85, Fet81b, FP84b, Gla82, Gla87, Gra82, HS81, Ioa83, JS81, JN80, JM85, KP88, KS87, Lub83, Lub85, Lun83, McL86, Net80, PP85a, Sch81, Slo81, Tem86, Ven86, Zag88, vK83f, vK83g, vt85a, vt85b]. Integrals [Car87a, Car87b, Car88a, Car88b, Car89, Fet80, Fet84a, Gat80, Gla83, Hea85, KCL82, Lev82b, Lev83, Lin82, Lin86, BLL87, Car80, CM87b, DiD82, Fet81a, Fet81c, Fet81e, Fet82b, Fet82a, Hea86, HC81, Ioa85a, Ioa85b, IL88, Kru83, Las82b, LA82, LA84, Lev80, Lev82a, LW82, LCC88, LCC89, Mas83, Mas89, Mei85, Oke87, Oku83, Rab83, Rab85, Rab88b, Rab88a, RL89, Sid80b, Sid82a, Sid82b, Sid88, Sol86, TL81, WM81, vK82, Dav82]. integrand [LM80]. integrands [LR84, Rab86b, TL81, VL80]. Integrating [Cor80, Hab83]. Integration [Cas83a, Cas83b, CG88, CR82, Gla82, JP82, LS86, Mon86, Rab80, Rab83, Rab85, Rab86a, Rab86b, Rab88b, RL89, Sch81, SSO83, SM82, TL85, Che83]. integration-collocation [CG88]. integrators [Fat80]. Integro [Mil85a, Bru84, Bru89, Cha82b, Mak80, Mak81, TZ89, vt85a, vt85b]. integro-differential [Bru84, Bru89, Cha82b, Mak80, Mak81, TZ89, vt85a, vt85b]. Integro-Exponential [Mil85a]. interesting [Sha82b]. Interfaces [Ber85, Tre85]. interior [LS86]. Intermediate [LeV86]. internally [DM82a]. interpolates [Olm86]. Interpolating [HW84, Luc83a, Nie83]. Interpolation [CJW87, CDR88, Cos86, DL88a, Fra82, Loc81, Mar82b, Met86, SW86a, Wer84a, DEH89, EK89, Far86, FR89, HTN83, Has86, SW82b, WZ88, dD85]. interpolatory [CM87b]. intersections [LS85]. interval [Mon86]. Intervals [Mar82a, MS84, LS86, MR83a, Mat87, Ral83, Sch86, Slo81, SSO83]. Invariant [van88, Bru82, EM87, GR89a, Gir87]. Invariants [CL83]. Inverse [MZ88, II89, LC88]. Inverse-average-type [MZ88]. Inverses [DMS84]. inversion [ALS89, LG86]. inversions [Nie89]. invertible [Fie81]. investigation [BBG86, Wil81a]. involutions [WWL81]. Involving [GM85b, Lin86, Bai88b, BB82, Cos82, GM85a, Kat80, LW82, Net80, VL80]. irrationalities [WD84]. irrationality [Eas86]. Irreducibility [AO83, Bri80]. Irreducible [Dar86, PP80a, PP80b, PP80c, BBG82, Mar81, SW86b]. Irregular [Ger83c, KMS^{+}86, Lev82a]. irregularities [Hud85]. isolated [Zul88]. Isomorphisms [Poh87]. isoparametric [Fie81]. Isotropic [DMG89a]. Issues [Min80b]. Iterated [Gfr87b, Gfr87a]. Iteration [Alf82, BB89, Pap83, BBG82, Vrs86, Cuy82]. Iterations [Hug86, BPST89, Hac80, HH88b]. Iterative [GK89, Ske80, BR82, BPS86b, EL89, Iga84, Mar80c, Mar81, Mie80, Mon88]. IV [BPS89, PP80b, vtW86]. Iwasawa [EM87].

J [Bac80, Bla82, Bri82, Che83, Fai84, Fet81b, Fet82a, Kel83a, Mei85, te 86c, te 86d]. Jacobi [BG84, CL84a, CL84b, Fer80, GN88, HJ81, KCL82, KE84, Lew85a, Lew86, Loc89]. Jacobian [Can87, DM82a, Fra81b, GT84, Kip84, Wri81a, Wri81b]. Janeiro [II89]. Jbuch [Stu81a]. Johnson [Bla82, Com89]. Julia [Vrs86]. June [GW84].

Keast [Fai84]. kernel [Mak81]. Kind [Car87a, Car87b, Car88a, Car88b, BJVZ89, BHN82, Egg84, Fet81b, Gla82, Gla87, Gra82, HS81, JM85, Lub83, Lub85, Oke87, Slo81, Spe86]. Kirkman [MPR81]. Kleinian [Gra83]. Knot [Mar82b]. knots [MNSS89, dS81b]. Kolmogorov [Puc89]. Korteweg [AW82, BDK83, DK85, Win80a]. Kronrod [Rab85, Rab88b, CGM86a, CGM86b, GN88, GR88, Rab80, Rab83, Rab86a]. Krylov [Saa81]. Kutta [BJVZ89, BHN82, Bru84, Cas83a, Cas83b, CS89, CS83a, DM82a, Kar86, Kee89, Lub83, Pia81, Sha86, Zen86, dW85a].

L [Ano88b, Bri82, GW84, Kni88]. Lab [Fet81b]. lag [Ana87]. Lagrange [Bra81, Pit80, Pit81, Wer84a]. Lánczos [ER80, PSS82, PTL85, Saa87, Sim84, Van82b]. Laplace [Kni88, ALS89, LG86]. Large [Bor81, Bre80, CW80, Ger83b, Har86, Kre86, Maj86b, Saa81, SW88a, Wei81, Zag87, Zag88, Dub89, ER80, Gal83, Gar89, Kel82c, Kel83b, Maj86a, PSS89, SWW83, Tad84a, Toi86, Wró84]. large-time [Tad84a]. Largest [PSS82, DL81, LS83]. Larkin [NS85]. Last [Sch83c, SW81, TW89]. lattice [FP85, LS89, SL89]. Lattices [PH84, PP85b, PP85a]. law [JS87b, Sze89]. Laws [Isa80, CS89, CM80, EO81, GL85, HTE81, Luc86a, Luc86b, OS82, OS83, OT88, San88, Shu87a, Shu87b, Spe87, Tad87, Vil88]. Lax [Tad84a]. layer [BSC^{+}80]. Leading [Gfr87b, Gfr87a]. Least [AKS85, LS81, MW86c, Mil87, BL88, EGL80, Hag80, Hag83, Mit83b, Ste84a]. least-change [Ste84a]. least-squares [Mit83b]. Lebesgue [SW82b]. Legendre [Bau86a, Bau86b, DiD82, Fet85, Lac84]. Lehmer [Bri82, Shi19, LMO85, Rot82, Shi86, Spo84]. lemma [Mei85]. Length [CT86, EHGL89, FP85, Wil81a]. Lenstra [Was86]. Less [PP80c, GO82, KSW86]. Level [Yse86, AG83]. Levine [Sal84]. Lie [MP87]. lifting [Dja81]. like [Kel86, Leh88a, SS85, Vrs86, Ypm83]. limit [Boy84, TW83]. Limited [Noc80]. line [Egg89, SS86a]. Linear [DW83, Gek82, Kin83, Le 81b, Man80, Mee80, Olv88, Saa81, SS85, de 87, vt85a, vt85b, Ada87, AW89, BL88, Cas80, CT87a, CR82, CR83, DM83, Dic80, EN85, EJ88, Fai84, GV87, Hal86, KFD83, Kin80, Lub85, Mar80c, Mon81, Neu87, NV88, Rab88a, RS82, RHD81, SL84, Ste86b, Tan88, WZ88, Wei84b, Zlá83, AW84]. linearly [Hof85]. Liouville [Gar84, Nas87, Pai82]. Lobatto [Rab86b]. Local [Pit80, Sar87, Wah84, Ypm83, dS81b, CS89, Eri85b]. locally [OS83]. Logarithmic [FM86, CS80b, CS83b, Sid80b]. London [Sal80]. Long [Löh89, PW85, Pri85, MR83a]. look [PTL85]. look-ahead [PTL85]. low [CTW89]. Lower [BN82a, BC89a, BC89b, EKS88, Par83, Wer84b, Let86]. Lowest [DL81]. Lucas [BW80, BMS88a, BMS88b, EKS88]. Luke [GW84]. Lune [te 86c, te 86d].

M [Bac80, Car80, Fet81c, Fet81e, Fet82b, Fet82a, Gra87a, Kel83a, Kru83, Sol86, vK82, Shi87]. Machines [Bra83, Wun85]. Macmillan [Sal80]. magnetic [Gar89, Zlá83]. magnetization [FP84b]. magnetohydrodynamic [JE82]. magnetostatic [Fri84, Pas84]. Magnus [Fet81d, Fet84b, KP88, Pex81, Tem81, vK83a, vK83c, vK83d, vK83e, vK83f, vK83g]. Mahler [Dea85]. Majorizing [Mie80]. Mandelbrot [Vrs86]. Mandelbrot-like [Vrs86]. manifolds [Rhe84]. manipulation [CD87]. Many [BB85b, WZ88]. mapping [Rei85]. mappings [GR89a]. Markoff [Zag82]. Marti [EN85]. Mass [Ano88b]. Mat [II89]. Math [AD88, Bac80, BCW82, Bla82, Bri82, Cha85, Fai84, Fet82b, Fet82a, Fra81a, Gra87a, GS80, Had82, Has86, Hea86, HH88b, Jae85, Kel82a, Kel82b, Kel82c, Kel83a, Kro80, LA84, Lew87, Mei85, Mor90, Per87, Phi87, Rab85, Rab88b, Sha82b, Ske86a, Vau84, Wil88, Zag88, te 86c, te 86d]. mathematical [Fet84b, Lon83, Pex81, Smi83, Tem81, vK83a, vK83b]. Mathematics [AMS84, Gau94, Gau94]. Mathieu [BR88]. Matrices [BP88a, Bun82, DMS84, Ger88, KE84, Noc80, Pap83, Ser80, BG84, Dic80, Fer80, Gal83, Ger83a, GT84, JS87a, Kal82, Mar81, PTL85]. Matrix [BC82a, DM82b, Hig86b, Bul80, BBG82, BC82b, EHGL89, EL89, Mar80c]. Maximal [Boy85a, Boy85b, PP80a, PP80b, PP80c, MNR89]. Maximum [Can88, WH83, DNW88, Sch80a].

Maxwell [Ben84a, Ben84b]. May [GW84]. McClure [Mei85]. McGraw [Fet81d, KP88, vK83c, vK83d, vK83e, vK83f, vK83g]. McGraw-Hill [Fet81d, KP88, vK83c, vK83d, vK83e, vK83f, vK83g]. Mean [BB89, FP84a, Mee80, Mey84, Smy84]. Measure [Boy80a, Boy89a, Boy89b, Smy81, Ziv82]. measures [Eas86]. Media [DR83, Ros83]. medium [Hof85]. Meissel [Shi19, LMO85, Shi86]. Mellin [FMPT84]. memoriam [GW84]. Mersenne [NN80, Wag83]. Mesh [BOP80, Ber85, Dup82, Mar82b, McC81a, Eri85b, Eri85c, Gar88c, Gol81, Jia87, Luc86a]. mesh-refinement [Eri85b]. mesh-refinements [Eri85c]. Meshes [MW86b, Bru85, Loc81, MW86a]. Method [Adl83, AB87b, BBP81, Bea86b, Ben84a, Ben84b, Bra84, BP89b, DMG89a, DMG89b, Gre86, Hig86b, LMO85, LL83b, McC81a, MT80, Pit80, Saa87, Tro80, VB81c, Wil82a, WF82, Alf80, AD82, AW82, Arn83, AB87a, AH81, BS88a, Bac80, BR82, Bea86a, BSC81, BHK84, BT87, BS88b, BG84, Bra81, BPS86b, Bre89a, BP89a, Bue84, CT87a, CN88, CS89, Com89, Cot87, CR83, Dja81, DW89, Dub89, EN85, ER80, Eri85b, Eri85c, EN87, EJ88, EM87, FO82, FG88, GL82, GR81, Gol80, Gol81, Gol82, Had82, Han82, HV89, Ioa83, JP86, JS87b, Kal82, KL85, Kin80, KS87, Kum88, Luc86a, Lun86, Lun80, Lus80, LR80, LG86, Mak80, Mak81, Mar80c, Mar81, MT86, McL86, Mei85, Mon88, Nak84, NS85, Nie83, Noc89, OS86b, OO84, PT85]. method [Pen81, Pit81, PS85, Por85, Puc89, Ric88, Rob89, Sch80a, SW83, Shi86, Shi19, Sid88, SS89, Sty80, SM82, Sze89, VB81a, VB81b, WT82, Wei84a, Wei86, WDS83, Win80a, Win81, WM81, Zul88]. Methods [AW83, AKS85, BOP80, BD89b, BLM89, BM82a, BS80, Bre82b, CGS89, Can88, Cas83a, Cas83b, Che83, Del89, DD86a, DD86b, DW83, DR83, Fra82, Gek82, Gir88, GK89, Got81a, Hei89, IN87, LS81, MORR81, Pas80, Pia81, Saa81, SS84, Sch84a, Sch84b, Ste84b, Tho81, Win80b, Yse86, Ana87, ALS89, AG83, BDS85, BD89a, BM82b, BJVZ89, BT87, BPST89, BK89, Bru84, Bru89, CG88, Cha82b, CGT88, CS80a, CS83a, CS87b, CES88, CR82, DHT81, DK85, DR85, Dut89, Egg84, Egg89, ET84, Eri85a, Fai84, FP85, Gla82, Gla87, GOT81b, Gra82, HK87, HS81, Iga84, Ise81, JN80, JP82, JS86, JSW87, Kar86, Kat80, KFD83, KG82, Lay81, LO83, LeV86, Lub85, MR83a, Mie80, Mil85b, Mon87, Par80, PS83, Ros83, SSP85, SW85]. methods [Sar87, Sch85a, Sch86, Sha80, Slo81, Spi85, ST87, Tho80, TZ89, Tsu86, Wer80, Wer82, Wol83, Ypm83, Zen86, dS81b, de 87, vt85a, vt85b, von84]. metric [Bru82, Sha80]. Microfiche [BSSW87, NT87]. Milne [Sal80]. Milne-Thomson [Sal80]. Mindlin [BF86, Pie88]. Minimal [Gau81, Ana87, Las82a, Tsu86]. minimization [CGT88, Kiw85]. Minimum [PP85b, BC86, Lue89, Mar80b, Nie83, PP85a]. minimum-norm [Mar80b]. Miscible [DR83]. Misstatements [Sal80]. Mixed [BOP80, MORR81, Mil85b, Ste84b, de 86, BP88b, BP88c, DR85, JP82, PS83]. Model [DR83, Gar88b]. models [Tre85]. modern [Stu81a]. modes [DLR81]. modest [Gra83]. Modification [Dup82, LG86]. Modified [Wol83, BG84, Kro80, Las82b, SS86a, Tem86, TW80, Whi83]. Modular [CD88, Coh88, Mon85b, Que81, Coh82, CD87, LS85]. Modulus [Boy85a, Boy85b, CS80b, CS83b, ELT88, Fet81b]. moment [FW84, Wri84]. moments [FW84]. Monotone [Cos86, CM80, GR89b, BBG82, Egg89, Luc86b, San83, Spe87]. monotonicity [DEH89]. monotonicity- [DEH89]. Monte [Bue84, SL84]. Mordell [Ste86a]. Morrison [Bac80]. Moving [LS81, Luc86a]. MR [AD88, Ano88b, Bac80, Bla82, Bri82, Car80, Cha85, Che83, Dav82, Fai84, Fet81d, Fet81c, Fet81e, Fet81b, Fet82b, Fet82a, Fet84b, Fra81a, Gra87a, GS80, Had82, Has86, Hea86, HH88b, II89, Jae85, Kel82a, Kel82b, Kel82c, Kel83a, KP88, Kni88, Kro80, Kru83, LA84, Lew87, Lon83, Mei85, Per87, Pex81, Phi87, Rab85, Rab88b, Sal81, Sal84, SL85, Sha82b, Ske86a, Smi83, Sol86, Tem81, Vau84, Wil88, Zag88, te 86c, te 86d, vK82, vK83a, vK83c, vK83d, vK83e, vK83b, vK83f, vK83g]. Multi [Sch84a, Sch84b, Yse86, Hac80]. Multi-Grid [Sch84a, Sch84b, Hac80]. Multi-Level [Yse86]. multidimensional [Hig86a, JR82, Mic83, Noc89]. Multigrid [Bra84, BP87, MR83b, Spe87, Auz87, AG83, BPX88, Bre89a, PS85]. multilag [Wol83]. multilevel [BR82]. multilinear [WZ88]. Multiple [Gor87, HS81, Sil87, IL88, Jia87, LCC88, Mei85, SL89, Smi89, WM81]. multiple-precision [Smi89]. Multiplication [Ger88, Mon85b, McC86]. Multiplications [New86]. multiplier [Bra81]. Multipliers [Pit80, Pit81]. Multistep [DD86a, DD86b, Gek82, LN89, Maj86b, Ske86b, Ske86c, BPST89, KG82, Lub85, Maj86a, Ske86a, de 87, vt85a, vt85b]. Multivariate [CW83, CJW87, CDR88, AD88, DM83, Gra88, Olm86, vK85].

N [Sal84, Stu81a]. N.J [vK86]. Nagell [PdW86]. Nat [Smi83]. Natural [BJVZ89, Zen86]. Navier [BDK82, BM81, FT84, Gev89, Gir88, JS86]. near [Mon86]. nearest [Jag82, WD84]. Nearly [Löh89, Ste86b]. Negative [Tho80, Gev84, Sey87]. neighborhood [Boy84]. Neighborhoods [Mos86]. Néron [TZ87]. Nested [Ana80]. network [Nie83]. Neumann [CT87a]. Newton [FR89, Hig86b, Luc83b, MT80, Noc80, ST87, Toi81, Wer84a, Ypm83, von84]. Newton-like [Ypm83]. Nine [PP80a]. Ninomiya [Has86]. No [Fet81b, AD88, BCW82, Bla82, Cha85, Fai84, Fet82b, Fra81a, Gra87a, Has86, Hea86, HH88b, Jae85, Kel82a, Kel82b, Kel82c, Kel83a, Kro80, LA84, Lew87, Mei85, Mor90, Per87, Phi87, Rab85, Rab88b, Sha82b, Ske86a, Vau84, Wil88, Zag88, te 86c, te 86d]. nodal [FT84]. nodes [Boj81]. noisy [GV87]. nonaveraging [Wró84]. noncompact [CG88]. Noncomputable [Bra83]. Nonconforming [CF89, Bre89a, Shi87]. nondiagonally [Toi81]. nonexpansive [San88]. Noninteger [Ste88]. Noninterpolatory [RL89]. noniterative [WT82]. Nonlinear [AJS83, BD89b, BE88b, DD86a, DD86b, ELS89, Kel86, Min80b, SS84, SH80, BD89a, BR82, BE88a, Bru89, Cuy82, DL88b, DRS85, ELT88, EO81, HK87, HS85, JS87b, LP86, Mak80, Mik89, OS83, Por85, San88, Sch84c, Sid80b, SF82, Tad84a, Toi86, Vér82, Vil88, Zlá83]. nonlocal [Luc86b]. nonnegative [BBG82]. Nonnegativity [DEH89]. Nonnegativity- [DEH89]. Nonoscillatory [CGS89]. nonrandomness [KLL88]. Nonselfadjoint [HT81]. nonsmooth [BPST89, CT87c, JLTW87, Kiw85]. Nonstandard [Gir88]. Nonstiff [Cas83a, HTE81]. Nonsymmetric [Pap83, Saa84, SS85, BPX88, EN87]. nontrivial [Rei86]. Nonuniform [MW86b, Gol81, MW86a]. nonzero [Col82]. Norm [Par83, Akr85, DNW88, EL89, Gaá88, Lue89, Mar80b, Nie83, Tho80, Toi81]. Normal [BC82a, LS87, DLR81]. Norms [BOP80, Dur87, Gev84, SH80]. Note [Bri80, Bun82, HW84, Mar81, Rüc87, Wri81b, Bro84a, Cas80, Cas85, FW85, SB84, SN83, SM82, TW86, WZ88, Wri84]. Nothing [New86]. Number [BP81, BP89b, CM87a, CT86, Gor89, Hei89, HW82, KR89, Mon85a, SW88a, Tez88, YB88, van82a, AD88, Bru82, BW88, BW89, BP89a, CW87, Col82, CT89, DW85b, DF81, ELT88, EP86, Ger83a, Hag80, Hag83, Let86, McC84b, Rei86, SWW83, Set80, SW88b, TW86, Vau85b, WDS83, Wil88, Zag82]. Number-Theoretic [CT86, CT89]. Numbers [Bor81, BH86, BC89a, BC89b, BSSW87, Ger83b, GM82, Hag81, IS86, Kis83, Kis84, Min80b, TW87, Wag83, Was87, te 81, AW89, BB88, Boy84, Bre82a, Bue84, Coh84, Dub89, Eas86, HP84, HHR^{+}87, KLL88, Kel83b, Kis81, Kis85, KSW86, MP85, McD89, Mol87, Nie89, SW88c, Wag80, Zag82]. Numerical [AB80, Ana80, Bai88b, BF89a, Ben84a, Ben84b, BF86, Cas83a, Cas83b, Cha87, CM87a, Cuy82, Dem88, DR83, Fat80, FO82, Hig87, IIMPL81, II89, Le 81a, Le 81b, LO83, LP86, Lue89, MW86b, MZM89, Mis84, NT87, OO84, OS83, Ric81, Ros83, SS84, SB89, Sid80b, Sid82b, Ske80, SF82, Tad84b, Toi86, ATZ83, AJ81, BHK84, BT87, BS88b, Bör89, Bro84a, BK89, Bru85, Cha82b, Che83, CT87a, CS80a, CR82, DF83, Ell83, ER80, Fal83, Gal83, Gla87, HK87, Ioa83, Luc86a, LG86, MW86a, Mos89, Pie88, Sal81, Sch83a, Shu87a, Sid82a, Spi85, SM82, Tad87, Wei86, Wil81a, Zlá83, dS81a]. Nyström [BT87, BS88b].

O.D.E. [dS81b]. Oberhettinger [Fet81d, Fet84b, KP88, Kni88, Pex81, Tem81, vK83a, vK83c, vK83d, vK83e, vK83f, vK83g]. Obrechkoff [Ana87]. obtained [BC86]. obtaining [Spe86]. Occurrence [Bre80, YP89]. OCI [BSC^{+}80]. Odd [BN82a, BC89a, BC89b, Jam80, Kis83, Kis84, Kis85, BB88, Hag80, Hag83, Kis81].

ODE [Sha81, Sha82a]. ODEs [Cor80]. off [Ziv82]. Office [Lon83]. Ohio [Fet81b]. Ojo [Fet82a]. old [Pri85]. Olver [Cas80]. One [DD86a, DD86b, EO81, KE84, BS88b, Boy81, BW88, CW87, ET84, EN87, FM85, God84a, God84b, Hal82, Hof85, Kin80, Lev82a, Loc81, OS86b, SW83, SS88, Spi85, SW88b, TW86, TH86, Wer80]. one- [Lev82a]. one-dimensional [BS88b, Hal82, Hof85, OS86b]. one-factorizations [SS88]. one-point [Kin80]. One-sided [EO81]. One-Step [DD86a, DD86b, Spi85, Wer80]. one-way [TH86]. onto [CT87b]. open [CES88, WWL81]. Operations [Min80a]. Operator [Par83, CG88, Fis88, FP84b, GV87, JE82, Neu87]. Operators [GGT87, DLR81, IIMPL81]. Optimal [AJ81, Gfr87b, GT84, MS84, MT80, Tez88, TL85, Wer82, BD81, Boj81, Bre89a, Bru84, CN86, Dur87, Eri85a, Gfr87a, Gla82, Gla87, Ral83, RS82, Ric88, Ste86b, VB81a]. optimal-order [Bre89a, Ric88]. Optimality [SW82a]. optimization [Luc83b, ST87, Tap88]. Optimum [AH81, Had82]. orbit [Gra83]. Order [Bal84a, BD89b, BS80, BSSW87, Gar87a, GR89b, Jam80, KG85b, MW86b, Rus89, Sch84a, Sch84b, Swa89b, Ada87, AD82, AW84, BD80, BDK82, Bal84b, BDS85, BD89a, BD81, BM82b, Bre89a, Bru84, Cas80, CN88, CS82, DK85, DR85, DL81, Eri85a, Eri85b, Fai84, GL82, Gar87b, Gev84, Gui87, HTE81, Kar80, Kat80, KFD83, KG85a, Lus80, LR80, MW86a, MZ88, MNR89, MPR81, Mil85b, RHD81, Ric88, San88, Sch85b, Shu87b, Slo83, Spe87, SF85, SS86b, Swa89a, Vil88, Vrs86, Wei84a, Wei86, Wer82, Win81, dS81b]. Orderings [Ana80]. Orders [Poh87, Buc87]. Ordinary [AV85b, Cas83a, Cas83b, DD86a, DD86b, Gfr87b, Maj86b, Alf80, AV85a, CS80a, CS83a, CR82, CR83, DHT81, FO82, Gfr87a, Hor80, Ise81, Maj86a, Sch83a, dS80a, dS81a]. Orlicz [Dur87]. Orthogonal [Gat80, Pai80, CQ82, EK89, Gau81, LS83, WT82]. Orthogonality [IN87]. orthogonalization [Cyb83]. Ortiz [CR83]. oscillating [BLL87, IL88]. oscillations [Lev82a]. oscillator [IIMPL81]. oscillators [IIMPL81]. oscillatory [Fat80, HC81, Oke87, Sch83a, Sid82a, Sid88, TL81]. other [Gol82]. Outline [Hag80]. Overrelaxation [Hug86, AH81, Had82, HH88b, Mar80c].

P [Ano88b, Che83, Fai84, Fet81c, Kel83a, Mei85, Pex81, Tem81, te 86c, vK83a]. P. [AD88, II89]. P1 [Bre89a]. Packing [Boy82a]. Padé [Bre89b, Bul80, McC83, NZ82, Und86]. Pairs [te 84, te 86a, te 86b, BBG86, CS87c]. paper [AD88, II89]. Parabolic [BS80, HT81, Win80b, BPST89, CT87c, CTW89, Eld82a, Eld82b, HS85, JLTW87, Kar86, Kee89, Las86, NV88, Noc89, Soc86, Tho80, TZ89, Wer86b]. parallel [WW87, Wun85]. Parameter [CT89, Gfr87b, Gfr87a]. Parameters [Hab83, Rot82]. parametric [Tsu86]. Parity [BBG87]. Part [Ben84a, Ben84b, Ben83, BS80, Cas83a, Cas83b, DMG89a, DMG89b, Fet81b, PdW86, RHD81, WF82, de 86]. Partial [Mah81, Mah85, BW85, HLR88, LO83, LeV86, Phi80, RHD81, Sim84]. Particle [DMG89a, DMG89b, Puc89]. Partition [Mac80, BBG87]. partitioned [BPS86b]. Partitioning [MT80]. Partitions [Gup84, Col82]. patterns [Sha82b]. Patterson [Fet81b]. Peaceman [HV89]. penalty [CMSY83, KL85]. pencils [Mit83b]. Penk [Kel83a]. Perfect [BC89a, BC89b, Gor87, Kis83, Hag80, Hag83, Kis81]. performance [Wun83]. Pergamon [Sal84, SL85]. period [EHGL89, Wil81a]. Periodic [PW85, Ana87, BG84, Fer80, Hab83, Luc83a, TW83]. Periods [LL83a, PW85, Leh88b]. permanents [Kal82]. Permutation [BC82a, Lan81, BC82b]. permutations [Leo80, Spo84]. Perron [HJ81]. Perturbation [Gar88b, Stu81b, AW84, BSC81]. Perturbed [Gar87a, Gar87b, Gar88c, OS86a, OS86b, OO84, SW83, Wei84b]. Peter [Tha81]. Petit [Dav82]. Petrov [Sar87]. phase [Alf80, Ana87, CS80b, CS83b, JR82]. phase-lag [Ana87]. physicists [Fet81c]. Physics [II89, Fet84b, Pex81, Tem81, vK83a]. Piecewise [Cho82, Yan87, Bro84a, MNSS89, Mit83b, Rab88a, RS82, WZ88, dS81b]. piecewise-polynomial [Mit83b]. Pisot [Boy84]. Pivoting [Ser80]. Pivots [Cha84, Cha85]. places [Sal84, SL85]. planar [DF81]. Plane [MS87, BS88b, CM83, Cor80, Ell83, PS83, Rei85, Sch80a]. plasma [JE82]. plate [Pie88]. Plateau [Tsu87]. Plates [BF86]. Pocket [Tha81]. Poincaré [Ioa85a, Ril83]. Poincaré-Bertrand [Ioa85a]. Point [Alf82, Gar87a, BHK84, Boy84, Fai84, FCM88, Gar87b, Gar88c, Kat80, KFD83, Kin80]. Points [BM82c, BT83, Zag87, AW84, FR89, Kum88, Zag88]. pointwise [JSW87]. Poisson [Bra84, FM86]. pole [Cor80]. poles [Mon86, Ste86b]. Pollard [Mon87]. Pólya [Lan81]. polygonal [Sch80a]. polygons [CS87b]. Polyhedra [AS86, Ril83, Spe86]. polyhedral [DF81]. Polynomial [DD86a, DD86b, DS80b, Gat80, GA81, KLL88, KG82, LL82, LV89, Mos86, NZ82, Sil87, Wer84a, AEM87, Boy82b, Ell83, Kim88, Mah85, Mit83b, Olm86, PT85, Rei85, SW82b, Wri85, Zul88, dS81b]. Polynomials [Ada84, BS89, Boy80a, Boy89a, Boy89b, Cho82, CS87a, CsAS87, How80, Lan81, Loc89, Mah82, Mal88, SW88a, AD88, AO83, Boy81, CZ81, CQ82, CT89, EK89, Gau81, Gir87, GO82, Iga84, Lew85a, Lew86, LS83, MNSS89, SSO83, vK85]. population [RHD81]. Porous [DR83, Hof85, Ros83]. Posed [Gfr87b, EN85, Gfr87a, Han82, Mar80b, Neu87]. posedness [TH86]. Positive [Man80, Pie84, Sha84, BBG82, Col82, EGL80, Gal83]. positivity [dD85]. possible [Net80]. postage [Mos81]. Posteriori [Gfr87b, VB81c, BW85, Gfr87a, Luc83a]. postprocessing [Com89]. Potential [BLM89, OHW85, Pas84]. Power [Eva83, GS89, Sha84, Boy81, ELT88, GL82, Sid80a]. powerful [SW88c]. Powers [Hud84, PdW86, de 86, Kip84, Leh88a]. Practical [Sha86]. preassigned [Ste86b]. precision [BM80, FW80, Fra81a, Rab80, Smi89]. preconditioners [BPS86a, BPS87, BPS88, BPS89]. Preconditioning [AG83, RW84, BP88b, BP88c]. predictor [JM85]. predictor-corrector [JM85]. Prentice [vK86]. Prentice-Hall [vK86]. Prescribed [CT86, PP85b, Fer80, PP85a, SS88]. Presence [Yse86]. preserving [DEH89]. Press [Ano88b, Car80, Che83, Fet81e, Fet82a, Kru83, SL85, Sol86, Stu81a, vK82]. pressure [CN86]. Primality [AS82, AL81, CL84a, CL87a, CL87b, Pom87, Tem80, WD86, Bre82a, Bri82, KSW86, PSS89, Wil87]. Prime [Bre80, Goe88, Jam80, PdW86, Wei81, YP89, de 86, te 81, Hag80, Hag83, KP89, Kis85, McC84b, Wun83]. prime-testing [Wun83]. Primes [Bor81, BCP82, CW80, GLS87, Kel83a, LM81, Löh89, Mor89, NN80, Pri83, HP84, Hud85, Kel82a, Kel82c, Kel83b, Mor90, PSS89, Pri85, Sha82b]. Primitive [Coh84, Lan81, LS87]. Principal [Wil82b, BW87, BW88, CM87b, Ioa85b, Mas89, Oke87, Rab83, Rab85, Rab88b, Rab88a, RL89]. Principle [Can88, Sch80a, dS80a]. Printing [Lon83]. priori [BHK84]. probabilistic [Sey87]. Probability [Dem88, KP89]. probable [KP89]. Problem [AK81, Ben84a, Ben84b, Dem88, Gar87a, Gar88b, KP87, Ste86a, Ste84b, Auz87, BCW81, BCW82, BSC81, BHK84, BR85, Bra81, Bul80, Coh82, CF89, DW89, DNW88, FM86, Fri84, Gar87b, Gui87, II89, JR82, Mos81, OS86a, OS86b, Pas84, Pie88, PS85, SW88c, Tsu87, Wei84a, Whi83, WWL81]. Problems [AV85b, BS80, Cas83a, Cas83b, Gfr87b, GR89b, MW86b, Mar82a, Mic87, Saa84, Wer86a, Win80b, Alf80, Ana87, AW84, AV85a, BO89, BSC^{+}80, BL88, BPS86a, BPS86b, BPS87, BPX88, BPS88, BP88b, BP88c, BPS89, BPST89, CMSY83, CT87a, CGT88, Cos82, DRS85, Dut89, DF83, EG87, ER80, ET84, Eri85a, EN87, EJ88, Fai84, Gar84, Gar88c, Gfr87a, GT81, GT85, GT87, GOT81b, HK87, HH88a, Han82, HTE81, HV89, JS85, Kar86, Kat80, KFD83, Las86, LC88, Lun86, MW86a, MR83a, MZ88, Mar80b, Mat87, Mic83, Mil85b, Mit83a, Mos89, Mun81, Nas87, NV88, Noc89, OO84, Par80, SW83, Sch85a, Sch86, Spi85, Tho80, Wei86, Wei84b, Wer82, Wer86b, Win81, dW85a]. Procedures [Bre89b, Lev82a]. process [BD81]. processes [DM82a]. processing [Cyb83]. producing [Dub89]. Product [BBG88, CG88, Sch81, SSO83, CM87b, LS86, Mas89, Sch84c, Sto81]. Products [AB80, Hud84, PdW86, de 86, Car80, Fet81e, Fet82b, Fet82a, vK86, Kru83, Sol86, vK82]. programming [DF83, Mon81, SN83]. Progression [Pri83]. progressions [Hud85, McC84b, Pri85, Rot82]. Projection [CT87b, SW82a, CS89, Par80, dS81b]. Projections [Wah84, Jia87]. proof [Hag80, Hag83, dD85]. Proofs [Pom87, Bre82a]. propagation [EG87].

Properties

[Kel82b, Lew86, BF88, Egg84, FP84b, IIMPL81, Lew85a, LS89, Pie88, Shi84, Und86, Wer82]. Property [Gla83, AD88]. Prove [CD86]. Pseudo [Pas80]. Pseudoprimes [BW80, EKS88, Gor89, MM89, PSW80, Pom81, Ada87, McD89, Rot82]. Pseudorandom [Tez88, AW89, EHGL89, ELT88, Nie89]. Pseudospectral [Got81a, FG88, GOT81b]. Pseudospectral-Chebyshev [Got81a]. pulse [LC88]. pulse-spectrum [LC88]. Pure [WCS80, TW86, Wil80c, WDS83]. pyramid [Spe86].

QR [Ste84c]. Quadratic

[Bue87a, DL88a, Del89, Ger83b, KR89, Mit83b, Sch83b, Sil87, de 86, AEM87, Bue84, CT87a, CW87, DF83, Fie81, LQ88a, Mol87, Rei86, Sey87, SW88b, TW86, Vau83, Vau84]. Quadratic-Exponential [de 86]. Quadratically [BB86]. Quadrature [CGM86a, CGM86b, FO86b, FO86a, För87, GM85b, GR88, Ise81, LM80, Mon86, Oke87, Slo81, Xu89, Akr85, BN82b, Boj81, BS84, Che82, Dyn84, Egg84, GM85a, GN88, HTN83, Has86, Ioa85a, Ioa85b, LR84, Lun83, Sid80b, Sid82b, SL89, Ven86, VL80]. Quadratures [KE84, Rab88a]. quadrilateral [Shi84]. Quartic [Ada84, Str89, Buc87, BW87, BF89b, BPv89, God84a, God84b, HHR^{+}87, Set80, Vau85a]. Quartically [Bai88a]. Quasi [Dur87, Le 81b, Mic87, Noc80, SW82a, WH83, Luc83b, ST87, Toi81]. Quasi-Initial-Boundary [Mic87]. Quasi-Linear [Le 81b]. Quasi-Newton [Noc80, Luc83b, ST87, Toi81]. Quasi-optimal [Dur87]. Quasi-Optimality [SW82a]. Quasi-Smooth [WH83]. quasilinear [Mil85b]. quasiperfect [Kis81]. quasistationary [Zlá83]. Quaternary [Hud84]. quick [PSS89]. Quintic [SW88a, DEH89]. Quotient [Leh81, MMM88, Sto80, McC83]. Quotient-Difference [Sto80, McC83].

R [vK86, Kel82a, Kel83a, Mei85, Pex81, Tem81, te 86c, vK83a]. Rabinowitz [Che83]. Rachford [HV89]. Radau [Rab86b]. Radii [Boy82a]. Rado [Bra83]. Ramanujan [BBG86, PdW86]. Random [Mar80a, Mon85a, AEM87, KP89, Puc89, Rob89]. range [Mos81, Sid82b]. Rank [Kre86, BW88, BGZ85, FM85, LS89]. rank- [LS89]. Rapid [Shr85, Lev82a, WDS83]. rapidly [BLL87, IL88, McC81b]. Rate [Boy80b, Bra84, DW81, Far86, Boy81, Boy82b, DRS85, Eri85b, Pen81]. rate-of-convergence [Boy82b]. Rates [BM81, DMS84, Gfr87b, LR84, Luk88, Rab86b, Gfr87a, Sch84c]. Rational [Bla82, Hea85, Hea86, IN83, Lar80, SW86a, WBR82, Dic80, EK89, Mah85, Ste86b, Vrs86]. Rationals [HW84]. ratios [TW80]. reaction [Gal83, OS86b, SW83]. reaction-diffusion [Gal83, OS86b, SW83]. Real [EMT85, ET85b, Gra87b, LO82, SW88a, Smy84, van82a, BF89b, BPv89, CS82, CS87c, Gau83, LQ88b, SWW83, Smy81, SW88b, TW86]. Reciprocal [Boy80a, Boy89a, Boy89b, FW84, Wri84]. reciprocals [Wró84]. reconstructing [BG84]. recovery [CN86]. rectangle [AW89]. rectangles [DL81]. rectangular [RHD81]. Recurrence [DiD82, Hor80, Kro80, Lew85b, Lew87, Olv88, PdW86, de 86, Gau81]. recurrences [Ada87]. Recursive [Bul80, CR83]. reduced [CS80b, CS83b, JP82, Por85]. Reduction [RW84, VB81c, VB81a, VB81b]. Reentrant [Wah84]. Refinement [Ber85, McC81a, Ske80, Eri85b]. refinements [Eri85c]. regions [BPS86b, WT82]. Regression [LL82]. regular [Gar84, HP84, Nas87, Olm86]. Regularity [Wer84b, CTW89]. Regularization [Gfr87b, ALS89, FCM88, Gfr87a]. Regularized [Luk88, Bör89, Neu87]. Regulator [WCS80, CMSY83, Wil80c, WDS83]. Reissner [BF86]. Related [BM82c, GW85a, FW80, Fra81a, GW85b, Gra87a, JP82, Loc81, McD89, Wer86b]. relating [Oku83]. Relations [Olv88, BF89a, DiD82, Gau81, Hor80, Kro80, Lew85b, Lew87]. Relative [Ziv82]. relatively [Hag83]. Relaxation [Bra84]. remain [DM82a]. Remark [Mei85, Tza86, AD88]. Remarkable [Gla83]. Remarks [PP80b, Was86, Vér82]. reorthogonalization [Sim84]. Rep [Fet81b]. repeated [Mah85]. Representation [Ste88, SL89].

Representations [Ral83]. Res [Fet81b]. Residue [Eva83, Len84, Mac80]. residues [WW87]. Resolution [Har86, Gaá88]. Respect [Bau86a, Bau86b]. restrictions [Spi85]. result [OS86a]. Resultants [LL87]. resulting [BP88b, BP88c]. Results [Wil81b, van88, Bai88b, BF89a, BBG87, CQ82, Ioa83, LP86, Rab88a, SW88c, VB81b]. Review [Tha81]. Riele [te 86c]. Riemann [te 86c, te 86d, Apo85, BvtW82, Eli86, Leh88a, vt83, vtW86]. Right [Saa87]. Right-Hand [Saa87]. Riley [Gra83]. Ring [GW85a, GW85b, Gra87a, KBF87]. rings [Mah80, von84]. Rio [II89]. Robust [Shr85]. roles [Par80]. Root [Hig86b, Lar80, Mos86, NS85]. Root-Finding [Lar80]. Rootfinder [Shr85]. rootfinding [Kim88]. Roots [Bau86a, Kel86, MT80, Sch85c, Bau86b, PT85]. rotating [DLR81]. Rotkiewicz [KP87]. rough [Las86]. round [Ziv82]. round-off [Ziv82]. row [JS87a]. rule [CM87b, Ven86]. Rules [SSO83, BS84, Ioa85a, Ioa85b, LS86, Lun83, LS89, Rab80, Rab83, Rab85, Rab86a, Rab86b, Rab88b, RL89, Sid80b, Sid82b, SL89]. Runge [CS89, BJVZ89, BHN82, Bru84, Cas83a, Cas83b, CS83a, DM82a, Kar86, Kee89, Lub83, Pia81, Sha86, Zen86, dW85a]. Russian [Car80, Fet81e]. Ryzhik [Fet81e, Fet82a, Kru83, Sol86, vK82]. Ryzik [Fet81e, Fet82b, Car80].

S [AD88, Car80, Fet81e, Fet82b, Fet82a, Fra81a, Kel82b, Kru83, Sol86, vK82]. Sadiku [Fet82a]. Salamin [New85]. Sander [Shi84]. satisfying [EO80, OHW85]. scalar [CM80, GL85, JP86, Luc86b, OT88, Pas84, Sze89, Tad84a]. Scaled [Ste84c]. scales [Sch85a]. Scan [CD86]. Scattered [Fra82, Nie83]. Scattering [AK81, BLM89]. Scheme [Gar88b, GT81, Har86, Kel86, Gev89, HTE81, Hof85, NV88, San88, Slo83, Tad84a, Van82b, Whi83, Zlá83]. Scheme-independent [GT81]. Schemes [Gar87a, HT81, JT81, KG85b, KMS^{+}86, Le 81b, Lor86, BSC^{+}80, Gar87b, Gar88c, Gev84, GL85, Hal82, JM85, Kar80, KG85a, Luc86b, Mos89, Mun81, OS82, San83, Sch85b, Shu87b, Soc86, Sta81, Swa80, Tad84b, Tad87, Vil88, Wei84b, dW85a]. Schröder [Vrs86]. Schrödinger [IIMPL81, SS84].

Sci [Phi87]. Secant [DM82b, Ste84a, Tap88]. Second [Bal84a, BD89b, Bue87b, Car87a, Car87b, GM84, KG85b, MW86b, Sch85b, Swa89b, BD80, Bal84b, BDS85, BD89a, BJVZ89, BHN82, Cas80, DR85, Fai84, Fet81c, Fet84b, Gev84, Gra82, HS81, JM85, Kar80, KFD83, KG85a, KG82, Lub83, Lub85, MW86a, MZ88, Mil85b, RHD81, Slo81, Spe87, Swa89a, Wei84a, Wei86, Wer82]. Second-Order [BD89b, KG85b, MW86b, Swa89b, Sch85b, BD80, BDS85, BD89a, Cas80, Fai84, Gev84, Kar80, KG85a, MW86a, MZ88, Mil85b, Spe87, Swa89a]. section [SS86a]. Sectors [CS87a, GO82, CsAS87]. segmented [OO84]. segmented-adaptive [OO84]. Seidel [Bra84]. selected [Gar84]. Selection [Del81, VB81a]. self [LS85]. self-intersections [LS85]. selfadjoint [BO89, MZ88]. Selfridge [Bri82]. semi [SN83]. semi-infinite [SN83]. semiconductor [Zlá86]. semicontinuity [HLR88]. Semidiscrete [Bal84a, Bal84b, HT81, Par80, TZ89]. semidiscretizations [JS85]. Semigroups [WWL81, HLR88]. semilinear [CT87c, CTW89, JLTW87]. Seminar [II89]. semisimple [MP87]. separably [Alf80]. Sequence [Boy82a, Del81, Mah82, GS80, Kel82b, NS84, Sid80b]. Sequences [Kin83, PdW86, de 86, EHGL89, MP85, Mie80]. serial [Nie89]. Series [Boy80b, BC82a, Fra81b, GM85b, Lav87, NT87, Pie84, Scr88, Sha84, Wri81a, Wri81b, AJS83, Boy81, Car80, Cyb83, Ell83, Fet81e, Fet82b, Fet82a, GL82, GM85a, vK86, Hor80, Kip84, Kru83, Lav86, Lev80, LCC88, LCC89, McC81b, NS84, Phi80, Sid80a, Sol86, vK82]. set [FT84, Leo80, Wró84]. Sets [Bro85, Eva83, Lag83, Lun88, GR89a, Ger83c, PSS89, Toi86, Vrs86, Wag80]. Several [Saa87, Hab83]. Sextic [EMT85, Gra87b]. Shamanskiĭ [Kel86]. Shamanskiĭ-like [Kel86]. shaped [Dja81]. Shapiro [Tro85, Tro89]. Sharp [CMSY83, DNW88, Sch84c]. Sharpening [Swa89b, Swa89a]. Sharpness [Wah84]. sheeted [LS85]. Shepard [Far86]. shift [NS85]. Shock [CGS89, ELS89, Lor86, Sze89]. shock-capturing [Sze89]. shocks [Kar80]. Shooting [MS84, Tro80]. Short [LL83a, Pom87, FP85]. side [Swa80]. Sided [Mar82b, EO81]. Sides [Saa87]. Sierpiński [BCW82, BCW81]. Sieve [Ger83b, Sil87]. Sign [KE84, te 87]. signal [Cyb83]. signature [God84a, God84b]. similar [BBG87]. Simple [Yam86b, CGT88, Kum88, PS85, Wun83, Yam86a, Zul88]. Simplest [Was87]. simplex [Gra88, LM80, Olm86]. Simplified [LL82, LL83b, New85]. Simulation [MR83b]. simultaneously [PT85]. Sinc [Kea83, Lun83, BS88b, Lun86]. sines [Sal84, SL85]. Single [Bal84a, BS80, ADT81, Bal84b, San88, IIMPL81]. Singular [Gar88b, Kel86, Ste84c, Tho81, AW84, BSC81, BS88b, Bru85, ET84, Eri85a, EN87, HC81, Ioa83, JS81, LR84, Mak81, Nas87, Rab86b, Sch81, TL81, Ven86, Wei84a, Wei86, dW85a]. Singular-Perturbation [Gar88b]. Singularities [MMM88, Yse86, Gra82, LS86, LM80, Sid80b]. Singularly [Gar87a, Gar87b, Gar88c, OS86a, OS86b, OO84, SW83, Wei84b]. Six [GM82, PH84, GO82]. Six-Dimensional [PH84]. sixth [Kat80]. size [JM85]. Sizes [Maj86b, Gol81, Maj86a]. Sketch [Hag83]. Skew [Bun82]. skipping [Alf80]. Small [Boy80a, Boy89a, Boy89b, Cha84, BF89b, Cha85, God84a, God84b]. smallest [Jae83, Jae85]. smear [JSW87]. Smooth [CW83, Swa89b, WH83, BS88b, BPST89, Bro84a, Swa89a]. Sobolev [ADT81, CQ82, DS80b]. Soc [II89]. solenoidal [Fri84]. Solution [Dic80, MW86b, Mis84, SS84, Wil80a, Alf80, ADT81, BT87, BS88b, Bör89, Bro84a, Bru85, Cas80, Cha82b, Cha87, CS80a, Cuy82, DF83, Eld82a, Eld82b, ER80, FO82, Gal83, Gla87, Gre82, HS81, Ioa83, JS81, LP86, Lun80, MW86a, MZM89, OO84, Pie88, Por85, Sch83a, Sch86, Spe87, Spi85, Toi86, Tsu87, Wei86, WWL81, Zlá83, Zlá86]. Solutions [Hud84, Luk88, McI86a, McI86b, Str89, Zaj83, Auz87, BL88, CL84b, FT84, Gau81, Hor80, Lue89, Mar80b, Mat87, Neu87, Puc89, Shu87a, SJI85, Vér82, Wri85, Zul88]. solvability [Mol87]. Solving [AEM87, Le 81b, Saa81, Saa84, SS85, Saa87, Yse86, BD81, CGT88, DL88b, EN85, Gol82, Hil82, Ven86]. Some [BW85, BBG88, Bor81, BM80, Bre82b, CW80, Dud87, Fet80, Fra82, Gat80, HT81, Kel82c, Ker83, Lav86, Lav87, Le 81a, LS89, McD89, Mey84, MS87, Oku83, Pai80, PW85, Que81, RS82, SW86a, Scr88, Sha86, Sha82b, SW88c, Und86, Vér82, VB81b, Was86, Wil81b, BR85, Bre82a, Bue84, Com89, DK85, Fet81a, FW80, Fra81a, FW85, JP82, KLL88, MZM89, Mos89, Nak84, PS83, Pri85, Sar87, SW86b, Sid82b, Wil87]. Soni [Pex81, Tem81, vK83a]. SOR [EL89]. Space [PH84, AD82, ADT81, AM89, ET84, EN87, Fri84, Gra83, LC88, OS83, Sze89]. space-time [LC88]. Spaced [LL82, For88]. Spaces [CH89, CT87b, SW82a, Wah84, CQ82, DS80b, Vér82]. spacings [Odl87]. sparse [DL88b, ER80, Luc83b, Sha80, Ste84a, Toi81]. spatial [San83]. spatially [CTW89, JLTW87, Ric81]. Special [Gra87b, How80, Mah82, AH81, Fet84b, Had82, Lue89, McD89, Pex81, Spe86, Tem81, vK83a]. spectra [Fer80]. Spectral [CGS89, Can88, FP84b, Hei89, NOPEJ87, Pas80, Dut89, ER80, IIMPL81, McL86]. spectrum [DLR81, Fis88, JE82, LC88]. speed [Jag82, Wil80c]. Speeding [Mon87]. Spline [CW83, Cos86, Mar82b, Arn83, SW85, dD85]. spline-trigonometric [Arn83]. Splines [CJW87, CDR88, DL88a, Met86, BC86, CH84, DM83, Gra88, JS81, Jia87, LA82, LA84, Luc83a, MNSS89]. split [LeV86]. Splitting [Ada84, BM81]. splittings [LO83]. Sporadic [Yam86b, Yam86a]. Springer [Fet81c, Fet84b, Kni88, Pex81, Sal81, Tem81, vK83a]. Square [Bro85, Hig86b, Sch85c, Lag83]. squared [DL81]. Squares [AKS85, Bue87b, LL86, LS81, MW86c, BL88, DL81, EGL80, Lag83, Mil87, Mit83b, Shi86, Shi19, Ano88b]. Stability [Ana80, Ber85, CT87b, Dut89, Got81a, GOT81b, Hic81, HV89, Le 81a, Le 81b, Maj86a, Maj86b, Mic83, Pia81, Pit80, Ske80, Tre85, CR82, Cuy82, Elm86, Fis88, GT81, GT85, GT87, JE82, Par80, Sch80a, Spi85]. stabilized [DW89]. Stable [Bun82, EO80, Gek82, Sha81, Sto80, Ana87, DM82a, FM85, GL85, Gra88, Noc89, Sha82a, Tad87, Win81]. stamp [Mos81]. Standards [Smi83]. State [Bra83, EG87]. States [BB85b]. Stationary [Lor86]. statistics [Bue84]. Steady [BB85b, EG87, Kar80]. Stefan [JR82, Whi83]. Steiner [MPR81, SF85, SS86b]. Stenger [BN82b]. Step [Bal84a, BS80, DD86a, DD86b, Har86, IN87, Maj86b, Bal84b, JM85, Maj86a, Spi85, Wer80]. Stepsize [Ske86b, Ske86c, Spi85]. Stieltjes [BB82, BF88, Boy82b, Cha80]. Stiff [Cas83b, Maj86b, Alf80, CS83a, Fat80, Ise81, Maj86a, OO84]. stochastic [Cha87]. Stokes [BDK82, BM81, BR85, CF89, DW89, DNW88, FT84, Gev89, Gir88, Gui87, JS86, KL85, PS85, Ste84b]. Stolarsky [Ste86a]. Storage [Noc80, SL84]. stream [CN86]. Streamline [JS86, JSW87, Sze89]. Stretching [Bea86b, Bea86a]. strictly [Mar80c]. strip [BvtW82, te 86c, te 86d, vt83, vtW86]. Strong [AS82, Leo80, Rot82]. strongly [HTE81, Sar87]. structure [BR88]. structures [Rei86]. Struve [New84]. study [Fai84, GN88, KFD83, Vrs86]. Sturm [Gar84, Nas87, Pai82]. Subcycling [Hic81]. subgroup [BR88, LQ88a, Sto81]. Subgroups [BC82a, PP80a, PP80b, PP80c, Sto81]. Subset [Lun88]. Subset-Sums [Lun88]. Subspace [Saa81]. substructures [BPS86b]. substructuring [BPS86a, BPS87, BPS88, BPS89]. subsystem [SS86b]. success [Bue84]. successive [Pen81]. Succinct [Bre82a]. such [Jae83, Jae85]. Sufficient [AS82]. Sum [Min80a, EGL80, Lag83, Leh88a, Wró84]. Summation [GM85b, Lav87, Sch83c, GM85a, Phi87]. Summations [Gla81]. Sums [CL84a, Hud84, Lun88, JS87a, Shi86, Shi19, SL89]. Superconvergence [ADT81, Kum88, LN89, OS86a, Tho80]. superconvergent [AW82]. superlinear [ST87]. Superspline [CH89]. Supplement [AB87b, AV85b, Bal84b, BD89b, Bau86a, Bea86b, Boy85b, Boy89b, BE88b, BC89b, BMS88a, BP89b, BSSW87, CGM86b, Car87a, Car88a, CL87b, CsAS87, DD86b, FO86b, Gar87a, GM85b, Gfr87b, GW85a, KG85b, Maj86b, MW86b, McI86b, NT87, PP85b, Sch84b, Ske86c, Swa89b, Yam86b, te 86b, vt85b]. Supra [KMS^{+}86]. Supra-Convergent [KMS^{+}86]. surface [LS85]. Surfaces [LS81, KBF87, Tsu86]. surprising [AD88]. Swan [FHS86]. Swinnerton [BGZ85]. Swinnerton-Dyer [BGZ85]. Sylow [LQ88a]. symbolic [CD87]. Symm [CES88]. Symmetric [Bun82, Saa87, Ser80, Swa86, ER80, MNR89, Ste84a]. symmetrical [Phi87].

Symmetrization [Lun86]. Symposium [Gau94]. system [Cuy82, Gar89, Wri85]. Systems [AKS85, BE88b, Hud84, Man80, Saa81, SS85, Saa87, AW84, BP88b, BP88c, BE88a, CR82, CR83, DL88b, FO82, Gal83, HS85, Ise81, Lus80, Mar80c, MPR81, OS82, Rhe84, SF85, SS86b, Tad83, Tad87, Vil88, Wer80, Win81, Zul88].

T [Has86, Stu81a, te 86c]. Table [Ano88b, Bac80, Bau86a, Bri82, Car87a, Car87b, Car88a, Car88b, Car89, Car80, Che83, Dav82, Fet81d, Fet81c, Fet81e, Fet81b, Fet82b, Fet82a, Fet84b, vK86, Kel82a, Kel82b, Kel82c, KP88, Kni88, Kru83, Lon83, Pex81, Phi87, Sal84, SL85, Smi83, Sol86, Stu81a, Tem81, vK82, vK83a, vK83c, vK83d, vK83e, vK83b, vK83f, vK83g, CS87c, McC83, WZ88]. Tables [Bau86b, Ben83, BMS88a, BMS88b, Dav82, KP88, Kni88, SL85, vK83f, vK83g, Lon83, Smi83, vK83b]. tangent [Ven86]. Tate [TZ87]. tau [CR83, OO84, FO82]. Taylor [Fra81b, Scr88, Wri81a, Wri81b]. te [te 86c, te 86d]. Technique [Man80, BP88b, BP88c, CN86, Cyb83, LC88, Mah81]. Techniques [AV85b, Pai80, Saa84, AV85a]. Ten [PP80c, Sal84, SL85]. term [Gau81, McC84b]. Terminal [MMM88]. termination [Iga84]. Test [CL87a, CL87b, Nie89, Shi87]. Testing [AL81, Bri80, CL84a, CGT88, AO83, BW87, Wun83]. Tests [AS82, Fra82, KSW86, PSS89, Wil87]. th [NN80]. their [BJVZ89, Gir87, Ioa85b, Lag83]. Theorem [Mee80, Mic87, Tza86, Boy82b, FHS86, McC84b, Ril83, SSP85, TW89, Sch83c, SW81]. theorems [Fet84b, Mah80, Pex81, Tem81, vK83a]. Theoretic [CT86, CT89, VB81b]. Theory [DD86a, DD86b, BHN82, Ger83a, Gir87, GV87, Las86, Lub83, Mic83, SSP85, Stu81b]. there [BB88]. Theta [Par83, BBG87]. thing [dS80a]. Third [Car88a, Car88b, Ada87, Gui87, Pex81, San88, Tem81]. third-order [Ada87, Gui87, San88]. Thomson [Sal80]. thousandths [Sal84, SL85]. Three [BM82a, BM82c, Sch83c, te 81, AJ81, BB88, BM82b, Gau81, Gui87]. three-term [Gau81]. Thue [Str89]. Tikhonov [Gfr87a, Gfr87b, Neu87]. Tikhonov-regularized [Neu87]. Time [Bal84a, Ben84a, Ben84b, DW83, Eld82a, Eld82b, Har86, HT81, AEM87, AD82, AM89, Bal84b, BDS85, CT87c, Cyb83, Gal83, Gev89, Kar86, Kee89, LeV86, LC88, OS83, Sch85a, Tad84a]. Time-Dependent [Bal84a, HT81, DW83, Bal84b, BDS85, Kar86, Kee89]. Time-Harmonic [Ben84a, Ben84b]. time-split [LeV86]. Time-Step [Har86]. Toeplitz [BP88a, Dic80]. Torii [Has86]. torsion [Rei86]. total [dD85]. Totally [ET85b, Smy84, Buc87, BW87, BF89b, BPv89, CS87c, LQ88b, Smy81]. transcendence [Bai88b]. Transcendental [Mis84, Fet81d, KLL88, vK83c, vK83d, vK83e]. transform [ALS89, For81, LG86]. Transformation [NOPEJ87, ER80, Sid80a]. Transformations [Del81, AJS83, CD87, Fie81, NS85, Sid80b]. Transforms [CT86, FMPT84, BB82, CT89, FW85, KP88, Mah80, vK83f, vK83g, Kni88]. transient [Alf80]. translates [Loc81]. translation [Car80, Dav82, Fet81e]. transonic [EO80, Fis88, Mos89, OHW85]. trapezoidal [Wei84b]. treatment [Shu87a]. Triads [LL86]. Trial [Mon85b]. Triangles [Adl83, Sty80]. triangular [FM85]. Tricomi [Fet81d, KP88, vK83c, vK83d, vK83e, vK83f, vK83g]. Trigonometric [Fet80, Fet84a, Arn83, Fet81a, Mas83, Sch80b]. Triperfect [Kis84, Kis85]. Triperfects [BN82a]. triple [MPR81, SF85, SS86b]. Trivial [Str89]. Truncation [TW83]. turbulence [BK89]. Turing [Bra83]. turning [AW84, BHK84, Kum88]. TVB [CS89, Shu87a, Shu87b]. twelfth [Coh82]. Twentieth [YB88]. Twenty [Eva83]. Twenty-Fourth [Eva83]. Two [AB80, Ben83, CD88, Coh88, CL84b, DM82a, Gar87a, IN87, LL86, Leh81, Mar82b, Mis84, Olm86, AG83, BM82b, BK89, CT87a, CN88, Coh82, CD87, Cot87, Dja81, ELT88, Fai84, Gaá88, Gar87b, Gar88c, JR82, Kat80, KFD83, KBF87, Lev82a, Mas89, SW83, Shi86, Shi19, Sze89, Tre85, Tsu86]. two-dimensional [BK89, CT87a, CN88, Cot87, Dja81, Lev82a, Mas89]. two-level [AG83]. two-phase [JR82]. Two-Point [Gar87a, Fai84, Gar87b, Gar88c, Kat80, KFD83]. Two-Sided [Mar82b]. Two-Step [IN87].

Two-Valued [Coh88]. Type [FO86b, FO86a, HT81, HW82, Lin86, NZ82, Sha81, Sha82a, Dyn84, Gol82, Gre82, HC81, KS87, Kum88, LW82, MZ88, Ioa83, Ioa85a]. Type-Insensitive [Sha81, Sha82a].

U.S [Lon83]. unbounded [Gol81, Gol82, HK87, LS86]. unconditional [Tad83]. Unified [Ste84b, Sid80b]. Uniform [Egg89, Gar87a, Gar87b, GR89b, Mun81, Ray88a, Ioa85b, Loc81, Ray88b]. Uniformly [Gar88b, BSC81, OS86b, Shu87b]. Unigrid [MR83b]. unilateral [SB89]. Uniqueness [Boj81, Cho82]. Unit [Mae86, Buc87, BW88, BPv89, Lew85b, Lew87]. Unitary [Hag81]. Units [BP89b, Gra87b, PZ82, PWZ82, BP89a, CS87c, Leh88b]. Univ [Stu81a]. unramified [Vau85a]. Unsymmetric [Saa81, EL89, PTL85]. update [Luc83b, Ste84a, Toi81]. Updates [DM82b, Tap88]. Updating [Noc80, Pai80, FM85]. upon [Nie83]. Upper [HLR88, Cha82a]. Upwind [Lor86, OS82]. Use [CD86, Gal83, Tap88]. used [Iga84]. user [Sid88]. user-friendly [Sid88]. Using [BOP80, Bai88a, BLM89, AJ81, BF89a, Bue84, CT89, Dur87, LA82, LA84].

V [Kel82c, PP80c]. validity [Tro89]. valuation [von84]. Value [AV85b, Ben84a, Ben84b, Bra83, Cas83a, Cas83b, Gar87a, GR89b, MW86b, Mee80, Mic87, Ste84c, Win80b, Alf80, Ana87, AV85a, BL88, CM87b, Dut89, ET84, EN87, Fai84, Gar87b, Gar88c, GT81, GT85, GT87, HV89, Ioa85b, II89, Kat80, KFD83, Las86, Lun86, MW86a, MR83a, Mas89, Mat87, Mic83, Mun81, OS86a, Oke87, OO84, Par80, Rab83, Rab85, Rab88b, Rab88a, RL89, Sch86, Spi85, Wei84a, Wei86, Wei84b, Win81, dW85a]. Valued [Coh88, GR81]. Values [Alf82, LL82, Mey84, NT87, Smy84, Fet81b, FT84, FW80, Fra81a]. Vancouver [Gau94]. Variable [Cho82, DW83, JM85, LL82, Ske86b, Ske86c, ALS89, ADT81, Fal83, GOT81b, KL85, San83, Sha80]. variable-metric [Sha80]. Variable-Stepsize [Ske86b, Ske86c]. Variables [CD88, Mar80a, Coh82, CD87, CGT88, Gaá88, Hab83]. variation [San88, TL85]. Variational [Gol80, Mit83a]. variationally [Toi81]. Varies [CL83]. varying [OS83, Ric81]. vaulting [Cor80]. vector [EHGL89, For81, Fri84]. Vectors [Ger88, FP85]. version [BS88a, EN85, New85, PS85, SS89]. versus [Wer84a]. vertices [LM80]. Very [CW80, Pom87, Sch80b, Kel82c, Sid82a]. via [EK89, Ven86]. Viscosity [DMG89a, Tad84b, Tad87]. viscous [BM81, CN86]. Vol [Fet81d, KP88, vK83c, vK83d, vK83e, vK83f, vK83g]. Volterra [vt85b, BJVZ89, BHN82, Bru84, Bru85, Bru89, Egg84, Egg89, Gla82, Gla87, JM85, Lub83, Lub85, Mak80, Mak81, Wol83, vt85a]. volume [Spe86]. Volumes [AMS84, AS86]. Voronoı [Wil81b]. Vortex [BBP81, BM82a, BM82b, Bea86b, Del89, Gre86, Bea86a, CN88, Cot87]. Vries [AW82, BDK83, DK85, Win80a].

W [Fet81d, Fet84b, Kel82b, KP88, Pex81, Tem81, Wil88, vK83a, vK83c, vK83d, vK83e, vK83f, vK83g]. W. [Ker83, Tza86]. walk [Rob89]. Wang [AD88]. Washington [Lon83, Smi83]. Waterflood [DW83]. Watson [Stu81a]. Wave [CGS89, Le 81b, ATZ83, Bro84a, EG87, Hal82, Hig86a, Hig87, SB89, TH86]. waveguides [Gol82]. way [TH86]. Weak [Ray88a, Ray88b, Sch80a]. weakly [Bru85, Mak81, Mik89, Sch81]. weakly-singular [Mak81]. Weber [Gra87a]. Weeks [LG86]. Weierstrass [Las82a]. Weight [För87, KE84, SW86b, Xu89, BS84, GN88]. Weighted [DMG89a, DMG89b, FO86b, FO86a, Ioa83, Toi81]. weights [LS83]. Well [TH86, IIMPL81]. Well-posedness [TH86]. Which [Bro85, Boy81, DM82a, Lag83, Olm86, dS81b]. Whittaker [Stu81a]. whose [Mah85]. Williams [Sha82b, Kel82c, Wil88]. Winter [te 86c]. Without [Mon85b, Ser80, AW84]. witnesses [EP86]. Wong [Mei85]. Wrigge [Fra81a]. Wright [Fet81b].

York [Car80, Che83, Dav82, Fet81d, Fet81c, Fet81e, Fet82a, KP88, Kni88, Kru83, Pex81, Sal81, Sal84, SL85, Sol86, Tem81, vK82, vK83a, vK83c, vK83d, vK83e, vK83b, vK83f, vK83g]. Yudell [GW84].

Zero [Pie84, Cha82a]. zeroes [LS83]. Zeros [Ben83, CS82, FMPT84, Lac84, Mah82, BvtW82, Gar84, Iga84, Kim88, Leh88a, Odl87, SSO83, te 86c, te 86d, vt83, vtW86]. zeta [Apo85, BvtW82, Eli86, Leh88a, Odl87, te 86c, te 86d, vt83, vtW86].

References

Allasia:1980:NET

[AB80] G. Allasia and F. Bonardo. On the numerical evaluation of two infinite products. Mathematics of Computation, 35(151):917–931, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Atkinson:1987:DGM

[AB87a] Kendall Atkinson and Alex Bogomolny.

The discrete Galerkin method for integral equations. Mathematics of Computation, 48(178):595–616, S11–S15, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Atkinson:1987:SDG

[AB87b] Kendall Atkinson and Alex Bogomolny.

Supplement to the discrete Galerkin method for integral equations. Mathematics of Computation, 48(178):S11– S15, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Agrawal:1980:ECC

[ACHv80] M. K. Agrawal, J. H. Coates, D. C.

Hunt, and A. J. van der Poorten. Elliptic curves of conductor 11. Mathematics of Computation, 35(151):991–1002, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Archer:1982:CGM

[AD82] David Archer and Julio César Díaz. A collocation-Galerkin method for a first order hyperbolic equation with space and time dependent coefficient. Mathematics of Computation, 38(157):37–53, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Abbott:1988:ASP

[AD88] J. A. Abbott and J. H. Davenport. Another surprising property of 42. A remark on a paper by P. S. Wang: “Factoring multivariate polynomials over algebraic number fields” [Math. Comp. 30 (1976), no. 134, 324–336, MR 58 #27887a]. Mathematics of Computation, 51(184):837–839, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Adams:1984:SQP

[Ada84] William W. Adams. Splitting of quartic polynomials. Mathematics of Computation, 43(167):329–343, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Adams:1987:CPT

[Ada87] William W. Adams. Characterizing pseudoprimes for third-order linear recurrences. Mathematics of Computation, 48(177):1–15, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Adler:1983:BMT

[Adl83] Andrew Adler. On the bisection method for triangles. Mathematics of Computation, 40(162):571–574, April 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Arnold:1981:SFE

[ADT81] Douglas N. Arnold, Jim Douglas, Jr., and Vidar Thomée. Superconvergence of a finite element approximation to the solution of a Sobolev equation in a single space variable. Mathematics of Computation, 36(153):53–63, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Adleman:1987:SBQ

[AEM87] Leonard M. Adleman, Dennis R. Estes, and Kevin S. McCurley. Solving bivariate quadratic congruences in random polynomial time. Mathematics of Computation, 48(177):17–28, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Axelsson:1983:PTL

[AG83] O. Axelsson and I. Gustafsson. Preconditioning and two-level multigrid methods of arbitrary degree of approximation. Mathematics of Computation, 40 (161):219–242, January 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Avdelas:1981:OAO

[AH81] G. Avdelas and A. Hadjidimos. Optimum accelerated overrelaxation method in a special case. Mathematics of Computation, 36(153):183–187, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ash:1981:OND

[AJ81] J. Marshall Ash and Roger L. Jones.

Optimal numerical differentiation using three function evaluations. Mathematics of Computation, 37(155):159–167, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Albertsen:1983:NTA

[AJS83] N. C. Albertsen, G. Jacobsen, and S. B. Sørensen. Nonlinear transformations for accelerating the convergence of M-dimensional series. Mathematics of Computation, 41(164):623–634, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Aziz:1981:FEA

[AK81] A. K. Aziz and R. Bruce Kellogg. Finite element analysis of a scattering problem. Mathematics of Computation, 37(156): 261–272, October 1981. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Akrivis:1985:ENC

[Akr85] G. Akrivis. The error norm of certain Gaussian quadrature formulae. Mathematics of Computation, 45(172):513– 519, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Aziz:1985:LSM

[AKS85] A. K. Aziz, R. B. Kellogg, and A. B. Stephens. Least squares methods for elliptic systems. Mathematics of Computation, 44(169):53–70, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Adleman:1981:PTA

[AL81] Leonard Adleman and Frank Thomson Leighton. An O(n^{1/10.89}) primality testing algorithm. Mathematics of Computation, 36(153):261–266, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Alex:1985:DE

[Ale85] Leo J. Alex. On the Diophantine equation 1 + 2^{a} = 3^{b}5^{c} + 2^{d}3^{e}5^{f}. Mathematics of Computation, 44(169):267– 278, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Alfeld:1980:MST

[Alf80] Peter Alfeld. A method of skipping the transient phase in the solution of separably stiff ordinary initial value problems. Mathematics of Computation, 35(152): 1173–1176, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Alfeld:1982:FPI

[Alf82] Peter Alfeld. Fixed point iteration with inexact function values. Mathematics of Computation, 38(157):87–98, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ang:1989:CVR

[ALS89] D. D. Ang, John Lund, and Frank Stenger. Complex variable and regularization methods of inversion of the Laplace transform. Mathematics of Computation, 53(188):589–608, October 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Aziz:1989:CFE

[AM89] A. K. Aziz and Peter Monk. Continuous finite elements in space and time for the heat equation. Mathematics of Computation, 52(186):255–274, April 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

AMS:1984:CIM

[AMS84] AMS. Cumulative Index to Mathematics of Computation, Volumes 24–43, 1970– 1984. Mathematics of Computation. American Mathematical Society, Providence, RI, USA, October 1984. CO-DEN MCMPAF. ISBN 0-8218-0093-0. ISSN 0025-5718 (print), 1088-6842 (electronic). vii + 503 pp.

Anand:1980:NSN

[Ana80] Indu Mati Anand. Numerical stability of nested dissection orderings. Mathematics of Computation, 35(152):1235– 1249, October 1980. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ananthakrishnaiah:1987:SOM

[Ana87]  U. Ananthakrishnaiah.  P -stable

Obrechkoff methods with minimal phase-lag for periodic initial value problems. Mathematics of Computation, 49 (180):553–559, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1980:C

[Ano80] Anonymous. Corrigenda. Mathematics of Computation, 34(150):652, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1988:C

[Ano88a] Anonymous. Corrigendum. Mathematics of Computation, 50(181):359, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1988:TEB

[Ano88b] Anonymous. Table errata: The book of squares [Academic Press, Boston, Mass., 1987, MR 88m:01021a] by L. P. Fibonacci. Mathematics of Computation, 50(182):654, April 1988. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Adleman:1983:ITF

[AO83] Leonard M. Adleman and Andrew M. Odlyzko. Irreducibility testing and factorization of polynomials. Mathematics of Computation, 41(164):699–709, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Apostol:1985:FHD

[Apo85] Tom M. Apostol. Formulas for higher derivatives of the Riemann zeta function. Mathematics of Computation, 44 (169):223–232, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Arnold:1983:STG

[Arn83]  Douglas N. Arnold.  A spline-

trigonometric Galerkin method and an exponentially convergent boundary integral method. Mathematics of Computation, 41(164):383–397, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Adams:1982:SPT

[AS82] William Adams and Daniel Shanks.

Strong primality tests that are not sufficient. Mathematics of Computation, 39(159):255–300, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Allgower:1986:CVP

[AS86] Eugene L. Allgower and Phillip H.

Schmidt. Computing volumes of polyhedra. Mathematics of Computation, 46 (173):171–174, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Arscott:1983:NCE

[ATZ83] F. M. Arscott, P. J. Taylor, and R. V. M.

Zahar. On the numerical construction of ellipsoidal wave functions. Mathematics of Computation, 40(161):367– 380, January 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Auzinger:1987:DCM

[Auz87] Winfried Auzinger. Defect corrections for multigrid solutions of the Dirichlet problem in general domains. Mathematics of Computation, 48(178):471– 484, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Axelsson:1985:BVT

[AV85a] A. O. H. Axelsson and J. G. Verwer. Boundary value techniques for initial value problems in ordinary differential equations. Mathematics of Computation, 45(171):153–171, S1–S4, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Axelsson:1985:SBV

[AV85b] A. O. H. Axelsson and J. G. Verwer. Supplement to boundary value techniques for initial value problems in ordinary differential equations. Mathematics of Computation, 45(171):S1–S4, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Arnold:1982:SFE

[AW82] Douglas N. Arnold and Ragnar Winther. A superconvergent finite element method for the Korteweg–de Vries equation. Mathematics of Computation, 38 (157):23–36, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Arnold:1983:ACC

[AW83] Douglas N. Arnold and Wolfgang L. Wendland. On the asymptotic convergence of collocation methods. Mathematics of Computation, 41(164):349– 381, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ascher:1984:CSP

[AW84] U. Ascher and R. Weiss. Collocation for singular perturbation problems. II. Linear first order systems without turning points. Mathematics of Computation, 43(167):157–187, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Afflerbach:1989:EDR

[AW89] Lothar Afflerbach and Rainer Weilbächer.

The exact determination of rectangle discrepancy for linear congruential pseudorandom numbers. Mathematics of Computation, 53(187):343–354, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Backstrom:1980:TEM

[Bac80] Robert P. Backstrom. Table errata: “A method of factoring and the factorization of F_{7}” [Math. Comp. 29 (1975), 183–205, MR 51 #8017] by M. A. Morrison and J. Brillhart. Mathematics of Computation, 35(152):1444, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bailey:1988:CDD

[Bai88a] David H. Bailey. The computation of π to 29,360,000 decimal digits using Borweins’ quartically convergent algorithm. Mathematics of Computation, 50(181): 283–296, January 1988. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bailey:1988:NRT

[Bai88b] David H. Bailey. Numerical results on the transcendence of constants involving π, e, and Euler’s constant. Mathematics of Computation, 50(181):275– 281, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bales:1984:SSSa

[Bal84a] Laurence A. Bales. Semidiscrete and single step fully discrete approximations for second order hyperbolic equations with time-dependent coefficients. Mathematics of Computation, 43(168):383– 414, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bales:1984:SSSb

[Bal84b] Laurence A. Bales. Supplement to: “Semidiscrete and single step fully discrete approximations for second order hyperbolic equations with timedependent coefficients”. Mathematics of Computation, 43(168):S1–S12, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bauer:1986:STR

[Bau86a] Helmut F. Bauer. Supplement to table of the roots of the associated Legendre function with respect to the degree. Mathematics of Computation, 46(174): S29–S41, April 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bauer:1986:TRA

[Bau86b] Helmut F. Bauer. Tables of the roots of the associated Legendre function with respect to the degree. Mathematics of Computation, 46(174):601–602, S29– S41, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Belevitch:1982:SIT

[BB82] V. Belevitch and J. Boersma. On Stieltjes integral transforms involving Γ-functions. Mathematics of Computation, 38(157):223–226, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bigge:1985:DBD

[BB85a] J. Bigge and E. Bohl. Deformations of the bifurcation diagram due to discretization. Mathematics of Computation, 45(172):393–403, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bigge:1985:SSF

[BB85b] J. Bigge and E. Bohl. On the steady states of finitely many chemical cells. Mathematics of Computation, 44(170): 405–415, April 1985. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Borwein:1986:MQC

[BB86] J. M. Borwein and P. B. Borwein. More quadratically converging algorithms for π. Mathematics of Computation, 46 (173):247–253, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Battiato:1988:TOA

[BB88] S. Battiato and W. Borho. Are there odd amicable numbers not divisible by three. Mathematics of Computation, 50(182): 633–637, April 1988. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Borwein:1989:MI

[BB89] J. M. Borwein and P. B. Borwein. On the mean iteration (a, b) ← (\frac{a+3b}{4}, \frac{\sqrt{ab}+b}{2}). Mathematics of Computation, 53(187): 311–326, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bunse:1982:CBP

[BBG82] W. Bunse and A. Bunse-Gerstner. Computation of bounds for the positive eigenvector of a nonnegative irreducible matrix by monotone iteration. Mathematics of Computation, 39(159):125– 131, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Blecksmith:1986:CAI

[BBG86] Richard Blecksmith, John Brillhart, and Irving Gerst. A computer-assisted investigation of Ramanujan pairs. Mathematics of Computation, 46(174):731– 749, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Blecksmith:1987:PRC

[BBG87] Richard Blecksmith, John Brillhart, and Irving Gerst. Parity results for certain partition functions and identities similar to theta function identities. Mathematics of Computation, 48(177):29–38, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Blecksmith:1988:SIP

[BBG88] Richard Blecksmith, John Brillhart, and Irving Gerst. Some infinite product identities. Mathematics of Computation, 51(183):301–314, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bardos:1981:VMF

[BBP81] Claude Bardos, Michel Bercovier, and Olivier Pironneau. The vortex method with finite elements. Mathematics of Computation, 36(153):119–136, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Butler:1982:CPMa

[BC82a] Gregory Butler and John J. Cannon. Computing in permutation and matrix groups. I. normal closure, commutator subgroups, series. Mathematics of Computation, 39(160):663–670, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Butler:1982:CPMb

[BC82b] Gregory Butler and John J. Cannon.

Computing in permutation and matrix groups. II. Backtrack algorithm. Mathematics of Computation, 39(160):671– 680, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bremner:1984:E

[BC84] A. Bremner and J. W. S. Cassels. On the equation Y^{2} = X(X^{2} + p). Mathematics of Computation, 42(165):257– 264, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bini:1986:CCS

[BC86] D. Bini and M. Capovani. A class of cubic splines obtained through minimum conditions. Mathematics of Computation, 46(173):191–202, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brent:1989:NLB

[BC89a] Richard P. Brent and Graeme L. Cohen. A new lower bound for odd perfect numbers. Mathematics of Computation, 53(187):431–437, S7–S24, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brent:1989:SNL

[BC89b] Richard P. Brent and Graeme L. Cohen. Supplement to A new lower bound for odd perfect numbers. Mathematics of Computation, 53(187):S7–S24, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buhler:1982:PF

[BCP82] J. P. Buhler, R. E. Crandall, and M. A. Penk. Primes of the form n! ± 1 and 2 · 3 · 5 ⋯ p ± 1. Mathematics of Computation, 38(158):639–643, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Baillie:1981:PSC

[BCW81] Robert Baillie, G. Cormack, and H. C. Williams. The problem of Sierpiński concerning k · 2^{n} + 1. Mathematics of Computation, 37(155):229–231, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Baillie:1982:CPS

[BCW82] Robert Baillie, G. Cormack, and H. C. Williams. Corrigenda: “The problem of Sierpiński concerning k · 2^{n} + 1” [Math. Comp. 37 (1981), no. 155, 229–231]. Mathematics of Computation, 39(159): 308, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Baker:1980:CGA

[BD80] Garth A. Baker and Vassilios A. Dougalis. On the L^{∞}-convergence of Galerkin approximations for secondorder hyperbolic equations. Mathematics of Computation, 34(150):401– 424, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bank:1981:OOP

[BD81] Randolph E. Bank and Todd Dupont. An optimal order process for solving finite element equations. Mathematics of Computation, 36(153):35–51, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bales:1989:CMN

[BD89a] Laurence A. Bales and Vassilios A.

Dougalis. Cosine methods for nonlinear second-order hyperbolic equations. Mathematics of Computation, 52(186): 299–319, S15–S33, April 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bales:1989:SCM

[BD89b] Laurence A. Bales and Vassilios A. Dougalis. Supplement to cosine methods for nonlinear second-order hyperbolic equations. Mathematics of Computation, 52 (186):S15–S33, April 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Baker:1982:HOA

[BDK82] Garth A. Baker, Vassilios A. Dougalis, and Ohannes A. Karakashian. On a higher order accurate fully discrete Galerkin approximation to the Navier–Stokes equations. Mathematics of Computation, 39(160):339–375, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Baker:1983:CGA

[BDK83] Garth A. Baker, Vassilios A. Dougalis, and Ohannes A. Karakashian. Convergence of Galerkin approximations for the Korteweg–de Vries equation. Mathematics of Computation, 40(162):419– 433, April 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bales:1985:CMS

[BDS85] Laurence A. Bales, Vassilios A. Dougalis, and Steven M. Serbin. Cosine methods for second-order hyperbolic equations with time-dependent coefficients.

Mathematics of Computation, 45(171): 65–89, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brenan:1988:BDA

[BE88a] Kathryn E. Brenan and Björn E. Engquist. Backward differentiation approximations of nonlinear differential/ algebraic systems. Mathematics of Computation, 51(184):659–676, S7–S16, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brenan:1988:SBD

[BE88b] Kathryn E. Brenan and Bjorn E. Engquist. Supplement to backward differentiation approximations of nonlinear differential/algebraic systems. Mathematics of Computation, 51(184):S7–S16, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Beale:1986:CDV

[Bea86a] J. Thomas Beale. A convergent 3-D vortex method with grid-free stretching. Mathematics of Computation, 46(174): 401–424, S15–S20, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Beale:1986:SCD

[Bea86b] J. Thomas Beale. Supplement to A convergent 3-D vortex method with gridfree stretching. Mathematics of Computation, 46(174):S15–S20, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Benton:1983:CZT

[Ben83] T. C. Benton. Common zeros of two Bessel functions. part II. approximations and tables. Mathematics of Computation, 41(163):203–217, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bendali:1984:NAEa

[Ben84a] A. Bendali. Numerical analysis of the exterior boundary value problem for the time-harmonic Maxwell equations by a boundary finite element method. part 1: The continuous problem. Mathematics of Computation, 43(167):29–46, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bendali:1984:NAEb

[Ben84b] A. Bendali. Numerical analysis of the exterior boundary value problem for the time-harmonic Maxwell equations by a boundary finite element method. part 2: The discrete problem. Mathematics of Computation, 43(167):47–68, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Berger:1985:SIM

[Ber85] Marsha J. Berger. Stability of interfaces with mesh refinement. Mathematics of Computation, 45(172):301–318, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brezzi:1986:NAM

[BF86] F. Brezzi and M. Fortin. Numerical approximation of Mindlin–Reissner plates. Mathematics of Computation, 47(175): 151–158, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bohman:1988:SFD

[BF88] Jan Bohman and Carl-Erik Fröberg. The Stieltjes function—definition and properties. Mathematics of Computation, 51(183):281–289, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bailey:1989:NRR

[BF89a] David H. Bailey and Helaman R. P. Ferguson. Numerical results on relations between fundamental constants using a new algorithm. Mathematics of Computation, 53(188):649–656, October 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buchmann:1989:CTR

[BF89b] Johannes Buchmann and David Ford. On the computation of totally real quartic fields of small discriminant. Mathematics of Computation, 52(185):161– 174, January 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boley:1984:MMR

[BG84] Daniel Boley and Gene H. Golub. A modified method for reconstructing periodic Jacobi matrices. Mathematics of Computation, 42(165):143–150, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buhler:1985:CBS

[BGZ85] Joe P. Buhler, Benedict H. Gross, and Don B. Zagier. On the conjecture of Birch and Swinnerton-Dyer for an elliptic curve of rank 3. Mathematics of Computation, 44(170):473–481, April 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Borho:1986:BAN

[BH86] W. Borho and H. Hoffmann. Breeding amicable numbers in abundance. Mathematics of Computation, 46(173):281– 293, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Berger:1984:PEA

[BHK84] Alan E. Berger, Hou De Han, and R. Bruce Kellogg. A priori estimates and analysis of a numerical method for a turning point problem. Mathematics of Computation, 42(166):465–492, April 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brunner:1982:RKT

[BHN82] H. Brunner, E. Hairer, and S. P. Nørsett.

Runge–Kutta theory for Volterra integral equations of the second kind. Mathematics of Computation, 39(159):147– 163, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bustoz:1986:GFI

[BI86] Joaquin Bustoz and Mourad E. H. Ismail. On gamma function inequalities. Mathematics of Computation, 47(176): 659–667, October 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bellen:1989:NCE

[BJVZ89] A. Bellen, Z. Jackiewicz, R. Vermiglio, and M. Zennaro. Natural continuous extensions of Runge–Kutta methods for Volterra integral equations of the second kind and their applications. Mathematics of Computation, 52(185):49–63, January 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Browning:1989:CNM

[BK89] G. L. Browning and H.-O. Kreiss. Comparison of numerical methods for the calculation of two-dimensional turbulence. Mathematics of Computation, 52(186): 369–388, April 1989. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bohmer:1988:AED

[BL88] Klaus Böhmer and John Locker. Asymptotic expansions for the discretization error of least squares solutions of linear boundary value problems. Mathematics of Computation, 51(183):75–91, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Blair:1982:CRC

[Bla82] J. M. Blair. Corrigenda: “Rational Chebyshev approximations for the Bickley functions Ki_{n}(x)” [Math. Comp. 32 (1978), no. 143, 876–886 and C. A. Edwards MR 57 #10999] by the author and J. H. Johnson. Mathematics of Computation, 38(158):657, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Banerjee:1987:AEI

[BLL87] U. Banerjee, L. J. Lardy, and A. Lutoborski. Asymptotic expansions of integrals of certain rapidly oscillating functions. Mathematics of Computation, 49(179):243–249, July 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bayliss:1989:SPU

[BLM89] Alvin Bayliss, Yanyan Li, and Cathleen Synge Morawetz. Scattering by a potential using hyperbolic methods. Mathematics of Computation, 52(186): 321–338, April 1989. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brent:1980:SNA

[BM80] Richard P. Brent and Edwin M. McMillan. Some new algorithms for highprecision computation of Euler’s constant. Mathematics of Computation, 34 (149):305–312, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Beale:1981:RCV

[BM81] J. Thomas Beale and Andrew Majda.

Rates of convergence for viscous splitting of the Navier–Stokes equations. Mathematics of Computation, 37(156):243– 259, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Beale:1982:VMC

[BM82a] J. Thomas Beale and Andrew Majda.

Vortex methods. I: Convergence in three dimensions. Mathematics of Computation, 39(159):1–27, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Beale:1982:VMI

[BM82b] J. Thomas Beale and Andrew Majda.

Vortex methods. II. higher order accuracy in two and three dimensions. Mathematics of Computation, 39(159): 29–52, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bremner:1982:IPT

[BM82c] Andrew Bremner and Patrick Morton.

The integer points on three related elliptic curves. Mathematics of Computation, 39(159):235–238, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brillhart:1988:STF

[BMS88a] John Brillhart, Peter L. Montgomery, and Robert D. Silverman. Supplement to tables of Fibonacci and Lucas factorizations. Mathematics of Computation, 50(181):S1–S15, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brillhart:1988:TFL

[BMS88b] John Brillhart, Peter L. Montgomery, and Robert D. Silverman. Tables of Fibonacci and Lucas factorizations. Mathematics of Computation, 50(181):251– 260, S1–S15, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Beck:1982:LBO

[BN82a] Walter E. Beck and Rudolph M. Najar. A lower bound for odd triperfects. Mathematics of Computation, 38(157): 249–251, January 1982. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Beighton:1982:EES

[BN82b] S. Beighton and B. Noble. An error estimate for Stenger’s quadrature formula. Mathematics of Computation, 38(158): 539–545, April 1982. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Babuska:1989:FEG

[BO89] I. Babuška and J. E. Osborn. Finite element-Galerkin approximation of the eigenvalues and eigenvectors of selfadjoint problems. Mathematics of Computation, 52(186):275–297, April 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bojanov:1981:UON

[Boj81] Borislav D. Bojanov. Uniqueness of the optimal nodes of quadrature formulae. Mathematics of Computation, 36(154): 525–546, April 1981. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Babuska:1980:AMM

[BOP80] I. Babuška, J. Osborn, and J. Pitkäranta.

Analysis of mixed methods using mesh dependent norms. Mathematics of Computation, 35(152):1039–1062, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Borho:1981:SLP

[Bor81] W. Borho. Some large primes and amicable numbers. Mathematics of Computation, 36(153):303–304, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Borgers:1989:NSR

[Bör89] Christoph Börgers. On the numerical solution of the regularized Birkhoff equations. Mathematics of Computation, 53(187):141–156, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boyd:1980:RPH

[Boy80a] David W. Boyd. Reciprocal polynomials having small measure. Mathematics of Computation, 35(152):1361– 1377, October 1980. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boyd:1980:RCH

[Boy80b] John P. Boyd. The rate of convergence of Hermite function series. Mathematics of Computation, 35(152):1309– 1316, October 1980. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boyd:1981:RCC

[Boy81] John P. Boyd. The rate of convergence of Chebyshev polynomials for functions which have asymptotic power series about one endpoint. Mathematics of Computation, 37(155):189–195, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boyd:1982:SRA

[Boy82a] David W. Boyd. The sequence of radii of the Apollonian packing. Mathematics of Computation, 39(159):249–254, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boyd:1982:CPR

[Boy82b] John P. Boyd. A Chebyshev polynomial rate-of-convergence theorem for Stieltjes functions. Mathematics of Computation, 39(159):201–206, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boyd:1984:PNN

[Boy84] David W. Boyd. Pisot numbers in the neighborhood of a limit point. II. Mathematics of Computation, 43(168):593– 602, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boyd:1985:MMA

[Boy85a] David W. Boyd. The maximal modulus of an algebraic integer. Mathematics of Computation, 45(171):243–249, S17– S20, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boyd:1985:SMM

[Boy85b] David W. Boyd. Supplement to the maximal modulus of an algebraic integer. Mathematics of Computation, 45(171): S17–S20, July 1985. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boyd:1989:RPH

[Boy89a] David W. Boyd. Reciprocal polynomials having small measure. II. Mathematics of Computation, 53(187):355– 357, S1–S5, July 1989. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boyd:1989:SRP

[Boy89b] David W. Boyd. Supplement to reciprocal polynomials having small measure. II. Mathematics of Computation, 53 (187):S1–S5, July 1989. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brent:1981:FEF

[BP81] Richard P. Brent and John M. Pollard. Factorization of the eighth Fermat number. Mathematics of Computation, 36(154):627–630, April 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1987:NCE

[BP87] James H. Bramble and Joseph E. Pasciak. New convergence estimates for multigrid algorithms. Mathematics of Computation, 49(180):311–329, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bini:1988:EAE

[BP88a] D. Bini and V. Pan. Efficient algorithms for the evaluation of the eigenvalues of (block) banded Toeplitz matrices. Mathematics of Computation, 50(182):431– 448, April 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1988:CPT

[BP88b] James H. Bramble and Joseph E. Pasciak. Corrigenda: “A preconditioning technique for indefinite systems resulting from mixed approximations of elliptic problems”. Mathematics of Computation, 51(183):387–388, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1988:PTI

[BP88c] James H. Bramble and Joseph E. Pasciak. A preconditioning technique for indefinite systems resulting from mixed approximations of elliptic problems. Mathematics of Computation, 50(181):1–17, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buchmann:1989:CIU

[BP89a] Johannes Buchmann and Attila Pethö. Computation of independent units in number fields by Dirichlet’s method. Mathematics of Computation, 52(185): 149–159, S1–S14, January 1989. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buchmann:1989:SCI

[BP89b] Johannes Buchmann and Attila Petho. Supplement to computation of independent units in number fields by Dirichlet’s method. Mathematics of Computation, 52(185):S1–S14, January 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1986:CPE

[BPS86a] James H. Bramble, Joseph E. Pasciak, and Alfred H. Schatz. The construction of preconditioners for elliptic problems by substructuring. I. Mathematics of Computation, 47(175):103–134, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1986:IME

[BPS86b] James H. Bramble, Joseph E. Pasciak, and Alfred H. Schatz. An iterative method for elliptic problems on regions partitioned into substructures. Mathematics of Computation, 46(174):361– 369, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1987:CPE

[BPS87] James H. Bramble, Joseph E. Pasciak, and Alfred H. Schatz. The construction of preconditioners for elliptic problems by substructuring, II. Mathematics of Computation, 49(179):1–16, July 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1988:CPE

[BPS88] James H. Bramble, Joseph E. Pasciak, and Alfred H. Schatz. The construction of preconditioners for elliptic problems by substructuring. III. Mathematics of Computation, 51(184):415–430, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1989:CPE

[BPS89] James H. Bramble, Joseph E. Pasciak, and Alfred H. Schatz. The construction of preconditioners for elliptic problems by substructuring. IV. Mathematics of Computation, 53(187):1–24, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1989:IIM

[BPST89] James H. Bramble, Joseph E. Pasciak, Peter H. Sammon, and Vidar Thomée. Incomplete iterations in multistep backward difference methods for parabolic problems with smooth and nonsmooth data. Mathematics of Computation, 52 (186):339–367, April 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buchmann:1989:CUG

[BPv89] J. Buchmann, M. Pohst, and J. von Schmettow. On the computation of unit groups and class groups of totally real quartic fields. Mathematics of Computation, 53(187):387–397, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1988:AMA

[BPX88] James H. Bramble, Joseph E. Pasciak, and Jin Chao Xu. The analysis of multigrid algorithms for nonsymmetric and indefinite elliptic problems. Mathematics of Computation, 51(184):389– 414, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bank:1982:AMI

[BR82] Randolph E. Bank and Donald J. Rose. Analysis of a multilevel iterative method for nonlinear finite element equations. Mathematics of Computation, 39(160): 453–465, October 1982. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bernardi:1985:ASF

[BR85] Christine Bernardi and Geneviève Raugel. Analysis of some finite elements for the Stokes problem. Mathematics of Computation, 44(169):71–79, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buekenhout:1988:SSM

[BR88] Francis Buekenhout and Sarah Rees. The subgroup structure of the Mathieu group M_{12}. Mathematics of Computation, 50(182):595–605, April 1988.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1981:LMM

[Bra81] James H. Bramble. The Lagrange multiplier method for Dirichlet’s problem. Mathematics of Computation, 37(155): 1–11, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brady:1983:DVR

[Bra83] Allen H. Brady. The determination of the value of Rado’s noncomputable function Σ(k) for four-state Turing machines. Mathematics of Computation, 40(162): 647–665, April 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Braess:1984:CRM

[Bra84] Dietrich Braess. The convergence rate of a multigrid method with Gauss– Seidel relaxation for the Poisson equation. Mathematics of Computation, 42 (166):505–519, April 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brent:1980:FOC

[Bre80] Richard P. Brent. The first occurrence of certain large prime gaps. Mathematics of Computation, 35(152):1435– 1436, October 1980. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brent:1982:SPP

[Bre82a] Richard P. Brent. Succinct proofs of primality for the factors of some Fermat numbers. Mathematics of Computation, 38(157):253–255, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brezinski:1982:SNC

[Bre82b] Claude Brezinski. Some new convergence acceleration methods. Mathematics of Computation, 39(159):133– 145, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brenner:1989:OOM

[Bre89a] Susanne C. Brenner. An optimal-order multigrid method for P1 nonconforming finite elements. Mathematics of Computation, 52(185):1–15, January 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brezinski:1989:PEE

[Bre89b] Claude Brezinski. Procedures for estimating the error in Padé approximation. Mathematics of Computation, 53(188): 639–648, October 1989. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brillhart:1980:NIT

[Bri80] John Brillhart. Note on irreducibility testing. Mathematics of Computation, 35(152):1379–1381, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brillhart:1982:TEN

[Bri82] John Brillhart. Table errata: “New primality criteria and factorizations of 2^{m}± and1” [Math. Comp. 29 (1975), 620–647 and D. H. Lehmer MR 52 #5546] by the author and J. L. Selfridge. Mathematics of Computation, 39(160):747, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brown:1984:NNS

[Bro84a] David L. Brown. A note on the numerical solution of the wave equation with piecewise smooth coefficients. Mathematics of Computation, 42(166):369– 391, April 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brown:1984:DE

[Bro84b] M. L. Brown. On the Diophantine equation ∑ X_{i} = ∏ X_{i}. Mathematics of Computation, 42(165):239–240, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brown:1985:SWA

[Bro85] Ezra Brown. Sets in which xy + k is always a square. Mathematics of Computation, 45(172):613–620, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brunotte:1982:CCM

[Bru82] Horst Brunotte. The computation of a certain metric invariant of an algebraic number field. Mathematics of Computation, 38(158):627–632, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brunner:1984:IRK

[Bru84] Hermann Brunner. Implicit Runge– Kutta methods of optimal order for Volterra integro-differential equations. Mathematics of Computation, 42(165): 95–109, January 1984. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brunner:1985:NSW

[Bru85] Hermann Brunner. The numerical solution of weakly singular Volterra integral equations by collocation on graded meshes. Mathematics of Computation, 45(172):417–437, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brunner:1989:CMN

[Bru89] Hermann Brunner. Collocation methods for nonlinear Volterra integrodifferential equations with infinite delay. Mathematics of Computation, 53(188): 571–587, October 1989. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1980:EHO

[BS80] James H. Bramble and Peter H. Sammon. Efficient higher order single step methods for parabolic problems: Part I. Mathematics of Computation, 35(151): 655–677, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Byrd:1984:CQR

[BS84] Paul F. Byrd and Lawrence Stalla. Chebyshev quadrature rules for a new class of weight functions. Mathematics of Computation, 42(165):173–181, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Babuska:1988:VFE

[BS88a] I. Babuška and Manil Suri. The pversion of the finite element method for constraint boundary conditions. Mathematics of Computation, 51(183):1– 13, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bialecki:1988:SNM

[BS88b] Bernard Bialecki and Frank Stenger. Sinc–Nyström method for numerical solution of one-dimensional Cauchy singular integral equation given on a smooth arc in the complex plane. Mathematics of Computation, 51(183):133–165, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bach:1989:FCP

[BS89] Eric Bach and Jeffrey Shallit. Factoring with cyclotomic polynomials. Mathematics of Computation, 52(185):201– 219, January 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Berger:1980:GOS

[BSC^{+}80] Alan E. Berger, Jay M. Solomon, Melvyn Ciment, Stephen H. Leventhal, and Bernard C. Weinberg. Generalized OCI schemes for boundary layer problems. Mathematics of Computation, 35(151):695–731, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Berger:1981:AUA

[BSC81] Alan E. Berger, Jay M. Solomon, and Melvyn Ciment. An analysis of a uniformly accurate difference method for a singular perturbation problem. Mathematics of Computation, 37(155):79– 94, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buck:1987:CNO

[BSSW87] Nicholas Buck, Lones Smith, Blair K.

Spearman, and Kenneth S. Williams. The cyclotomic numbers of order fifteen. with microfiche supplement. Mathematics of Computation, 48(177):67–83, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bremner:1983:IP

[BT83] Andrew Bremner and Nicholas Tzanakis. Integer points on y^{2} = x^{3} − 7x + 10. Mathematics of Computation, 41(164): 731–741, October 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Berrut:1987:ENM

[BT87] Jean-Paul Berrut and Manfred R. Trummer. Equivalence of Nyström’s method and Fourier methods for the numerical solution of Fredholm integral equations. Mathematics of Computation, 48(178): 617–623, April 1987. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buchmann:1987:CFU

[Buc87] Johannes Buchmann. The computation of the fundamental unit of totally complex quartic orders. Mathematics of Computation, 48(177):39–54, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buell:1984:ESU

[Bue84] Duncan A. Buell. The expectation of success using a Monte Carlo factoring method—some statistics on quadratic class numbers. Mathematics of Computation, 43(167):313–327, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buell:1987:CGQ

[Bue87a] Duncan A. Buell. Class groups of quadratic fields. II. Mathematics of Computation, 48(177):85–93, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buell:1987:ISC

[Bue87b] Duncan A. Buell. Integer squares with constant second difference. Mathematics of Computation, 49(180):635–644, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bultheel:1980:RAM

[Bul80] Adhemar Bultheel. Recursive algorithms for the matrix Padé problem. Mathematics of Computation, 35(151):875– 892, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bunch:1982:NSD

[Bun82] James R. Bunch. A note on the stable decomposition of skew symmetric matrices. Mathematics of Computation, 38(158):475–479, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brent:1982:ZRZ

[BvtW82] R. P. Brent, J. van de Lune, H. J. J.

te Riele, and D. T. Winter. On the zeros of the Riemann zeta function in the critical strip. II. Mathematics of Computation, 39(160):681–688, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Baillie:1980:LP

[BW80] Robert Baillie and Samuel S. Wagstaff, Jr. Lucas pseudoprimes. Mathematics of Computation, 35(152):1391–1417, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bank:1985:SPE

[BW85] R. E. Bank and A. Weiser. Some a posteriori error estimators for elliptic partial differential equations. Mathematics of Computation, 44(170):283–301, April 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buchmann:1987:PIT

[BW87] Johannes Buchmann and H. C. Williams.

On principal ideal testing in totally complex quartic fields and the determination of certain cyclotomic constants. Mathematics of Computation, 48(177):55– 66, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buchmann:1988:IPI

[BW88] Johannes Buchmann and H. C. Williams. On the infrastructure of the principal ideal class of an algebraic number field of unit rank one. Mathematics of Computation, 50(182):569–579, April 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buchmann:1989:CCN

[BW89] Johannes Buchmann and H. C. Williams. On the computation of the class number of an algebraic number field. Mathematics of Computation, 53(188):679– 688, October 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cantor:1987:CJH

[Can87] David G. Cantor. Computing in the Jacobian of a hyperelliptic curve. Mathematics of Computation, 48(177):95–101, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Canuto:1988:SMM

[Can88] Claudio Canuto. Spectral methods and a maximum principle. Mathematics of Computation, 51(184):615–629, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

CarlsonGallas:1980:TET

[Car80] Jason A. Carlson Gallas. Table errata: Table of integrals, series Gallas and products [English translation of the fourth Russian edition, Academic Press, New York, 1965 and MR 33 #5952] by I. S. Gradšteĭn and I. M. Ryžik. Mathematics of Computation, 35(152): 1444, October 1980. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carlson:1987:STE

[Car87a] B. C. Carlson. Supplement to A table of elliptic integrals of the second kind. Mathematics of Computation, 49(180): S13–S17, October 1987. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carlson:1987:TEI

[Car87b] B. C. Carlson. A table of elliptic integrals of the second kind. Mathematics of Computation, 49(180):595–606, S13– S17, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carlson:1988:STE

[Car88a] B. C. Carlson. Supplement to A table of elliptic integrals of the third kind. Mathematics of Computation, 51(183): S1–S5, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carlson:1988:TEI

[Car88b] B. C. Carlson. A table of elliptic integrals of the third kind. Mathematics of Computation, 51(183):267–280, S1– S5, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carlson:1989:TEI

[Car89] B. C. Carlson. A table of elliptic integrals: Cubic cases. Mathematics of Computation, 53(187):327–333, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cash:1980:NOA

[Cas80] J. R. Cash. A note on Olver’s algorithm for the solution of second-order linear difference equations. Mathematics of Computation, 35(151):767–772, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cash:1983:BRKa

[Cas83a] J. R. Cash. Block Runge–Kutta methods for the numerical integration of initial value problems in ordinary differential equations. part I. the nonstiff case. Mathematics of Computation, 40(161): 175–191, January 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cash:1983:BRKb

[Cas83b] J. R. Cash. Block Runge–Kutta methods for the numerical integration of initial value problems in ordinary differential equations. part II. the stiff case. Mathematics of Computation, 40(161): 193–206, January 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cassels:1985:NDE

[Cas85] J. W. S. Cassels. A note on the Diophantine equation x^{3} + y^{3} + z^{3} = 3. Mathematics of Computation, 44(169): 265–266, January 1985. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohn:1986:UCS

[CD86] Harvey Cohn and Jesse Deutsch. Use of a computer scan to prove Q(\sqrt{2 + \sqrt{2}}) and Q(\sqrt{3 + \sqrt{2}}) are Euclidean. Mathematics of Computation, 46(173):295– 299, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohn:1987:ASM

[CD87] Harvey Cohn and Jesse Deutsch. Application of symbolic manipulation to Hecke transformations of modular forms in two variables. Mathematics of Computation, 48(177):139–146, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohn:1988:EME

[CD88] Harvey Cohn and Jesse Deutsch. An explicit modular equation in two variables for Q(\sqrt{3}). Mathematics of Computation, 50(182):557–568, April 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chui:1988:IMS

[CDR88] Charles K. Chui, Harvey Diamond, and Louise A. Raphael. Interpolation by multivariate splines. Mathematics of Computation, 51(183):203–218, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Costabel:1988:CCM

[CES88] M. Costabel, V. J. Ervin, and E. P.

Stephan. On the convergence of collocation methods for Symm’s integral equation on open curves. Mathematics of Computation, 51(183):167–179, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Crouzeix:1989:NFE

[CF89] Michel Crouzeix and Richard S. Falk.

Nonconforming finite elements for the Stokes problem. Mathematics of Computation, 52(186):437–456, April 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chandler:1988:PIC

[CG88] G. A. Chandler and I. G. Graham. Product integration-collocation methods for noncompact integral operator equations. Mathematics of Computation, 50(181): 125–138, January 1988. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Calio:1986:CGK

[CGM86a] Franca Caliò, Walter Gautschi, and Elena Marchetti. On computing Gauss– Kronrod quadrature formulae. Mathematics of Computation, 47(176):639– 650, S57–S63, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Calio:1986:SCG

[CGM86b] Franca Calio, Walter Gautschi, and Elena Marchetti. Supplement to on computing Gauss–Kronrod quadrature formulae. Mathematics of Computation, 47 (176):S57–S63, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cai:1989:ENS

[CGS89] Wei Cai, David Gottlieb, and Chi-Wang Shu. Essentially nonoscillatory spectral Fourier methods for shock wave calculations. Mathematics of Computation, 52(186):389–410, April 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Conn:1988:TCM

[CGT88] Andrew R. Conn, Nicholas I. M. Gould, and Philippe L. Toint. Testing a class of methods for solving minimization problems with simple bounds on the variables. Mathematics of Computation, 50(182):399–430, April 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chen:1984:ACH

[CH84] Han Lin Chen and Tron Hvaring. Approximation of complex harmonic functions by complex harmonic splines. Mathematics of Computation, 42(165): 151–164, January 1984. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chui:1989:DBS

[CH89] Charles K. Chui and Tian Xiao He. On the dimension of bivariate superspline spaces. Mathematics of Computation, 53(187):219–234, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Char:1980:SCF

[Cha80] Bruce W. Char. On Stieltjes’ continued fraction for the gamma function. Mathematics of Computation, 34(150): 547–551, April 1980. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chambers:1982:UBF

[Cha82a] Ll. G. Chambers. An upper bound for the first zero of Bessel functions. Mathematics of Computation, 38(158):589– 591, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chang:1982:CEM

[Cha82b] S. H. Chang. On certain extrapolation methods for the numerical solution of integro-differential equations. Mathematics of Computation, 39(159):165– 171, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chan:1984:ECF

[Cha84] Tony F. Chan. On the existence and computation of LU-factorizations with small pivots. Mathematics of Computation, 42(166):535–547, April 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). Citation communicated by Per Christian Hansen.

Chan:1985:CEC

[Cha85] Tony F. Chan. Corrigendum: “On the existence and computation of LUfactorizations with small pivots” [Math. Comp. 42 (1984), no. 166, 535–547, MR 85i:65037]. Mathematics of Computation, 44(169):282, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chang:1987:NSS

[Cha87] Chien-Cheng Chang. Numerical solution of stochastic differential equations with constant diffusion coefficients. Mathematics of Computation, 49(180):523– 542, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chen:1982:AEE

[Che82] T. H. Charles Chen. Asymptotic error estimates for Gaussian quadrature formulas. Mathematics of Computation, 38 (157):143–152, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chen:1983:TEM

[Che83] T. Charles Chen. Table errata: Methods of numerical integration [Academic Press, New York, 1975, MR 56 #7119] by P. J. Davis and P. Rabinowitz. Mathematics of Computation, 40(161):417, January 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chow:1982:UBA

[Cho82] Jeff Chow. On the uniqueness of best L_{2}[0, 1] approximation by piecewise polynomials with variable breakpoints. Mathematics of Computation, 39(160): 571–585, October 1982. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chui:1987:CIM

[CJW87] C. K. Chui, K. Jetter, and J. D. Ward.

Cardinal interpolation by multivariate splines. Mathematics of Computation, 48(178):711–724, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohn:1983:EFG

[CL83] H. Cohn and J. C. Lagarias. On the existence of fields governing the 2-invariants of the classgroup of Q(\sqrt{dp}) as p varies. Mathematics of Computation, 41(164): 711–730, October 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohen:1984:PTJ

[CL84a] H. Cohen and H. W. Lenstra, Jr. Primality testing and Jacobi sums. Mathematics of Computation, 42(165):297– 330, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Crandall:1984:TAS

[CL84b] M. G. Crandall and P. L. Lions. Two approximations of solutions of Hamilton– Jacobi equations. Mathematics of Computation, 43(167):1–19, July 1984. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohen:1987:INP

[CL87a] H. Cohen and A. K. Lenstra. Implementation of a new primality test. Mathematics of Computation, 48(177):103– 121, S1–S4, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohen:1987:SIN

[CL87b] H. Cohen and A. K. Lenstra. Supplement to implementation of a new primality test. Mathematics of Computation, 48(177):s1–s4, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Crandall:1980:MDA

[CM80] Michael G. Crandall and Andrew Majda. Monotone difference approximations for scalar conservation laws. Mathematics of Computation, 34(149):1–21, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Coleman:1983:CEB

[CM83] J. P. Coleman and A. J. Monaghan. Chebyshev expansions for the Bessel function J_{n}(z) in the complex plane. Mathematics of Computation, 40(161): 343–366, January 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohen:1987:CGN

[CM87a] H. Cohen and J. Martinet. Class groups of number fields: Numerical heuristics. Mathematics of Computation, 48(177): 123–137, January 1987. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Criscuolo:1987:CIP

[CM87b] Giuliana Criscuolo and Giuseppe Mastroianni. On the convergence of an interpolatory product rule for evaluating Cauchy principal value integrals. Mathematics of Computation, 48(178):725– 735, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chen:1983:SEE

[CMSY83] Goong Chen, Wendell H. Mills, Jr., Shun Hua Sun, and David A. Yost. Sharp error estimates for a finite element-penalty approach to a class of regulator problems. Mathematics of Computation, 40(161):151–173, January 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cayco:1986:FET

[CN86] M. E. Cayco and R. A. Nicolaides. Finite element technique for optimal pressure recovery from stream function formulation of viscous flows. Mathematics of Computation, 46(174):371–377, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chiu:1988:CHO

[CN88] C. Chiu and R. A. Nicolaides. Convergence of a higher-order vortex method for two-dimensional Euler equations. Mathematics of Computation, 51(184): 507–534, October 1988. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohn:1982:EME

[Coh82] Harvey Cohn. An explicit modular equation in two variables and Hilbert’s twelfth problem. Mathematics of Computation, 38(157):227–236, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohen:1984:PAN

[Coh84] Graeme L. Cohen. Primitive α-abundant numbers. Mathematics of Computation, 43(167):263–270, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohn:1988:FTV

[Coh88] Harvey Cohn. Fricke’s two-valued modular equations. Mathematics of Computation, 51(184):787–807, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Colman:1982:NPI

[Col82] W. J. A. Colman. The number of partitions of the integer N into M nonzero positive integers. Mathematics of Computation, 39(159):213–224, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Comodi:1989:HHJ

[Com89] M. I. Comodi. The Hellan–Herrmann– Johnson method: some new error estimates and postprocessing. Mathematics of Computation, 52(185):17–29, January 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Corliss:1980:IOC

[Cor80] George F. Corliss. Integrating ODEs in the complex plane—pole vaulting. Mathematics of Computation, 35(152):1181– 1189, October 1980. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Costello:1982:DPI

[Cos82] Patrick J. Costello. Density problems involving p_{r}(n). Mathematics of Computation, 38(158):633–637, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Costantini:1986:MCS

[Cos86] Paolo Costantini. On monotone and convex spline interpolation. Mathematics of Computation, 46(173):203–214, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cottet:1987:CVC

[Cot87] G. H. Cottet. Convergence of a vortex in cell method for the two-dimensional Euler equations. Mathematics of Computation, 49(180):407–425, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Canuto:1982:ARO

[CQ82] C. Canuto and A. Quarteroni. Approximation results for orthogonal polynomials in Sobolev spaces. Mathematics of Computation, 38(157):67–86, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Crisci:1982:SCM

[CR82] M. R. Crisci and E. Russo. A-stability of a class of methods for the numerical integration of certain linear systems of ordinary differential equations. Mathematics of Computation, 38(158):431–435, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Crisci:1983:EOR

[CR83] M. R. Crisci and E. Russo. An extension of Ortiz’ recursive formulation of the tau method to certain linear systems of ordinary differential equations. Mathematics of Computation, 41(163):27–42, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cooper:1980:AMN

[CS80a] G. J. Cooper and A. Sayfy. Additive methods for the numerical solution of ordinary differential equations. Mathematics of Computation, 35(152):1159– 1172, October 1980. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cruz:1980:MPR

[CS80b] Andrés Cruz and Javier Sesma. Modulus and phase of the reduced logarithmic derivative of the cylindrical Bessel function. Mathematics of Computation, 35(152):1317–1324, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cruz:1982:ZHF

[CS82] Andrés Cruz and Javier Sesma. Zeros of the Hankel function of real order and of its derivative. Mathematics of Computation, 39(160):639–645, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cooper:1983:ARK

[CS83a] G. J. Cooper and A. Sayfy. Additive Runge–Kutta methods for stiff ordinary differential equations. Mathematics of Computation, 40(161):207–218, January 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cruz:1983:MPR

[CS83b] Andrés Cruz and Javier Sesma. Modulus and phase of the reduced logarithmic derivative of the Hankel function. Mathematics of Computation, 41(164): 597–605, October 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Coleman:1987:FPC

[CS87a] John P. Coleman and Russell A. Smith. The Faber polynomials for circular sectors. Mathematics of Computation, 49 (179):231–241, S1–S4, July 1987. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Costabel:1987:CCM

[CS87b] Martin Costabel and Ernst P. Stephan.

On the convergence of collocation methods for boundary integral equations on polygons. Mathematics of Computation, 49(180):461–478, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cusick:1987:TFP

[CS87c] T. W. Cusick and Lowell Schoenfeld. A table of fundamental pairs of units in totally real cubic fields. Mathematics of Computation, 48(177):147–158, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cockburn:1989:TRK

[CS89] Bernardo Cockburn and Chi-Wang Shu.

TVB Runge–Kutta local projection discontinuous Galerkin finite element method for conservation laws. II. General framework. Mathematics of Computation, 52(186):411–435, April 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Coleman:1987:SFP

[CsAS87] John P. Coleman and Russel [sic] A. Smith. Supplement to “The Faber Polynomials for Circular Sectors”. Mathematics of Computation, 49(179):S1– S4, July 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Creutzburg:1986:NTT

[CT86] R. Creutzburg and M. Tasche. Numbertheoretic transforms of prescribed length. Mathematics of Computation, 47 (176):693–701, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chen:1987:BEN

[CT87a] Goong Chen and Ying-Liang Tsai. The boundary element numerical method for two-dimensional linear quadratic elliptic problems. I. Neumann control. Mathematics of Computation, 49(180):479– 498, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Crouzeix:1987:SPF

[CT87b] M. Crouzeix and V. Thomée. The stability in L_{p} and W_{p}^{1} of the L_{2}- projection onto finite element function spaces. Mathematics of Computation, 48(178):521–532, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Crouzeix:1987:DTS

[CT87c] Michel Crouzeix and Vidar Thomée. On the discretization in time of semilinear parabolic equations with nonsmooth initial data. Mathematics of Computation, 49(180):359–377, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Creutzburg:1989:PDC

[CT89] R. Creutzburg and M. Tasche. Parameter determination for complex numbertheoretic transforms using cyclotomic polynomials. Mathematics of Computation, 52(185):189–200, January 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Crouzeix:1989:EES

[CTW89] M. Crouzeix, V. Thomée, and L. B. Wahlbin. Error estimates for spatially discrete approximations of semilinear parabolic equations with initial data of low regularity. Mathematics of Computation, 53(187):25–41, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cuyt:1982:NSH

[Cuy82] Annie A. M. Cuyt. Numerical stability of the Halley-iteration for the solution of a system of nonlinear equations. Mathematics of Computation, 38(157): 171–179, January 1982. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cormack:1980:SVL

[CW80] G. V. Cormack and H. C. Williams.

Some very large primes of the form k · 2^{m} + 1. Mathematics of Computation, 35(152):1419–1421, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chui:1983:SMS

[CW83] Charles K. Chui and Ren Hong Wang.

On smooth multivariate spline functions. Mathematics of Computation, 41(163): 131–142, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cherubini:1987:CAI

[CW87] Juergen M. Cherubini and Rolf V. Wallisser. On the computation of all imaginary quadratic fields of class number one. Mathematics of Computation, 49 (179):295–299, July 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cybenko:1983:GOT

[Cyb83] George Cybenko. A general orthogonalization technique with applications to time series analysis and signal processing. Mathematics of Computation, 40 (161):323–336, January 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cantor:1981:NAF

[CZ81] David G. Cantor and Hans Zassenhaus. A new algorithm for factoring polynomials over finite fields. Mathematics of Computation, 36(154):587–592, April 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Darafsheh:1986:CIC

[Dar86] M. R. Darafsheh. Computing the irreducible characters of the group GL_{6}(2). Mathematics of Computation, 46(173): 301–319, January 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Davenport:1982:TET

[Dav82] J. H. Davenport. Table errata: Tables of indefinite integrals [English translation, Dover, New York, 1961, MR 23 #A256] by G. Petit Bois. Mathematics of Computation, 38(157):335– 336, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

deBoor:1985:GPT

[dD85] C. de Boor and R. DeVore. A geometric proof of total positivity for spline interpolation. Mathematics of Computation, 45(172):497–504, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Delfour:1986:DPA

[DD86a] M. C. Delfour and F. Dubeau. Discontinuous polynomial approximations in the theory of one-step, hybrid and multistep methods for nonlinear ordinary differential equations. Mathematics of Computation, 47(175):169–189, S1– S8, July 1986. CODEN MCMPAF.

ISSN 0025-5718 (print), 1088-6842 (electronic).

Delfour:1986:SDP

[DD86b] M. C. Delfour and F. Dubeau. Supplement to discontinuous polynomial approximations in the theory of one-step, hybrid and multistep methods for nonlinear ordinary differential equations. Mathematics of Computation, 47(175): S1–S8, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

deWeger:1986:PPP

[de 86] B. M. M. de Weger. Products of prime powers in binary recurrence sequences. part II: The elliptic case, with an application to a mixed quadratic-exponential equation. Mathematics of Computation, 47(176):729–739, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

deGee:1987:LMM

[de 87] Maarten de Gee. Linear multistep methods for functional-differential equations. Mathematics of Computation, 48(178): 633–649, April 1987. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Deanin:1985:CCM

[Dea85] Alice A. Deanin. A counterexample to a conjecture of Mahler on best p-adic Diophantine approximation constants. Mathematics of Computation, 45(172): 621–632, October 1985. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dougherty:1989:NMC

[DEH89] Randall L. Dougherty, Alan S. Edelman, and James M. Hyman. Nonnegativity- , monotonicity-, or convexity-preserving cubic and quintic Hermite interpolation. Mathematics of Computation, 52(186): 471–494, April 1989. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Delahaye:1981:ASS

[Del81] J. P. Delahaye. Automatic selection of sequence transformations. Mathematics of Computation, 37(155):197–204, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

DelPrete:1989:QCV

[Del89] Vincenza Del Prete. Quadratic convergence of vortex methods. Mathematics of Computation, 52(186):457–470, April 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Demmel:1988:PNA

[Dem88] James W. Demmel. The probability that a numerical analysis problem is difficult. Mathematics of Computation, 50(182): 449–480, April 1988. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Duijvestijn:1981:NPC

[DF81] A. J. W. Duijvestijn and P. J. Federico. The number of polyhedral (3-connected planar) graphs. Mathematics of Computation, 37(156):523–532, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dyn:1983:NSE

[DF83] Nira Dyn and Warren E. Ferguson, Jr. The numerical solution of equality constrained quadratic programming problems. Mathematics of Computation, 41(163):165–170, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Delfour:1981:DGM

[DHT81] M. Delfour, W. Hager, and F. Trochu.

Discontinuous Galerkin methods for ordinary differential equations. Mathematics of Computation, 36(154):455– 473, April 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dickinson:1980:SLE

[Dic80] Bradley W. Dickinson. Solution of linear equations with rational Toeplitz matrices. Mathematics of Computation, 34 (149):227–233, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

DiDonato:1982:RRI

[DiD82] A. R. DiDonato. Recurrence relations for the indefinite integrals of the associated Legendre functions. Mathematics of Computation, 38(158):547–551, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dixon:1981:AFF

[Dix81] John D. Dixon. Asymptotically fast factorization of integers. Mathematics of Computation, 36(153):255–260, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Djaoua:1981:MCL

[Dja81] M. Djaoua. A method of calculation of lifting flows around two-dimensional corner-shaped bodies. Mathematics of Computation, 36(154):405–425, April 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dougalis:1985:SHO

[DK85] Vassilios A. Dougalis and Ohannes A.

Karakashian. On some high-order accurate fully discrete Galerkin methods for the Korteweg–de Vries equation. Mathematics of Computation, 45(172):329– 345, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Duijvestijn:1981:LOS

[DL81] A. J. W. Duijvestijn and P. Leeuw. Lowest order squared rectangles and squares with the largest element not on the boundary. Mathematics of Computation, 37(155):223–228, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Daehlen:1988:BIQ

[DL88a] Morten Dæhlen and Tom Lyche. Bivariate interpolation with quadratic box splines. Mathematics of Computation, 51(183):219–230, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dennis:1988:HAS

[DL88b] J. E. Dennis, Jr. and Guang Ye Li. A hybrid algorithm for solving sparse nonlinear systems of equations. Mathematics of Computation, 50(181):155–166, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Descloux:1981:ASC

[DLR81] Jean Descloux, Mitchell Luskin, and Jacques Rappaz. Approximation of the spectrum of closed operators: the determination of normal modes of a rotating basin. Mathematics of Computation, 36(153):137–154, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Day:1982:TCI

[DM82a] J. D. Day and D. N. P. Murthy. Two classes of internally S-stable generalized Runge–Kutta processes which remain consistent with an inaccurate Jacobian. Mathematics of Computation, 39(160): 491–509, October 1982. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dennis:1982:DSU

[DM82b] J. E. Dennis, Jr. and Earl S. Marwil.

Direct secant updates of matrix factorizations. Mathematics of Computation, 38(158):459–474, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dahmen:1983:LIM

[DM83] Wolfgang A. Dahmen and Charles A.

Micchelli. On the linear independence of multivariate B-splines. II. complete configurations. Mathematics of Computation, 41(163):143–163, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Degond:1989:WPMa

[DMG89a] P. Degond and S. Mas-Gallic. The weighted particle method for convectiondiffusion equations. part 1: The case of an isotropic viscosity. Mathematics of Computation, 53(188):485–507, October 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Degond:1989:WPMb

[DMG89b] P. Degond and S. Mas-Gallic. The weighted particle method for convectiondiffusion equations. part 2: The anisotropic case. Mathematics of Computation, 53(188):509–525, October 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Demko:1984:DRI

[DMS84] Stephen Demko, William F. Moss, and Philip W. Smith. Decay rates for inverses of band matrices. Mathematics of Computation, 43(168):491–499, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Duran:1988:SMN

[DNW88] R. Durán, R. H. Nochetto, and Jun Ping Wang. Sharp maximum norm error estimates for finite element approximations of the Stokes problem in 2-D. Mathematics of Computation, 51(184):491– 506, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Douglas:1983:NMM

[DR83] Jim Douglas, Jr. and Jean E. Roberts. Numerical methods for a model for compressible miscible displacement in porous media. Mathematics of Computation, 41(164):441–459, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Douglas:1985:GEM

[DR85] Jim Douglas, Jr. and Jean E. Roberts. Global estimates for mixed methods for second order elliptic equations. Mathematics of Computation, 44(169):39–52, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Descloux:1985:RCA

[DRS85] J. Descloux, J. Rappaz, and R. Scholz. On the rate of convergence for the approximation of nonlinear problems. Mathematics of Computation, 45(171): 51–64, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

deBoor:1980:CAE

[dS80a] Carl de Boor and Blair Swartz. Collocation approximation to eigenvalues of an ordinary differential equation: the principle of the thing. Mathematics of Computation, 35(151):679–694, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dupont:1980:PAF

[DS80b] Todd Dupont and Ridgway Scott. Polynomial approximation of functions in Sobolev spaces. Mathematics of Computation, 34(150):441–463, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

deBoor:1981:CAE

[dS81a] Carl de Boor and Blair Swartz. Collocation approximation to eigenvalues of an ordinary differential equation: numerical illustrations. Mathematics of Computation, 36(153):1–19, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

deBoor:1981:LPP

[dS81b] Carl de Boor and Blair Swartz. Local piecewise polynomial projection methods for an O.D.E. which give highorder convergence at knots. Mathematics of Computation, 36(153):21–33, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dubner:1989:NMP

[Dub89] H. Dubner. A new method for producing large Carmichael numbers. Mathematics of Computation, 53(187):411–414, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dudley:1987:SIC

[Dud87] R. M. Dudley. Some inequalities for continued fractions. Mathematics of Computation, 49(180):585–593, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dupont:1982:MME

[Dup82]  Todd Dupont.  Mesh modification

for evolution equations. Mathematics of Computation, 39(159):85–107, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Duran:1987:QOE

[Dur87] Ricardo G. Durán. Quasi-optimal estimates for finite element approximations using Orlicz norms. Mathematics of Computation, 49(179):17–23, July 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dutt:1989:SCS

[Dut89] P. Dutt. Stability and convergence of spectral methods for hyperbolic initialboundary value problems. Mathematics of Computation, 53(188):547–561, October 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dunham:1981:RCD

[DW81] C. B. Dunham and Jack Williams. Rate of convergence of discretization in Chebyshev approximation. Mathematics of Computation, 37(155):135–139, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Douglas:1983:ITD

[DW83] Jim Douglas, Jr. and Mary Fanett Wheeler. Implicit, time-dependent variable grid finite difference methods for the approximation of a linear waterflood. Mathematics of Computation, 40(161): 107–121, January 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

deHoog:1985:ARK

[dW85a] Frank de Hoog and Richard Weiss. The application of Runge–Kutta schemes to singular initial value problems. Mathematics of Computation, 44(169):93–103, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dueck:1985:CCN

[DW85b] G. Dueck and H. C. Williams. Computation of the class number and class group of a complex cubic field. Mathematics of Computation, 45(171):223–231, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Douglas:1989:ASF

[DW89] Jim Douglas, Jr. and Jun Ping Wang.

An absolutely stabilized finite element method for the Stokes problem. Mathematics of Computation, 52(186):495– 508, April 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dyn:1984:CHB

[Dyn84] Nira Dyn. Composite Hermite–Birkhoff quadrature formulas of Gaussian type. Mathematics of Computation, 43(168): 535–541, October 1984. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Easton:1986:EIM

[Eas86] David Easton. Effective irrationality measures for certain algebraic numbers. Mathematics of Computation, 46(174): 613–622, April 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Engquist:1987:SSC

[EG87] Björn Engquist and Bertil Gustafsson. Steady state computations for wave propagation problems. Mathematics of Computation, 49(179):39–64, July 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eggermont:1984:APQ

[Egg84] P. P. B. Eggermont. Approximation properties of quadrature methods for Volterra integral equations of the first kind. Mathematics of Computation, 43 (168):455–471, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eggermont:1989:UEE

[Egg89] P. P. B. Eggermont. Uniform error estimates of Galerkin methods for monotone Abel–Volterra integral equations on the half-line. Mathematics of Computation, 53(187):157–189, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Evans:1980:LSE

[EGL80] John W. Evans, William B. Gragg, and Randall J. LeVeque. On least squares exponential sum approximation with positive coefficients. Mathematics of Computation, 34(149):203–211, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eichenauer-Herrmann:1989:PLP

[EHGL89] Jürgen Eichenauer-Herrmann, Holger Grothe, and Jürgen Lehn. On the period length of pseudorandom vector sequences generated by matrix generators. Mathematics of Computation, 52(185): 145–148, January 1989. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eriksson:1988:AFE

[EJ88] Kenneth Eriksson and Claes Johnson.

An adaptive finite element method for linear elliptic problems. Mathematics of Computation, 50(182):361–383, April 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Egecioglu:1989:FAR

[EK89] Ömer Eğecioğlu and Cetin K. Koc. A fast algorithm for rational interpolation via orthogonal polynomials. Mathematics of Computation, 53(187):249– 264, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Erdos:1988:LBC

[EKS88] P. Erdős, P. Kiss, and A. Sárközy. A lower bound for the counting function of Lucas pseudoprimes. Mathematics of Computation, 51(183):315–323, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Evans:1989:DNS

[EL89] D. J. Evans and C. Li. Determination of the D^{1/2}-norm of the SOR iterative matrix for the unsymmetric case. Mathematics of Computation, 53(187): 203–218, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Elden:1982:TDBa

[Eld82a] Lars Eldén. Time discretization in the backward solution of parabolic equations. I. Mathematics of Computation, 39(159):53–68, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Elden:1982:TDBb

[Eld82b] Lars Eldén. Time discretization in the backward solution of parabolic equations. II. Mathematics of Computation, 39(159):69–84, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Elizalde:1986:AEF

[Eli86] E. Elizalde. An asymptotic expansion for the first derivative of the generalized Riemann zeta function. Mathematics of Computation, 47(175):347–350, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Elkies:1988:XXX

[Elk88] Noam D. Elkies. On A^{4} +B^{4} +C^{4} = D^{4}. Mathematics of Computation, 51(184): 825–835, October 1988. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ellacott:1983:CFS

[Ell83] S. W. Ellacott. Computation of Faber series with application to numerical polynomial approximation in the complex plane. Mathematics of Computation, 40(162):575–587, April 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Elman:1986:SAI

[Elm86] Howard C. Elman. A stability analysis of incomplete LU factorizations. Mathematics of Computation, 47(175):191– 217, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Engquist:1989:NFE

[ELS89] Björn Engquist, Per Lötstedt, and Björn Sjögreen. Nonlinear filters for efficient shock computation. Mathematics of Computation, 52(186):509–537, April 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eichenauer:1988:NCP

[ELT88] Jürgen Eichenauer, Jürgen Lehn, and Alev Topuzoǧlu. A nonlinear congruential pseudorandom number generator with power of two modulus. Mathematics of Computation, 51(184):757– 759, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ernvall:1987:MCI

[EM87] R. Ernvall and T. Metsänkylä. A method for computing the Iwasawa λinvariant. Mathematics of Computation, 49(179):281–294, July 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ennola:1985:RCS

[EMT85] V. Ennola, S. Mäki, and R. Turunen. On real cyclic sextic fields. Mathematics of Computation, 45(172):591–611, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Engl:1985:IVM

[EN85] Heinz W. Engl and Andreas Neubauer.

An improved version of Marti’s method for solving ill-posed linear integral equations. Mathematics of Computation, 45 (172):405–416, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eriksson:1987:CAN

[EN87] Kenneth Eriksson and Yi Yong Nie.

Convergence analysis for a nonsymmetric Galerkin method for a class of singular boundary value problems in one space dimension. Mathematics of Computation, 49(179):167–186, July 1987.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Engquist:1980:SES

[EO80] Björn Engquist and Stanley Osher. Stable and entropy satisfying approximations for transonic flow calculations. Mathematics of Computation, 34(149): 45–75, January 1980. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Engquist:1981:OSD

[EO81] Björn Engquist and Stanley Osher. Onesided difference approximations for nonlinear conservation laws. Mathematics of Computation, 36(154):321–351, April 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Erdos:1986:NFW

[EP86] Paul Erdős and Carl Pomerance. On the number of false witnesses for a composite number. Mathematics of Computation, 46(173):259–279, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ericsson:1980:STL

[ER80] Thomas Ericsson and Axel Ruhe. The spectral transformation Lánczos method for the numerical solution of large sparse generalized symmetric eigenvalue problems. Mathematics of Computation, 35(152):1251–1268, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eriksson:1985:FEM

[Eri85a] Kenneth Eriksson. Finite element methods of optimal order for problems with singular data. Mathematics of Computation, 44(170):345–360, April 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eriksson:1985:HOL

[Eri85b] Kenneth Eriksson. High-order local rate of convergence by mesh-refinement in the finite element method. Mathematics of Computation, 45(171):109–142, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eriksson:1985:IAA

[Eri85c] Kenneth Eriksson. Improved accuracy by adapted mesh-refinements in the finite element method. Mathematics of Computation, 44(170):321–343, April 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eriksson:1984:GMS

[ET84] Kenneth Eriksson and Vidar Thomée.

Galerkin methods for singular boundary value problems in one space dimension. Mathematics of Computation, 42(166): 345–367, April 1984. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ennola:1985:CCF

[ET85a] Veikko Ennola and Reino Turunen. On cyclic cubic fields. Mathematics of Computation, 45(172):585–589, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ennola:1985:TRC

[ET85b] Veikko Ennola and Reino Turunen. On totally real cubic fields. Mathematics of Computation, 44(170):495–518, April 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Evans:1983:TFP

[Eva83] Ronald J. Evans. Twenty-fourth power residue difference sets. Mathematics of Computation, 40(162):677–683, April 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fairweather:1984:CCS

[Fai84] G. Fairweather. Corrigenda: “A computational study of finite element methods for second-order linear two-point boundary value problems” [Math. Comp. 40 (1983), no. 162, 499–518, MR 84c:65103] by Fairweather, P. Keast and J. C. Diaz. Mathematics of Computation, 43(167): 347, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Falk:1983:EEN

[Fal83] Richard S. Falk. Error estimates for the numerical identification of a variable coefficient. Mathematics of Computation, 40(162):537–546, April 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Farwig:1986:RCS

[Far86] Reinhard Farwig. Rate of convergence of Shepard’s global interpolation formula. Mathematics of Computation, 46(174): 577–590, April 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fatunla:1980:NIS

[Fat80] Simeon Ola Fatunla. Numerical integrators for stiff and highly oscillatory differential equations. Mathematics of Computation, 34(150):373–390, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fernandez-Cara:1988:CPA

[FCM88] Enrique Fernández-Cara and Carlos Moreno. Critical point approximation through exact regularization. Mathematics of Computation, 50(181):139– 153, January 1988. CODEN MCMPAF.

ISSN 0025-5718 (print), 1088-6842 (electronic).

Ferguson:1980:CJP

[Fer80] Warren E. Ferguson, Jr. The construction of Jacobi and periodic Jacobi matrices with prescribed spectra. Mathematics of Computation, 35(152):1203– 1220, October 1980. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1980:STI

[Fet80] Henry E. Fettis. On some trigonometric integrals. Mathematics of Computation, 35(152):1325–1329, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1981:CST

[Fet81a] Henry E. Fettis. Corrigendum: “On some trigonometric integrals”. Mathematics of Computation, 37(156):605, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1981:TETb

[Fet81b] Henry E. Fettis. Table errata: A table of the complete elliptic integral of the first kind for complex values of the modulus, Part I [Rep. No. ARL 69-0172, Aerospace Res. Lab., Wright–Patterson Air Force Base, Ohio, 1969; MR 40 #6725] by Fettis and J. C. Caslin. Mathematics of Computation, 36(153): 318, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1981:TEHb

[Fet81c] Henry E. Fettis. Table errata: Handbook of elliptic integrals for engineers and physicists [second edition, Springer, New York, 1971 and MR 43 #3506] by P. F. Byrd and M. D. Friedman. Mathematics of Computation, 36(153): 317, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1981:TEHa

[Fet81d] Henry E. Fettis. Table errata: Higher transcendental functions, Vol. II [McGraw-Hill, New York, 1953, F. Oberhettinger MR 15, 419] by A. Erdélyi, W. Magnus and F. G. Tricomi. Mathematics of Computation, 36(153):315, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1981:TETa

[Fet81e] Henry E. Fettis. Table errata: Table of integrals, series, and products [English translation of the fourth Russian edition, Academic Press, New York, 1965; MR 33 #5952] by I. S. Gradshteyn [I. S. Gradšteĭn] and I. M. Ryzhik [I. M. Ryžik]. Mathematics of Computation, 36(153):317–318, 320, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1982:CETb

[Fet82a] Henry E. Fettis. Corrigenda: “Erratum: Table of integrals, series, and products [Academic Press, New York, 1965 and MR 33 #5952] by I. S. Gradshteĭn and I. M. Ryzhik” [Math. Comp. 27 (1973), 451–452, MR 54 #4049] by A. Ojo and J. Sadiku. Mathematics of Computation, 38(158):657, April 1982. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1982:CETa

[Fet82b] Henry E. Fettis. Corrigendum: “Table errata: Table of integrals, series, and products by I. S. Gradšteĭn and I. M. Ryžik” [Math. Comp. 36 (1981), no.

153, 317–318], MR 82a:65009. Mathematics of Computation, 38(157):337, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1984:MTI

[Fet84a] Henry E. Fettis. More trigonometric integrals. Mathematics of Computation, 43 (168):557–564, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1984:TEF

[Fet84b] Henry E. Fettis. Table errata: Formulas and theorems for the special functions of mathematical physics [second edition, Springer, Berlin, 1948 and 38] by W. Magnus MR 10 and F. Oberhettinger. (German). Mathematics of Computation, 43(167):346, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1985:FEL

[Fet85] Henry E. Fettis. Further extensions of a Legendre function integral. Mathematics of Computation, 45(172):549– 552, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Funaro:1988:NMI

[FG88] D. Funaro and D. Gottlieb. A new method of imposing boundary conditions in pseudospectral approximations of hyperbolic equations. Mathematics of Computation, 51(184):599–613, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fredricksen:1986:GST

[FHS86] Harold M. Fredricksen, Alfred W. Hales, and Melvin M. Sweet. A generalization of Swan’s theorem. Mathematics of Computation, 46(173):321–331, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Field:1981:ADI

[Fie81] David A. Field. An algorithm for determining invertible quadratic isoparametric finite element transformations. Mathematics of Computation, 37(156):347– 360, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fishelov:1988:SSC

[Fis88] Dalia Fishelov. The spectrum and the stability of the Chebyshev collocation operator for transonic flow. Mathematics of Computation, 51(184):559– 579, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fletcher:1985:SAU

[FM85] R. Fletcher and S. P. J. Matthews. A stable algorithm for updating triangular factors under a rank one change. Mathematics of Computation, 45(172):471– 485, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Falk:1986:LCD

[FM86] R. S. Falk and P. B. Monk. Logarithmic convexity for discrete harmonic functions and the approximation of the Cauchy problem for Poisson’s equation. Mathematics of Computation, 47(175): 135–149, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ferguson:1984:ZMT

[FMPT84] H. R. P. Ferguson, R. D. Major, K. E. Powell, and H. G. Throolin. On zeros of Mellin transforms of SL_{2}(Z) cusp forms.

Mathematics of Computation, 42(165): 241–255, January 1984. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Freilich:1982:NSS

[FO82] J. H. Freilich and E. L. Ortiz. Numerical solution of systems of ordinary differential equations with the Tau method: an error analysis. Mathematics of Computation, 39(160):467–479, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Forster:1986:WCT

[FO86a] Klaus-Jürgen Förster and Georg-Peter Ostermeyer. On weighted Chebyshevtype quadrature formulas. Mathematics of Computation, 46(174):591–599, S21– S27, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Forster:1986:SWC

[FO86b] Klaus-Jurgen Forster and Georg-Peter Ostermeyer. Supplement to “On Weighted Chebyshev-Type Quadrature Formulas”. Mathematics of Computation, 46(174):S21–S27, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fornberg:1981:VIF

[For81] Bengt Fornberg. A vector implementation of the fast Fourier transform algorithm. Mathematics of Computation, 36 (153):189–191, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Forster:1987:WFA

[För87] Klaus-Jürgen Förster. On weight functions admitting Chebyshev quadrature. Mathematics of Computation, 49(179): 251–258, July 1987. CODEN MCMPAF.

ISSN 0025-5718 (print), 1088-6842 (electronic).

Fornberg:1988:GFD

[For88] Bengt Fornberg. Generation of finite difference formulas on arbitrarily spaced grids. Mathematics of Computation, 51 (184):699–706, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Foster:1984:AHM

[FP84a] D. M. E. Foster and G. M. Phillips. The arithmetic-harmonic mean. Mathematics of Computation, 42(165):183– 191, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Friedman:1984:SPM

[FP84b] Mark J. Friedman and Joseph E. Pasciak. Spectral properties for the magnetization integral operator. Mathematics of Computation, 43(168):447– 453, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fincke:1985:IMC

[FP85] U. Fincke and M. Pohst. Improved methods for calculating vectors of short length in a lattice, including a complexity analysis. Mathematics of Computation, 44(170):463–471, April 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fischer:1989:NIF

[FR89] Bernd Fischer and Lothar Reichel. Newton interpolation in Fejér and Chebyshev points. Mathematics of Computation, 53(187):265–278, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fransen:1981:ACH

[Fra81a] Arne Fransén. Addendum and corrigendum to: “High-precision values of the gamma function and of some related coefficients” [Math. Comp. 34 (1980), no. 150, 553–566, MR 81f:65004] by Fransén and S. Wrigge. Mathematics of Computation, 37(155):233–235, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [FW80].

Fransen:1981:CTS

[Fra81b] Arne Fransén. Conjectures on the Taylor series expansion coefficients of the Jacobian elliptic function sn(x, k). Mathematics of Computation, 37(156):475– 494, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Franke:1982:SDI

[Fra82] Richard Franke. Scattered data interpolation: Tests of some methods. Mathematics of Computation, 38(157):181– 200, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Friedman:1984:FEF

[Fri84] Mark J. Friedman. Finite element formulation of the general magnetostatic problem in the space of solenoidal vector functions. Mathematics of Computation, 43(168):415–431, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Foias:1984:DSN

[FT84] Ciprian Foiaş and Roger Temam. Determination of the solutions of the Navier– Stokes equations by a set of nodal values. Mathematics of Computation, 43(167): 117–133, July 1984. CODEN MCMPAF.

ISSN 0025-5718 (print), 1088-6842 (electronic).

Fransen:1980:HPV

[FW80] Arne Fransén and Staffan Wrigge. Highprecision values of the gamma function and of some related coefficients. Mathematics of Computation, 34(150):553– 566, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See addendum and corrigendum [Fra81a].

Fransen:1984:CMM

[FW84] Arne Fransén and Staffan Wrigge. Calculation of the moments and the moment generating function for the reciprocal gamma distribution. Mathematics of Computation, 42(166):601–616, April 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Frenzen:1985:NAE

[FW85] C. L. Frenzen and R. Wong. A note on asymptotic evaluation of some Hankel transforms. Mathematics of Computation, 45(172):537–548, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gunji:1981:PFF

[GA81] Hiroshi Gunji and Dennis Arnon. On polynomial factorization over finite fields. Mathematics of Computation, 36 (153):281–287, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gaal:1988:RIN

[Gaá88] István Gaál. On the resolution of inhomogeneous norm form equations in two dominating variables. Mathematics of Computation, 51(183):359–373, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Galeone:1983:UPM

[Gal83] Luciano Galeone. The use of positive matrices for the analysis of the large time behavior of the numerical solution of reaction-diffusion systems. Mathematics of Computation, 41(164):461– 472, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gartland:1984:AAE

[Gar84] Eugene C. Gartland, Jr. Accurate approximation of eigenvalues and zeros of selected eigenfunctions of regular Sturm–Liouville problems. Mathematics of Computation, 42(166):427–439, April 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gartland:1987:SUH

[Gar87a] Eugene C. Gartland, Jr. Supplement to uniform high-order difference schemes for a singularly perturbed two-point boundary value problem. Mathematics of Computation, 48(178):S5–S9, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gartland:1987:UHO

[Gar87b] Eugene C. Gartland, Jr. Uniform high-order difference schemes for a singularly perturbed two-point boundary value problem. Mathematics of Computation, 48(178):551–564, S5–S9, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Garnett:1988:CAD

[Gar88a] Lucy Garnett. A computer algorithm for determining the Hausdorff dimension of certain fractals. Mathematics of Computation, 51(183):291–300, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gartland:1988:AUC

[Gar88b] Eugene C. Gartland, Jr. An analysis of a uniformly convergent finite difference/finite element scheme for a model singular-perturbation problem. Mathematics of Computation, 51(183):93– 106, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gartland:1988:GMD

[Gar88c] Eugene C. Gartland, Jr. Graded-mesh difference schemes for singularly perturbed two-point boundary value problems. Mathematics of Computation, 51 (184):631–657, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Garabedian:1989:DLH

[Gar89] P. R. Garabedian. Design of a large helical system for magnetic fusion experiments. Mathematics of Computation, 52(186):539–543, April 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gatteschi:1980:SOP

[Gat80] Luigi Gatteschi. On some orthogonal polynomial integrals. Mathematics of Computation, 35(152):1291–1298, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1981:MST

[Gau81] Walter Gautschi. Minimal solutions of three-term recurrence relations and orthogonal polynomials. Mathematics of Computation, 36(154):547–554, April 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1983:CBC

[Gau83] Walter Gautschi. On the convergence behavior of continued fractions with real elements. Mathematics of Computation, 40(161):337–342, January 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1994:MCH

[Gau94] Walter Gautschi, editor. Mathematics of computation, 1943–1993: a half-century of computational mathematics: Mathematics of Computation 50th Anniversary Symposium, August 9–13, 1993, Vancouver, British Columbia, volume 48 of Proceedings of Symposia in Applied Mathematics. American Mathematical Society, Providence, RI, USA, 1994. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Gekeler:1982:LMM

[Gek82] Eckart Gekeler. Linear multistep methods for stable differential equations y^{..} = Ay+B(t)y^{.}c(t). Mathematics of Computation, 39(160):481–490, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gerth:1983:AMF

[Ger83a] Frank Gerth, III. An application of matrices over finite fields to algebraic number theory. Mathematics of Computation, 41(163):229–234, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gerver:1983:FLN

[Ger83b] Joseph L. Gerver. Factoring large numbers with a quadratic sieve. Mathematics of Computation, 41(163):287–294, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gerver:1983:ISI

[Ger83c] Joseph L. Gerver. Irregular sets of integers generated by the greedy algorithm. Mathematics of Computation, 40(162): 667–676, April 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gerasoulis:1988:FAM

[Ger88] A. Gerasoulis. A fast algorithm for the multiplication of generalized Hilbert matrices with vectors. Mathematics of Computation, 50(181):179–188, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Geveci:1984:CGA

[Gev84] Tunc Geveci. On the convergence of Galerkin approximation schemes for second-order hyperbolic equations in energy and negative norms. Mathematics of Computation, 42(166):393–415, April 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Geveci:1989:CTD

[Gev89] T. Geveci. On the convergence of a time discretization scheme for the Navier–Stokes equations. Mathematics of Computation, 53(187):43–53, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gfrerer:1987:PPC

[Gfr87a] Helmut Gfrerer. An a posteriori parameter choice for ordinary and iterated Tikhonov regularization of ill-posed problems leading to optimal convergence rates. Mathematics of Computation, 49 (180):507–522, S5–S12, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gfrerer:1987:SPP

[Gfr87b] Helmut Gfrerer. Supplement to an A posteriori parameter choice for ordinary and iterated Tikhonov regularization of ill-posed problems leading to optimal convergence rates. Mathematics of Computation, 49(180):S5–S12, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gray:1988:NDF

[GfZ88] Henry L. Gray and Nien fan Zhang. On a new definition of the fractional difference. Mathematics of Computation, 50(182):513–529, April 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gordon:1987:HOF

[GGT87] Daniel Gordon, Douglas Grenier, and Audrey Terras. Hecke operators and the fundamental domain for SL(3, Z). Mathematics of Computation, 48(177):159– 178, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Girstmair:1987:IPT

[Gir87] Kurt Girstmair. On invariant polynomials and their application in field theory. Mathematics of Computation, 48(178): 781–797, April 1987. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Girault:1988:IFE

[Gir88] V. Girault. Incompressible finite element methods for Navier–Stokes equations with nonstandard boundary conditions in R^{3}. Mathematics of Computation, 51(183):55–74, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Golub:1989:EEI

[GK89] Gene H. Golub and Mark D. Kent. Estimates of eigenvalues for iterative methods. Mathematics of Computation, 53 (188):619–626, October 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gabutti:1982:AMP

[GL82] B. Gabutti and J. N. Lyness. An acceleration method for the power series of entire functions of order 1. Mathematics of Computation, 39(160):587–597, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goodman:1985:ASS

[GL85] Jonathan B. Goodman and Randall J. LeVeque. On the accuracy of stable schemes for 2D scalar conservation laws. Mathematics of Computation, 45(171): 15–21, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Glasser:1981:CBS

[Gla81] M. L. Glasser. A class of Bessel summations. Mathematics of Computation, 37 (156):499–501, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gladwin:1982:OIM

[Gla82] C. J. Gladwin. On optimal integration methods for Volterra integral equations of the first kind. Mathematics of Computation, 39(160):511–518, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Glasser:1983:RPD

[Gla83] M. L. Glasser. A remarkable property of definite integrals. Mathematics of Computation, 40(162):561–563, April 1983.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gladwin:1987:ACO

[Gla87] C. J. Gladwin. An algorithm for the construction of optimal methods for the numerical solution of Volterra integral equations of the first kind. Mathematics of Computation, 48(178):625–632, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Guy:1987:PG

[GLS87] R. K. Guy, C. B. Lacampagne, and J. L.

Selfridge. Primes at a glance. Mathematics of Computation, 48(177):183– 202, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gostin:1982:SNF

[GM82] Gary B. Gostin and Philip B. McLaughlin, Jr. Six new factors of Fermat numbers. Mathematics of Computation, 38 (158):645–649, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Grone:1984:ASI

[GM84] Robert Grone and Russell Merris. An algorithm for the second immanant. Mathematics of Computation, 43(168):589– 591, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1985:GQI

[GM85a] Walter Gautschi and Gradimir V. Milovanović. Gaussian quadrature involving Einstein and Fermi functions with an application to summation of series. Mathematics of Computation, 44(169): 177–190, January 1985. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1985:SGQ

[GM85b] Walter Gautschi and Gradimir V. Milovanovic. Supplement to Gaussian quadrature involving Einstein and Fermi functions with an application to summation of series. Mathematics of Computation, 44(169):S1–S11, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1988:ASG

[GN88] Walter Gautschi and Sotirios E. Notaris. An algebraic study of Gauss–Kronrod quadrature formulae for Jacobi weight functions. Mathematics of Computation, 51(183):231–248, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Grothkopf:1982:CCP

[GO82] U. Grothkopf and G. Opfer. Complex Chebyshev polynomials on circular sectors with degree six or less. Mathematics of Computation, 39(160):599– 615, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Godwin:1984:CQF

[God84a] H. J. Godwin. Corrigenda: “On quartic fields of signature one with small discriminant. II”. Mathematics of Computation, 43(168):621, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Godwin:1984:QFS

[God84b] H. J. Godwin. On quartic fields of signature one with small discriminant. II. Mathematics of Computation, 42(166): 707–711, April 1984. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goetgheluck:1988:PDB

[Goe88] Pierre Goetgheluck. On prime divisors of binomial coefficients. Mathematics of Computation, 51(183):325–329, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goldstein:1980:VCE

[Gol80] Charles I. Goldstein. Variational crimes and L^{∞} error estimates in the finite element method. Mathematics of Computation, 35(152):1131–1157, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goldstein:1981:FEM

[Gol81] C. I. Goldstein. The finite element method with nonuniform mesh sizes for unbounded domains. Mathematics of Computation, 36(154):387–404, April 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goldstein:1982:FEM

[Gol82] Charles I. Goldstein. A finite element method for solving Helmholtz type equations in waveguides and other unbounded domains. Mathematics of Computation, 39(160):309–324, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gordon:1987:PME

[Gor87] Daniel M. Gordon. Perfect multiple error-correcting arithmetic codes. Mathematics of Computation, 49(180):621– 633, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gordon:1989:NEP

[Gor89] Daniel M. Gordon. On the number of elliptic pseudoprimes. Mathematics of Computation, 52(185):231–245, January 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gostin:1980:F

[Gos80] Gary B. Gostin. A factor of F_{17}. Mathematics of Computation, 35(151):975– 976, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gottlieb:1981:SPC

[Got81a] David Gottlieb. The stability of pseudospectral-Chebyshev methods. Mathematics of Computation, 36(153): 107–118, January 1981. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gottlieb:1981:SPF

[GOT81b] David Gottlieb, Steven A. Orszag, and Eli Turkel. Stability of pseudospectral and finite-difference methods for variable coefficient problems. Mathematics of Computation, 37(156):293– 305, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Glashoff:1981:NMC

[GR81] K. Glashoff and K. Roleff. A new method for Chebyshev approximation of complex-valued functions. Mathematics of Computation, 36(153):233– 239, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1988:FGK

[GR88] Walter Gautschi and Theodore J. Rivlin. A family of Gauss–Kronrod quadrature formulae. Mathematics of Computation, 51(184):749–754, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gelman:1989:CFI

[GR89a] Alex Gelman and Werner C. Rheinboldt.

On the computation of finite invariant sets of mappings. Mathematics of Computation, 52(186):545–551, April 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Grossmann:1989:UEH

[GR89b] Ch. Großmann and H.-G. Roos. Uniform enclosure of high order for boundary value problems by monotone discretization. Mathematics of Computation, 53(188):609–617, October 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Graham:1982:GMS

[Gra82] Ivan G. Graham. Galerkin methods for second kind integral equations with singularities. Mathematics of Computation, 39(160):519–533, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Grayson:1983:OSK

[Gra83] Matthew A. Grayson. The orbit space of a Kleinian group: Riley’s modest example. Mathematics of Computation, 40(162):633–646, April 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gramain:1987:CCA

[Gra87a] François Gramain. Corrigendum:

“Computing an arithmetic constant related to the ring of Gaussian integers” [Math. Comp. 44 (1985), no. 169, 241– 250, MR 87a:11028] by Gramain and M. Weber. Mathematics of Computation, 48 (178):854, April 1987. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gras:1987:SUR

[Gra87b] Marie-Nicole Gras. Special units in real cyclic sextic fields. Mathematics of Computation, 48(177):179–182, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Grandine:1988:SEM

[Gra88] Thomas A. Grandine. The stable evaluation of multivariate simplex splines. Mathematics of Computation, 50(181): 197–205, January 1988. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Greaves:1982:ASC

[Gre82] G. Greaves. An algorithm for the solution of certain differential-difference equations of advanced type. Mathematics of Computation, 38(157):237– 247, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Greengard:1986:CVF

[Gre86] Claude Greengard. Convergence of the vortex filament method. Mathematics of Computation, 47(176):387–398, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Guy:1980:CWD

[GS80] Richard K. Guy and J. L. Selfridge. Corrigendum to: “What drives an aliquot sequence?” [Math. Comp. 29 (1975), 101–107, MR 52 #5542]. Mathematics of Computation, 34(149):319–321, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gaal:1989:CAP

[GS89] I. Gaál and N. Schulte. Computing all power integral bases of cubic fields.

Mathematics of Computation, 53(188): 689–696, October 1989. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goldberg:1981:SIS

[GT81] Moshe Goldberg and Eitan Tadmor.

Scheme-independent stability criteria for difference approximations of hyperbolic initial-boundary value problems. II. Mathematics of Computation, 36 (154):603–626, April 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goldfarb:1984:OEJ

[GT84] D. Goldfarb and Ph. L. Toint. Optimal estimation of Jacobian and Hessian matrices that arise in finite difference calculations. Mathematics of Computation, 43 (167):69–88, July 1984. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goldberg:1985:CSC

[GT85] Moshe Goldberg and Eitan Tadmor.

Convenient stability criteria for difference approximations of hyperbolic initial-boundary value problems. Mathematics of Computation, 44(170):361– 377, April 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goldberg:1987:CSC

[GT87] Moshe Goldberg and Eitan Tadmor.

Convenient stability criteria for difference approximations of hyperbolic initial-boundary value problems. II. Mathematics of Computation, 48(178): 503–520, April 1987. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Guirguis:1987:TOB

[Gui87] Georges H. Guirguis. A third-order boundary condition for the exterior Stokes problem in three dimensions. Mathematics of Computation, 49(180): 379–389, October 1987. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gupta:1984:DEP

[Gup84] Hansraj Gupta. Diophantine equations in partitions. Mathematics of Computation, 42(165):225–229, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Groetsch:1987:ATF

[GV87] C. W. Groetsch and C. R. Vogel. Asymptotic theory of filtering for linear operator equations with discrete noisy data. Mathematics of Computation, 49(180): 499–506, October 1987. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1984:MYL

[GW84] Walter Gautschi and Jet Wimp. In memoriam: Yudell L. Luke, June 26, 1918–May 6, 1983. Mathematics of Computation, 43(168):349–352, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Graimain:1985:SCA

[GW85a] F. Graimain and M. Weber. Supplement to computing an arithmetic constant related to the ring of Gaussian integers. Mathematics of Computation, 44(169): S13–S16, January 1985. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gramain:1985:CAC

[GW85b] F. Gramain and M. Weber. Computing an arithmetic constant related to the ring of Gaussian integers. Mathematics of Computation, 44(169):241–250, S13– S16, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Haber:1983:PIP

[Hab83] Seymour Haber. Parameters for integrating periodic functions of several variables. Mathematics of Computation, 41(163):115–129, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hackbusch:1980:CMG

[Hac80] Wolfgang Hackbusch. Convergence of multi-grid iterations applied to difference equations. Mathematics of Computation, 34(150):425–440, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hadjidimos:1982:COA

[Had82] A. Hadjidimos. Corrigenda: “Optimum accelerated overrelaxation method in a special case” [Math. Comp. 36 (1981), 183–187, MR 82c:65016] by G. Avdelas and the author. Mathematics of Computation, 38(158):657, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hagis:1980:OPE

[Hag80] Peter Hagis, Jr. Outline of a proof that every odd perfect number has at least eight prime factors. Mathematics of Computation, 35(151):1027–1032, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hagis:1981:UHN

[Hag81] Peter Hagis, Jr. Unitary hyperperfect numbers. Mathematics of Computation, 36(153):299–301, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hagis:1983:SPO

[Hag83] Peter Hagis, Jr. Sketch of a proof that an odd perfect number relatively prime to 3 has at least eleven prime factors. Mathematics of Computation, 40(161): 399–404, January 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Halpern:1982:ABC

[Hal82] Laurence Halpern. Absorbing boundary conditions for the discretization schemes of the one-dimensional wave equation. Mathematics of Computation, 38(158): 415–429, April 1982. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Halpern:1986:ABC

[Hal86] Laurence Halpern. Artificial boundary conditions for the linear advection diffusion equation. Mathematics of Computation, 46(174):425–438, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Han:1982:FEM

[Han82] Houde Han. The finite element method in a family of improperly posed problems. Mathematics of Computation, 38 (157):55–65, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Harten:1986:LTS

[Har86] Ami Harten. On a large time-step high resolution scheme. Mathematics of Computation, 46(174):379–399, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hasegawa:1986:CGC

[Has86] Takemitsu Hasegawa. Corrigendum:

“Generalized Chebyshev interpolation and its application to automatic quadrature” [Math. Comp. 41 (1983), no. 164, 537–553, MR 84m:65037] by Hasegawa, T. Torii and I. Ninomiya. Mathematics of Computation, 47(175):385, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hsu:1981:AFT

[HC81] L. C. Hsu and Y. S. Chou. An asymptotic formula for a type of singular oscillatory integrals. Mathematics of Computation, 37(156):503–507, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Heald:1985:RAF

[Hea85] Mark A. Heald. Rational approximations for the Fresnel integrals. Mathematics of Computation, 44(170):459– 461, April 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Heald:1986:CRA

[Hea86] Mark A. Heald. Corrigenda: “Rational approximations for the Fresnel integrals” [Math. Comp. 44 (1985), no. 170, 459–461, MR 86b:65017]. Mathematics of Computation, 46(174):771, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Heinrichs:1989:ICN

[Hei89] Wilhelm Heinrichs. Improved condition number for spectral methods. Mathematics of Computation, 53(187):103– 119, July 1989. CODEN MCMPAF.

ISSN 0025-5718 (print), 1088-6842 (electronic).

Hagstrom:1988:ABC

[HH88a] Thomas Hagstrom and S. I. Hariharan. Accurate boundary conditions for exterior problems in gas dynamics. Mathematics of Computation, 51(184):581– 597, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hughes-Hallett:1988:CCA

[HH88b] A. J. Hughes-Hallett. Corrigenda: “The convergence of accelerated overrelaxation iterations” [Math. Comp. 47 (1986), no. 175, 219–223, MR 87g:65046]. Mathematics of Computation, 51(183):387, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hardy:1987:CCN

[HHR^{+}87] Kenneth Hardy, R. H. Hudson, D. Richman, Kenneth S. Williams, and N. M. Holtz. Calculation of the class numbers of imaginary cyclic quartic fields. Mathematics of Computation, 49(180): 615–620, October 1987. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hicks:1981:HSS

[Hic81] D. L. Hicks. Hydrocode subcycling stability. Mathematics of Computation, 37 (155):69–78, July 1981. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Higdon:1986:ABC

[Hig86a] Robert L. Higdon. Absorbing boundary conditions for difference approximations to the multidimensional wave equation. Mathematics of Computation, 47(176):

437–459, October 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Higham:1986:NMM

[Hig86b] Nicholas J. Higham. Newton’s method for the matrix square root. Mathematics of Computation, 46(174):537–549, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Higdon:1987:NAB

[Hig87] Robert L. Higdon. Numerical absorbing boundary conditions for the wave equation. Mathematics of Computation, 49 (179):65–90, July 1987. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hilliker:1982:ASC

[Hil82] David Lee Hilliker. An algorithm for solving a certain class of Diophantine equations. I. Mathematics of Computation, 38(158):611–626, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hendy:1981:JPA

[HJ81] M. D. Hendy and N. S. Jeans. The Jacobi–Perron algorithm in integer form. Mathematics of Computation, 36(154): 565–574, April 1981. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hagstrom:1987:ABC

[HK87] T. M. Hagstrom and H. B. Keller.

Asymptotic boundary conditions and numerical methods for nonlinear elliptic problems on unbounded domains. Mathematics of Computation, 48(178):449– 470, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hale:1988:USA

[HLR88] Jack K. Hale, Xiao Biao Lin, and Geneviève Raugel. Upper semicontinuity of attractors for approximations of semigroups and partial differential equations. Mathematics of Computation, 50(181): 89–123, January 1988. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hoff:1985:LIF

[Hof85]  David Hoff.  A linearly implicit

finite-difference scheme for the onedimensional porous medium equation. Mathematics of Computation, 45(171): 23–33, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Horner:1980:RRC

[Hor80] T. S. Horner. Recurrence relations for the coefficients in Chebyshev series solutions of ordinary differential equations. Mathematics of Computation, 35(151): 893–905, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Howard:1980:SCB

[How80] F. T. Howard. A special class of Bell polynomials. Mathematics of Computation, 35(151):977–989, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hao:1984:GBN

[HP84] Fred H. Hao and Charles J. Parry. Generalized Bernoulli numbers and mregular primes. Mathematics of Computation, 43(167):273–288, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hemker:1981:MGM

[HS81] P. W. Hemker and H. Schippers. Multiple grid methods for the solution of Fredholm integral equations of the second kind. Mathematics of Computation, 36 (153):215–232, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hoff:1985:EBF

[HS85] David Hoff and Joel Smoller. Error bounds for finite-difference approximations for a class of nonlinear parabolic systems. Mathematics of Computation, 45(171):35–49, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Huang:1981:SCE

[HT81] Mingyou Huang and Vidar Thomée.

Some convergence estimates for semidiscrete type schemes for time-dependent nonselfadjoint parabolic equations. Mathematics of Computation, 37(156): 327–346, October 1981. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Harten:1981:FOA

[HTE81] Amiram Harten and Hillel Tal-Ezer. On a fourth order accurate implicit finite difference scheme for hyperbolic conservation laws. I. nonstiff strongly dynamic problems. Mathematics of Computation, 36(154):353–373, April 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hasegawa:1983:GCI

[HTN83] Takemitsu Hasegawa, Tatsuo Torii, and Ichizo Ninomiya. Generalized Chebyshev interpolation and its application to automatic quadrature. Mathematics of Computation, 41(164):537–553, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hudson:1984:PSP

[Hud84] Richard H. Hudson. Products and sums of powers of binomial coefficients mod p and solutions of certain quaternary Diophantine systems. Mathematics of Computation, 43(168):603– 613, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hudson:1985:AEI

[Hud85] Richard H. Hudson. Averaging effects on irregularities in the distribution of primes in arithmetic progressions. Mathematics of Computation, 44(170):561– 571, April 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

HughesHallett:1986:CAO

[Hug86] A. J. Hughes Hallett. The convergence of accelerated overrelaxation iterations. Mathematics of Computation, 47(175): 219–223, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hundsdorfer:1989:SCP

[HV89] W. H. Hundsdorfer and J. G. Verwer. Stability and convergence of the Peaceman–Rachford ADI method for initial-boundary value problems. Mathematics of Computation, 53(187):81– 101, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hudson:1982:CNF

[HW82] Richard H. Hudson and Kenneth S. Williams. Class number formulae of Dirichlet type. Mathematics of Computation, 39(160):725–732, October 1982.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Huang:1984:NAI

[HW84] Yougun Huang and Jack Williams. A note on an algorithm for interpolating rationals. Mathematics of Computation, 42(165):111–113, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Igarashi:1984:TCI

[Iga84] Masao Igarashi. A termination criterion for iterative methods used to find the zeros of polynomials. Mathematics of Computation, 42(165):165–171, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Isaacson:1989:CPC

[II89] David Isaacson and Eli L. Isaacson.

Comment on A.-P. Calderón’s paper: “On an inverse boundary value problem” [in Seminar on Numerical Analysis and its Applications to Continuum Physics (Rio de Janeiro, 1980), 65–73, Soc. Brasil. Mat., Rio de Janeiro, 1980, MR 81k:35160]. Mathematics of Computation, 52(186):553–559, April 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Isaacson:1981:NAS

[IIMPL81] D. Isaacson, E. L. Isaacson, D. Marchesin, and P. J. Paes-Leme. Numerical analysis of spectral properties of coupled oscillator Schrödinger operators. I. Single and double well anharmonic oscillators. Mathematics of Computation, 37 (156):273–292, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Iwaniec:1988:AEM

[IL88] T. Iwaniec and A. Lutoborski. Asymptotic expansions of multiple integrals of rapidly oscillating functions. Mathematics of Computation, 50(181):215– 228, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Iserles:1983:FFR

[IN83] A. Iserles and S. P. Nørsett. Frequency fitting of rational approximations to the exponential function. Mathematics of Computation, 40(162):547– 559, April 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Iserles:1987:TSM

[IN87] A. Iserles and S. P. Nørsett. Two-step methods and bi-orthogonality. Mathematics of Computation, 49(180):543– 552, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ioakimidis:1983:FCR

[Ioa83] N. I. Ioakimidis. Further convergence results for the weighted Galerkin method of numerical solution of Cauchy-type singular integral equations. Mathematics of Computation, 41(163):79–85, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ioakimidis:1985:AQR

[Ioa85a] N. I. Ioakimidis. Application of quadrature rules for Cauchy-type integrals to the generalized Poincaré-Bertrand formula. Mathematics of Computation, 44 (169):199–206, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ioakimidis:1985:UCG

[Ioa85b] N. I. Ioakimidis. On the uniform convergence of Gaussian quadrature rules for Cauchy principal value integrals and their derivatives. Mathematics of Computation, 44(169):191–198, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ireland:1986:CNC

[IS86] K. F. Ireland and R. D. Small. Class numbers of cyclotomic function fields. Mathematics of Computation, 46(173): 337–340, January 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Isaacs:1980:ELF

[Isa80] Godfrey L. Isaacs. Exponential laws for fractional differences. Mathematics of Computation, 35(151):933–936, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Iserles:1981:QMS

[Ise81] A. Iserles. Quadrature methods for stiff ordinary differential systems. Mathematics of Computation, 36(153):171– 182, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Iserles:1982:CEA

[Ise82] Arieh Iserles. Composite exponential approximations. Mathematics of Computation, 38(157):99–112, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Joshi:1982:ICH

[JA82] C. M. Joshi and J. P. Arya. Inequalities for certain hypergeometric functions. Mathematics of Computation, 38 (157):201–205, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jaeschke:1983:SSA

[Jae83] G. Jaeschke. On the smallest k such that all k · 2^{n} + 1 are composite. Mathematics of Computation, 40(161):381– 384, January 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jaeschke:1985:CSS

[Jae85] G. Jaeschke. Corrigendum: “On the smallest k such that all k · 2^{n} + 1 are composite” [Math. Comp. 40 (1983), no. 161, 381–384, MR 84k:10006]. Mathematics of Computation, 45(172):637, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jager:1982:SCN

[Jag82] H. Jager. On the speed of convergence of the nearest integer continued fraction. Mathematics of Computation, 39(160): 555–558, October 1982. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

James:1980:GOO

[Jam80] Rodney James. The groups of order p^{6} (p an odd prime). Mathematics of Computation, 34(150):613–637, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jaccard:1982:ASO

[JE82] Yves Jaccard and Hugo Evéquoz. Approximation of the spectrum of an operator given by the magnetohydrodynamic stability of a plasma. Mathematics of Computation, 39(160):443–452, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jing:1987:ACC

[JF87] Zhongqi Jing and Adly T. Fam. An algorithm for computing continuous Chebyshev approximations. Mathematics of Computation, 48(178):691–710, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jia:1987:BPS

[Jia87] Rong Qing Jia. L^{∞}-boundedness of L^{2}- projections on splines for a multiple geometric mesh. Mathematics of Computation, 48(178):675–690, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnson:1987:EES

[JLTW87] Claes Johnson, Stig Larsson, Vidar Thomée, and Lars B. Wahlbin. Error estimates for spatially discrete approximations of semilinear parabolic equations with nonsmooth initial data. Mathematics of Computation, 49(180):331– 357, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jones:1985:VSS

[JM85] H. M. Jones and S. McKee. Variable step size predictor-corrector schemes for second kind Volterra integral equations. Mathematics of Computation, 44(170): 391–404, April 1985. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnson:1980:CBI

[JN80] Claes Johnson and J.-Claude Nédélec.

On the coupling of boundary integral and finite element methods. Mathematics of Computation, 35(152):1063– 1079, October 1980. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnson:1982:ASM

[JP82] Claes Johnson and Juhani Pitkäranta. Analysis of some mixed finite element methods related to reduced integration. Mathematics of Computation, 38(158): 375–400, April 1982. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnson:1986:ADG

[JP86] C. Johnson and J. Pitkäranta. An analysis of the discontinuous Galerkin method for a scalar hyperbolic equation. Mathematics of Computation, 46(173): 1–26, January 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jerome:1982:EEM

[JR82] Joseph W. Jerome and Michael E. Rose. Error estimates for the multidimensional two-phase Stefan problem. Mathematics of Computation, 39(160):377–414, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jen:1981:CSA

[JS81] Erica Jen and R. P. Srivastav. Cubic splines and approximate solution of singular integral equations. Mathematics of Computation, 37(156):417–423, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jeltsch:1985:ABS

[JS85] Rolf Jeltsch and Klaus-Günther Strack. Accuracy bounds for semidiscretizations of hyperbolic problems. Mathematics of Computation, 45(172):365–376, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnson:1986:SDM

[JS86] Claes Johnson and Jukka Saranen.

Streamline diffusion methods for the incompressible Euler and Navier–Stokes equations. Mathematics of Computation, 47(175):1–18, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnsen:1987:CBM

[JS87a] Ben Johnsen and Eldar Straume. Counting binary matrices with given row and column sums. Mathematics of Computation, 48(178):737–750, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnson:1987:CFE

[JS87b] Claes Johnson and Anders Szepessy.

On the convergence of a finite element method for a nonlinear hyperbolic conservation law. Mathematics of Computation, 49(180):427–444, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnson:1987:CSP

[JSW87] C. Johnson, A. H. Schatz, and L. B.

Wahlbin. Crosswind smear and pointwise errors in streamline diffusion finite element methods. Mathematics of Computation, 49(179):25–38, July 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jameson:1981:ISD

[JT81] A. Jameson and E. Turkel. Implicit schemes and LU decompositions. Mathematics of Computation, 37(156):385– 397, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kallman:1982:MFP

[Kal82] Ralph Kallman. A method for finding permanents of 0, 1 matrices. Mathematics of Computation, 38(157):167– 170, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Karlsen:1980:CSS

[Kar80] Lasse K. Karlsen. Computation of steady shocks by second-order finitedifference schemes. Mathematics of Computation, 34(150):391–400, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Karakashian:1986:RKM

[Kar86] Ohannes A. Karakashian. On Runge– Kutta methods for parabolic problems with time-dependent coefficients. Mathematics of Computation, 47(175):77– 101, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Katti:1980:FDS

[Kat80] C. P. Katti. Five-diagonal sixth order methods for two-point boundary value problems involving fourth order differential equations. Mathematics of Computation, 35(152):1177–1179, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kuchel:1987:BCF

[KBF87] Philip W. Kuchel, Brian T. Bulliman, and Edward D. Fackerell. Bi-cyclide and flat-ring cyclide coordinate surfaces: correction of two expressions. Mathematics of Computation, 49(180):607– 613, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kalla:1982:IJF

[KCL82] Shyam L. Kalla, Salvador Conde, and Yudell L. Luke. Integrals of Jacobi functions. Mathematics of Computation, 38 (157):207–214, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kautsky:1984:GQJ

[KE84] J. Kautský and S. Elhay. Gauss quadratures and Jacobi matrices for weight functions not of one sign. Mathematics of Computation, 43(168):543–550, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kearfott:1983:SAI

[Kea83] Ralph Baker Kearfott. A sinc approximation for the indefinite integral. Mathematics of Computation, 41(164):559– 572, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kearfott:1987:AGB

[Kea87] R. Baker Kearfott. Abstract generalized bisection and a cost bound. Mathematics of Computation, 49(179):187–202, July 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Keeling:1989:GRK

[Kee89] Stephen L. Keeling. Galerkin/Runge– Kutta discretizations for parabolic equations with time-dependent coefficients. Mathematics of Computation, 52(186): 561–586, April 1989. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Keller:1982:TEN

[Kel82a] Wilfrid Keller. Table errata: “New primes of the form k · 2^{n} + 1” [Math.

Comp. 33 (1979), no. 148, 1333–1336], MR 80h:10009, by R. Baillie. Mathematics of Computation, 38(157):335, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Keller:1982:TEP

[Kel82b] Wilfrid Keller. Table errata: “Properties of the sequence 3 · 2^{n} + 1” [Math. Comp. 30 (1976), no. 135, 657–663, MR 53 #7933] by S. W. Golomb. Mathematics of Computation, 38(157):335, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Keller:1982:TES

[Kel82c] Wilfrid Keller. Table errata: “Some very large primes of the form k · 2^{m} + 1” [Math. Comp. 35 (1980), no. 152, 1419–1421], MR 81i:10011 by G. V. Cormack and H. C. Williams. Mathematics of Computation, 38(157):335, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Keller:1983:CPF

[Kel83a] Wilfred Keller. Corrigendum: “Primes of the form n! ± 1 and 2 · 3 · 5 ⋯ p ± 1” [Math. Comp. 38 (1982), no. 158, 639– 643 and MR 83c:10006] by J. P. Buhler, R. E. Crandall and M. A. Penk. Mathematics of Computation, 40(162): 727, April 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Keller:1983:FFN

[Kel83b] Wilfrid Keller. Factors of Fermat numbers and large primes of the form k · 2^{n} + 1. Mathematics of Computation, 41 (164):661–673, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kelley:1986:SLA

[Kel86] C. T. Kelley. A Shamanskiĭ-like acceleration scheme for nonlinear equations at singular roots. Mathematics of Computation, 47(176):609–623, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kershaw:1983:SEW

[Ker83] D. Kershaw. Some extensions of W.

Gautschi’s inequalities for the gamma function. Mathematics of Computation, 41(164):607–611, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Keast:1983:CSF

[KFD83] P. Keast, G. Fairweather, and J. C.

Diaz. A computational study of finite element methods for second order linear two-point boundary value problems. Mathematics of Computation, 40(162): 499–518, April 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kovvali:1982:PFS

[KG82] S. Kovvali and G. K. Gupta. Polynomial formulation of second derivative multistep methods. Mathematics of Computation, 38(158):447–458, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kok:1985:CGA

[KG85a] Barbara Kok and Tunc Geveci. The convergence of Galerkin approximation schemes for second-order hyperbolic equations with dissipation. Mathematics of Computation, 44(170):379–390, S17– S25, April 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kok:1985:SCG

[KG85b] Barbara Kok and Tunc Geveci. Supplement to the convergence of Galerkin approximation schemes for second-order hyperbolic equations with dissipation. Mathematics of Computation, 44(170): S17–S25, April 1985. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kim:1988:AZR

[Kim88] Myong-Hi Kim. On approximate zeros and rootfinding algorithms for a complex polynomial. Mathematics of Computation, 51(184):707–719, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

King:1980:EOP

[Kin80] Richard F. King. An efficient onepoint extrapolation method for linear convergence. Mathematics of Computation, 35(152):1285–1290, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

King:1983:ABL

[Kin83] Richard F. King. Anderson–Björck for linear sequences. Mathematics of Computation, 41(164):591–596, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kiper:1984:FSC

[Kip84] Ayşe Kiper. Fourier series coefficients for powers of the Jacobian elliptic functions. Mathematics of Computation, 43(167): 247–259, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kishore:1981:OPQ

[Kis81]  Masao Kishore.  On odd perfect,

quasiperfect, and odd almost perfect numbers. Mathematics of Computation, 36(154):583–586, April 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kishore:1983:OPN

[Kis83] Masao Kishore. Odd perfect numbers not divisible by 3. II. Mathematics of Computation, 40(161):405–411, January 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kishore:1984:OTN

[Kis84] Masao Kishore. Odd triperfect numbers.

Mathematics of Computation, 42(165): 231–233, January 1984. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kishore:1985:OTN

[Kis85] Masao Kishore. Odd triperfect numbers are divisible by eleven distinct prime factors. Mathematics of Computation, 44 (169):261–263, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kiwiel:1985:ANC

[Kiw85] Krzysztof C. Kiwiel. An algorithm for nonsmooth convex minimization with errors. Mathematics of Computation, 45(171):173–180, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kheshgi:1985:AFE

[KL85] Haroon Kheshgi and Mitchell Luskin.

Analysis of the finite element variable penalty method for Stokes equations. Mathematics of Computation, 45(172): 347–363, October 1985. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kannan:1988:PFN

[KLL88] R. Kannan, A. K. Lenstra, and L. Lovász. Polynomial factorization and nonrandomness of bits of algebraic and some transcendental numbers. Mathematics of Computation, 50(181):235– 250, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kreiss:1986:SCS

[KMS^{+}86] H.-O. Kreiss, T. A. Manteuffel, B. Swartz, B. Wendroff, and A. B. White, Jr. Supra-convergent schemes on irregular grids. Mathematics of Computation, 47(176):537–554, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Knight:1988:TETb

[Kni88] J. H. Knight. Table errata: Tables of Laplace transforms [Springer, New York, 1973; MR 50 #5375] by F. Oberhettinger and L. Badii. Mathematics of Computation, 50(182):653–654, April 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Koblitz:1987:ECC

[Kob87] Neal Koblitz. Elliptic curve cryptosystems. Mathematics of Computation, 48 (177):203–209, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kolbig:1982:CE

[Köl82] K. S. Kölbig. Closed expressions for ∫_{0}^{1} t^{−1} log^{n−1} t log^{p}(1 − t)dt. Mathematics of Computation, 39(160):647–654, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kolbig:1983:Ia

[Köl83a] K. S. Kölbig. On the integral ∫_{0}^{π/2} log^{n} cos x log^{p} sin xdx. Mathematics of Computation, 40(162):565–570, April 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kolbig:1983:Ib

[Köl83b] K. S. Kölbig. On the integral ∫_{0}^{∞} e^{−μt}t^{ν−1} log^{m} tdt. Mathematics of Computation, 41(163):171–182, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kiss:1987:PR

[KP87] Péter Kiss and Bui Minh Phong. On a problem of A. rotkiewicz. Mathematics of Computation, 48(178):751–755, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Knight:1988:TETa

[KP88] J. H. Knight and J. R. Philip. Table errata: Tables of integral transforms. Vol. 1 [McGraw-Hill, New York, 1954; MR 15, 868] by A. Erdélyi, W. Magnus, F. Oberhettinger and F. G. Tricomi. Mathematics of Computation, 50(182): 653, April 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kim:1989:PRP

[KP89] Su Hee Kim and Carl Pomerance. The probability that a random probable prime is composite. Mathematics of Computation, 53(188):721–741, October 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kaltofen:1989:CGC

[KR89] Erich Kaltofen and Heinrich Rolletschek.

Computing greatest common divisors and factorizations in quadratic number fields. Mathematics of Computation, 53 (188):697–720, October 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kretschmer:1986:CEC

[Kre86] Thomas J. Kretschmer. Construction of elliptic curves with large rank. Mathematics of Computation, 46(174):627– 635, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Krogh:1980:CRR

[Kro80] Fred T. Krogh. Corrigendum: “Recurrence relations for computing with modified divided differences” [Math. Comp. 33 (1979), no. 148, 1265–1271, MR 80f:65094]. Mathematics of Computation, 35(152):1445, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Krupnikov:1983:TET

[Kru83] Ernst D. Krupnikov. Table errata: Table of integrals, series, and products [Academic Press, New York, 1980 and MR 81g:33001] by I. S. Gradshteĭn and I. M. Ryzhik. Mathematics of Computation, 41(164):780–783, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kumar:1987:NCT

[KS87] Sunil Kumar and Ian H. Sloan. A new collocation-type method for Hammerstein integral equations. Mathematics of Computation, 48(178):585–593, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kurtz:1986:FPT

[KSW86] G. C. Kurtz, Daniel Shanks, and H. C. Williams. Fast primality tests for numbers less than 50 · 10^{9}. Mathematics of Computation, 46(174):691–701, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kumar:1988:SCT

[Kum88] Sunil Kumar. Superconvergence of a collocation-type method for simple turning points of Hammerstein equations. Mathematics of Computation, 50(182): 385–398, April 1988. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lax:1982:EFI

[LA82] M. Lax and G. P. Agrawal. Evaluation of Fourier integrals using B-splines. Mathematics of Computation, 39(160): 535–548, October 1982. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lax:1984:CEF

[LA84] M. Lax and G. P. Agrawal. Corrigenda: “Evaluation of Fourier integrals using B-splines” [Math. Comp. 39 (1982), no. 160, 535–548, MR 83i:65021]. Mathematics of Computation, 43(167): 347, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lacroix:1984:CZL

[Lac84] Norbert H. J. Lacroix. On common zeros of Legendre’s associated functions. Mathematics of Computation, 43(167): 243–245, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Laforgia:1984:FIG

[Laf84] Andrea Laforgia. Further inequalities for the gamma function. Mathematics of Computation, 42(166):597–600, April 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lagrange:1983:SSW

[Lag83] Jean Lagrange. Sets of n squares of which any n − 1 have their sum square. Mathematics of Computation, 41(164): 675–681, October 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Land:1981:CPP

[Lan81] Rudolf Land. Computation of Pólya polynomials of primitive permutation group. Mathematics of Computation, 36 (153):267–278, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Larkin:1980:RFF

[Lar80] F. M. Larkin. Root-finding by fitting rational functions. Mathematics of Computation, 35(151):803–816, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Laska:1982:AFM

[Las82a] Michael Laska. An algorithm for finding a minimal Weierstrass equation for an elliptic curve. Mathematics of Computation, 38(157):257–260, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lassey:1982:CCI

[Las82b] Keith R. Lassey. On the computation of certain integrals containing the modified Bessel function I_{0}(ξ). Mathematics of Computation, 39(160):625–637, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lasiecka:1986:GAA

[Las86] I. Lasiecka. Galerkin approximations of abstract parabolic boundary value problems with rough boundary data—L_{p} theory. Mathematics of Computation, 47 (175):55–75, July 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lavoie:1986:SEG

[Lav86] J. L. Lavoie. Some evaluations for the generalized hypergeometric series. Mathematics of Computation, 46(173):215– 218, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lavoie:1987:SSF

[Lav87] J. L. Lavoie. Some summation formulas for the series _{3}F_{2}(1). Mathematics of Computation, 49(179):269–274, July 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Layton:1981:EDD

[Lay81] William J. Layton. Estimates away from a discontinuity for dissipative Galerkin methods for hyperbolic equations. Mathematics of Computation, 36(153):87–92, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Liu:1988:CGP

[LC88] X. Y. Liu and Y. M. Chen. Convergence of a generalized pulse-spectrum technique (GPST) for inverse problems of 1-D diffusion equations in space-time domain. Mathematics of Computation, 51 (184):477–489, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Loiseau:1988:HIM

[LCC88] Jean-Francis Loiseau, Jean-Pierre Codaccioni, and Régis Caboz. Hyperelliptic integrals and multiple hypergeometric series. Mathematics of Computation, 50(182):501–512, April 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Loiseau:1989:IHI

[LCC89] J. F. Loiseau, J. P. Codaccioni, and R. Caboz. Incomplete hyperelliptic integrals and hypergeometric series. Mathematics of Computation, 53(187):335– 342, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

LeRoux:1981:NSS

[Le 81a] A. Y. Le Roux. Numerical stability for some equations of gas dynamics. Mathematics of Computation, 37(156):307– 320, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

LeRoux:1981:SNS

[Le 81b] A. Y. Le Roux. Stability of numerical schemes solving quasi-linear wave equations. Mathematics of Computation, 36 (153):93–105, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1981:FQB

[Leh81] D. H. Lehmer. On Fermat’s quotient, base two. Mathematics of Computation, 36(153):289–290, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1988:SLP

[Leh88a] D. H. Lehmer. The sum of like powers of the zeros of the Riemann zeta function. Mathematics of Computation, 50(181): 265–273, January 1988. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1988:CBG

[Leh88b] Emma Lehmer. Connection between Gaussian periods and cyclic units. Mathematics of Computation, 50(182):535– 541, April 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lenstra:1984:DRC

[Len84] H. W. Lenstra, Jr. Divisors in residue classes. Mathematics of Computation, 42 (165):331–340, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Leon:1980:AFB

[Leo80] Jeffrey S. Leon. On an algorithm for finding a base and a strong generating set for a group given by generating permutations. Mathematics of Computation, 35(151):941–974, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lettl:1986:LBC

[Let86] Günter Lettl. A lower bound for the class number of certain cubic number fields. Mathematics of Computation, 46(174): 659–666, April 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Levin:1980:ACI

[Lev80] David Levin. On accelerating the convergence of infinite double series and integrals. Mathematics of Computation, 35(152):1331–1345, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Levin:1982:PCO

[Lev82a] David Levin. Procedures for computing one- and two-dimensional integrals of functions with rapid irregular oscillations. Mathematics of Computation, 38(158):531–538, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Levin:1982:EDI

[Lev82b] Moshe Levin. On the evaluation of double integrals. Mathematics of Computation, 39(159):173–177, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Levin:1983:ACD

[Lev83] Moshe Levin. On the approximate calculation of double integrals. Mathematics of Computation, 40(161):273– 282, January 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

LeVeque:1986:IBC

[LeV86]  Randall J. LeVeque.  Intermediate

boundary conditions for time-split methods applied to hyperbolic partial differential equations. Mathematics of Computation, 47(175):37–54, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lewanowicz:1985:DDP

[Lew85a] S. Lewanowicz. On the differentialdifference properties of the extended Jacobi polynomials. Mathematics of Computation, 44(170):435–441, April 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lewanowicz:1985:RRH

[Lew85b] Stanisław Lewanowicz. Recurrence relations for hypergeometric functions of unit argument. Mathematics of Computation, 45(172):521–535, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lewanowicz:1986:PPA

[Lew86] S. Lewanowicz. Properties of the polynomials associated with the Jacobi polynomials. Mathematics of Computation, 47 (176):669–682, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lewanowicz:1987:CRR

[Lew87] Stanisław Lewanowicz. Corrigendum:

“Recurrence relations for hypergeometric functions of unit argument” [Math. Comp. 45 (1985), no. 172, 521–535, MR 86m:33004]. Mathematics of Computation, 48(178):853, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1986:MWM

[LG86] J. N. Lyness and G. Giunta. A modification of the Weeks method for numerical inversion of the Laplace transform. Mathematics of Computation, 47(175): 313–322, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ling:1982:EGH

[Lin82] Chih Bing Ling. Evaluation of generalized Howland integrals. Mathematics of Computation, 38(158):593–599, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ling:1986:EHT

[Lin86] Chih Bing Ling. Evaluation of Howlandtype integrals involving tanh and coth functions. Mathematics of Computation, 46(173):219–223, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Laderman:1982:SFP

[LL82] Jack Laderman and Julian D. Laderman. Simplified forecasting by polynomial regression with equally spaced values of the independent variable. Mathematics of Computation, 38(158):601– 610, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1983:CSP

[LL83a] D. H. Lehmer and Emma Lehmer. Cyclotomy with short periods. Mathematics of Computation, 41(164):743– 758, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Li:1983:SHC

[LL83b] Zi Cai Li and Guo Ping Liang. On the simplified hybrid-combined method. Mathematics of Computation, 41(163): 13–25, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lagrange:1986:TTS

[LL86] J. Lagrange and J. Leech. Two triads of squares. Mathematics of Computation, 46(174):751–758, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1987:CR

[LL87] D. H. Lehmer and Emma Lehmer. Cyclotomic resultants. Mathematics of Computation, 48(177):211–216, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1980:QEF

[LM80] J. N. Lyness and G. Monegato. Quadrature error functional expansions for the simplex when the integrand function has singularities at vertices. Mathematics of Computation, 34(149):213–225, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Leavitt:1981:PDF

[LM81] W. G. Leavitt and Albert A. Mullin.

Primes differing by a fixed integer. Mathematics of Computation, 37(156):581– 585, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lagarias:1985:CML

[LMO85] J. C. Lagarias, V. S. Miller, and A. M.

Odlyzko. Computing π(x): The Meissel– Lehmer method. Mathematics of Computation, 44(170):537–560, April 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1986:ID

[LMv86] D. H. Lehmer, K. Mahler, and A. J. van der Poorten. Integers with digits 0 or 1. Mathematics of Computation, 46(174): 683–689, April 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lie:1989:SMC

[LN89] Ivar Lie and Syvert P. Nørsett. Superconvergence for multistep collocation. Mathematics of Computation, 52(185): 65–79, January 1989. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Llorente:1982:RCF

[LO82] P. Llorente and A. V. Oneto. On the real cubic fields. Mathematics of Computation, 39(160):689–692, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

LeVeque:1983:NMB

[LO83] Randall J. LeVeque and Joseph Oliger. Numerical methods based on additive splittings for hyperbolic partial differential equations. Mathematics of Computation, 40(162):469–497, April 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Locher:1981:IUM

[Loc81] F. Locher. Interpolation on uniform meshes by the translates of one function and related attenuation factors. Mathematics of Computation, 37(156):403– 416, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Locher:1989:FGJ

[Loc89] F. Locher. A family of generalized Jacobi polynomials. Mathematics of Computation, 53(187):303–309, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Loh:1989:LCN

[Löh89] Günter Löh. Long chains of nearly doubled primes. Mathematics of Computation, 53(188):751–759, October 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Longman:1983:TEH

[Lon83] I. M. Longman. Table errata: Handbook of mathematical functions with formulas, graphs, and mathematical tables [U.S. Government Printing Office, Washington, D.C., 1964 and MR 29 #4914]. Mathematics of Computation, 41(164):783, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lorenz:1986:CUS

[Lor86] Jens Lorenz. Convergence of upwind schemes for a stationary shock. Mathematics of Computation, 46(173):45–57, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lotstedt:1986:NSN

[LP86] Per Lötstedt and Linda Petzold. Numerical solution of nonlinear differential equations with algebraic constraints. I. convergence results for backward differentiation formulas. Mathematics of Computation, 46(174):491–516, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Llorente:1988:SSC

[LQ88a] Pascual Llorente and Jordi Quer. On the 3-Sylow subgroup of the class group of quadratic fields. Mathematics of Computation, 50(181):321–333, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Llorente:1988:TRC

[LQ88b] Pascual Llorente and Jordi Quer. On totally real cubic fields with discriminant D < 10^{7}. Mathematics of Computation, 50(182):581–594, April 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lynch:1980:HOD

[LR80] Robert E. Lynch and John R. Rice. A high-order difference method for differential equations. Mathematics of Computation, 34(150):333–372, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lubinsky:1984:RCG

[LR84] D. S. Lubinsky and P. Rabinowitz. Rates of convergence of Gaussian quadrature for singular integrands. Mathematics of Computation, 43(167):219–242, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lancaster:1981:SGM

[LS81] P. Lancaster and K. Salkauskas. Surfaces generated by moving least squares methods. Mathematics of Computation, 37(155):141–158, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lubinsky:1983:LZO

[LS83] D. S. Lubinsky and A. Sharif. On the largest zeroes of orthogonal polynomials for certain weights. Mathematics of Computation, 41(163):199–202, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehner:1985:CSI

[LS85] J. Lehner and M. Sheingorn. Computing self-intersections of closed geodesics on finite-sheeted covers of the modular surface. Mathematics of Computation, 44(169):233–240, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lubinsky:1986:CPI

[LS86] D. S. Lubinsky and Avram Sidi. Convergence of product integration rules for functions with interior and endpoint singularities over bounded and unbounded intervals. Mathematics of Computation, 46(173):229–245, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lenstra:1987:PNB

[LS87] H. W. Lenstra, Jr. and R. J. Schoof.

Primitive normal bases for finite fields. Mathematics of Computation, 48(177): 217–231, January 1987. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1989:SPR

[LS89] J. N. Lyness and I. H. Sloan. Some properties of rank-2 lattice rules. Mathematics of Computation, 53(188):627– 637, October 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lubich:1983:RKT

[Lub83] Ch. Lubich. Runge–Kutta theory for Volterra and Abel integral equations of the second kind. Mathematics of Computation, 41(163):87–102, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lubich:1985:FLM

[Lub85] Ch. Lubich. Fractional linear multistep methods for Abel–Volterra integral equations of the second kind. Mathematics of Computation, 45(172):463– 469, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lucas:1983:PII

[Luc83a] Thomas R. Lucas. A posteriori improvements for interpolating periodic splines. Mathematics of Computation, 40(161): 243–251, January 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lucia:1983:EQN

[Luc83b] Angelo Lucia. An explicit quasi-Newton update for sparse optimization calculations. Mathematics of Computation, 40 (161):317–322, January 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lucier:1986:MMN

[Luc86a] Bradley J. Lucier. A moving mesh numerical method for hyperbolic conservation laws. Mathematics of Computation, 46(173):59–69, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lucier:1986:NMD

[Luc86b] Bradley J. Lucier. On nonlocal monotone difference schemes for scalar conservation laws. Mathematics of Computation, 47(175):19–36, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Luecke:1989:NAM

[Lue89] Glenn R. Luecke. Numerical approximation of minimum norm solutions of Kf = g for special K. Mathematics of Computation, 53(188):563–569, October 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lukas:1988:CRR

[Luk88] Mark A. Lukas. Convergence rates for regularized solutions. Mathematics of Computation, 51(183):107–131, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lundin:1980:CFM

[Lun80] L. R. Lundin. A cardinal function method of solution of the equation Δu = u − u^{3}. Mathematics of Computation, 35(151):747–756, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lund:1983:SFQ

[Lun83] John Lund. Sinc function quadrature rules for the Fourier integral. Mathematics of Computation, 41(163):103– 113, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lund:1986:SSG

[Lun86] John Lund. Symmetrization of the Sinc– Galerkin method for boundary value problems. Mathematics of Computation, 47(176):571–588, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lunnon:1988:ISD

[Lun88] W. F. Lunnon. Integer sets with distinct subset-sums. Mathematics of Computation, 50(181):297–320, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Luskin:1980:FEM

[Lus80] Mitchell Luskin. A finite element method for first-order hyperbolic systems. Mathematics of Computation, 35 (152):1093–1112, October 1980. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Landriani:1989:PAD

[LV89] Giovanni Sacchi Landriani and Hervé Vandeven. Polynomial approximation of divergence-free functions. Mathematics of Computation, 52(185):103– 130, January 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ling:1982:EIH

[LW82] Chih Bing Ling and Ming Jing Wu.

Evaluation of integrals of Howland type involving a Bessel function. Mathematics of Computation, 38(157):215– 222, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

MacLean:1980:RCP

[Mac80] D. W. MacLean. Residue classes of the partition function. Mathematics of Computation, 34(149):313–317, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Maeztu:1986:EGC

[Mae86] J. I. Maeztu. An estimate of goodness of cubatures for the unit circle in R^{2}. Mathematics of Computation, 47(176): 651–658, October 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Maher:1980:ETT

[Mah80] David P. Maher. Existence theorems for transforms over finite rings with applications to 2-D convolution. Mathematics of Computation, 35(151):757–765, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mahoney:1981:PFE

[Mah81] J. F. Mahoney. Partial fraction evaluation by an escalation technique. Mathematics of Computation, 36(153):241– 246, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mahler:1982:ZSS

[Mah82] K. Mahler. On the zeros of a special sequence of polynomials. Mathematics of Computation, 39(159):207–212, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mahoney:1985:PFE

[Mah85] J. F. Mahoney. Partial fraction evaluation and incomplete decomposition of a rational function whose denominator contains a repeated polynomial factor. Mathematics of Computation, 44(169): 167–175, January 1985. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Majda:1986:SLS

[Maj86a] George Majda. Stability with large step sizes for multistep discretizations of stiff ordinary differential equations. Mathematics of Computation, 47(176):473– 502, S41–S44, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Majda:1986:SSL

[Maj86b] George Majda. Supplement to stability with large step sizes for multistep discretizations of stiff ordinary differential equations. Mathematics of Computation, 47(176):S41–S44, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Makroglou:1980:CBB

[Mak80] Athena Makroglou. Convergence of a block-by-block method for nonlinear Volterra integro-differential equations. Mathematics of Computation, 35(151): 783–796, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Makroglou:1981:BBM

[Mak81] Athena Makroglou. A block-by-block method for Volterra integro-differential equations with weakly-singular kernel. Mathematics of Computation, 37(155): 95–99, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Malle:1988:PGG

[Mal88] Gunter Malle. Polynomials with Galois groups Aut(M_{22}), M_{22}, and PSL_{3}(𝔽_{4}) · 2_{2} over ℚ. Mathematics of Computation, 51(184):761–768, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Manteuffel:1980:IFT

[Man80] T. A. Manteuffel. An incomplete factorization technique for positive definite linear systems. Mathematics of Computation, 34(150):473–497, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mansfield:1984:CAC

[Man84] Richard Mansfield. A complete axiomatization of computer arithmetic. Mathematics of Computation, 42(166):623– 635, April 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.jstor.org/
                                              stable/2007607.

Marsaglia:1980:GRV

[Mar80a] George Marsaglia. Generating random variables with a t-distribution. Mathematics of Computation, 34(149):235– 236, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Marti:1980:CAC

[Mar80b] J. T. Marti. On the convergence of an algorithm computing minimum-norm solutions of ill-posed problems. Mathematics of Computation, 34(150):521– 527, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Martins:1980:AOI

[Mar80c] M. Madalena Martins. On an accelerated overrelaxation iterative method for linear systems with strictly diagonally dominant matrix. Mathematics of Computation, 35(152):1269–1273, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Martins:1981:NID

[Mar81] M. Madalena Martins. Note on irreducible diagonally dominant matrices and the convergence of the AOR iterative method. Mathematics of Computation, 37(155):101–103, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Markowich:1982:EPI

[Mar82a] Peter A. Markowich. Eigenvalue problems on infinite intervals. Mathematics of Computation, 39(160):421–441, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Marsden:1982:SIK

[Mar82b] M. J. Marsden. Spline interpolation at knot averages on a two-sided geometric mesh. Mathematics of Computation, 38(157):113–126, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Massidda:1983:ACC

[Mas83] Vittorio Massidda. Analytical calculation of a class of integrals containing exponential and trigonometric functions. Mathematics of Computation, 41(164): 555–557, October 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mastroianni:1989:CPF

[Mas89] Giuseppe Mastroianni. On the convergence of product formulas for the evaluation of certain two-dimensional Cauchy principal value integrals. Mathematics of Computation, 52(185):95–101, January 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mattheij:1987:CSB

[Mat87] R. M. M. Mattheij. On the computation of solutions of boundary value problems on infinite intervals. Mathematics of Computation, 48(178):533–549, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McCormick:1981:MRM

[McC81a] Stephen F. McCormick. A mesh refinement method for Ax = λBx. Mathematics of Computation, 36(154):485– 498, April 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McCullagh:1981:RCS

[McC81b] Peter McCullagh. A rapidly convergent series for computing ψ(z) and its derivatives. Mathematics of Computation, 36 (153):247–248, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McCabe:1983:QDA

[McC83] J. H. McCabe. The quotient-difference algorithm and the Padé table: an alternative form and a general continued fraction. Mathematics of Computation, 41(163):183–197, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McCurley:1984:EE

[McC84a] Kevin S. McCurley. Explicit estimates for θ(x; 3, l) and ψ(x; 3, l). Mathematics of Computation, 42(165):287–296, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McCurley:1984:EEE

[McC84b] Kevin S. McCurley. Explicit estimates for the error term in the prime number theorem for arithmetic progressions. Mathematics of Computation, 42(165): 265–285, January 1984. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McCarthy:1986:EIM

[McC86] D. P. McCarthy. Effect of improved multiplication efficiency on exponentiation algorithms derived from addition chains. Mathematics of Computation, 46(174): 603–608, April 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McDaniel:1989:SPR

[McD89] Wayne L. McDaniel. Some pseudoprimes and related numbers having special forms. Mathematics of Computation, 53(187):407–409, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McIntyre:1986:BIS

[McI86a] E. A. McIntyre, Jr. Boundary integral solutions of the heat equation. Mathematics of Computation, 46(173):71–79, S1–S14, January 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McIntyre:1986:SBI

[McI86b] E. A. McIntyre, Jr. Supplement to boundary integral solutions of the heat equations. Mathematics of Computation, 46(173):S1–S14, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McLean:1986:SGM

[McL86] W. McLean. A spectral Galerkin method for a boundary integral equation. Mathematics of Computation, 47(176):597– 607, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Meek:1980:MVT

[Mee80] D. Meek. A mean value theorem for linear functionals. Mathematics of Computation, 35(151):797–802, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Meinardus:1985:RLM

[Mei85] Günter Meinardus. Remark on a lemma: “On a method of asymptotic evaluation of multiple integrals” [Math. Comp. 37 (1981), no. 156, 509–521, MR 83b:41040] by R. Wong and J. P. McClure. Mathematics of Computation, 45(171):197– 198, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mettke:1986:CIS

[Met86] Holger Mettke. Convex interpolation by splines of arbitrary degree. Mathematics of Computation, 46(174):567–576, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Meyer:1984:SIE

[Mey84] Burnett Meyer. Some inequalities for elementary mean values. Mathematics of Computation, 42(165):193–194, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Michelson:1983:STD

[Mic83] Daniel Michelson. Stability theory of difference approximations for multidimensional initial-boundary value problems.

Mathematics of Computation, 40(161): 1–45, January 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Michelson:1987:CTD

[Mic87] Daniel Michelson. Convergence theorem for difference approximations of hyperbolic quasi-initial-boundary value problems. Mathematics of Computation, 49 (180):445–459, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Miel:1980:MSE

[Mie80] George J. Miel. Majorizing sequences and error bounds for iterative methods. Mathematics of Computation, 34(149): 185–202, January 1980. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Miklavcic:1989:AWN

[Mik89] Milan Miklavčič. Approximations for weakly nonlinear evolution equations. Mathematics of Computation, 53(188): 471–484, October 1989. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Milgram:1985:GIE

[Mil85a] M. S. Milgram. The generalized integroexponential function. Mathematics of Computation, 44(170):443–458, April 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Milner:1985:MFE

[Mil85b] F. A. Milner. Mixed finite element methods for quasilinear second-order elliptic problems. Mathematics of Computation, 44(170):303–320, April 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Milovanovic:1987:CLS

[Mil87]  G. V. Milovanović.  Corrigendum:

“Least squares approximation with constraints”. Mathematics of Computation, 48(178):854, April 1987. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Minoli:1980:IFG

[Min80a] Daniel Minoli. Inductive formulae for general sum operations. Mathematics of Computation, 34(150):543–545, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Minoli:1980:INH

[Min80b] Daniel Minoli. Issues in nonlinear hyperperfect numbers. Mathematics of Computation, 34(150):639–645, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Misici:1984:NST

[Mis84] Luciano Misici. Numerical solution of two transcendental equations. Mathematics of Computation, 42(166):589– 595, April 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mittelmann:1983:EAB

[Mit83a] H. D. Mittelmann. An efficient algorithm for bifurcation problems of variational inequalities. Mathematics of Computation, 41(164):473–485, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mityagin:1983:QPL

[Mit83b] Boris Mityagin. Quadratic pencils and least-squares piecewise-polynomial approximation. Mathematics of Computation, 40(161):283–300, January 1983.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Miyamoto:1989:EP

[MM89] Ian Miyamoto and M. Ram Murty. Elliptic pseudoprimes. Mathematics of Computation, 53(187):415–430, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mori:1988:FDT

[MMM88] Shigefumi Mori, David R. Morrison, and Ian Morrison. On four-dimensional terminal quotient singularities. Mathematics of Computation, 51(184):769– 786, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McCurdy:1984:ACD

[MNP84] A. McCurdy, K. C. Ng, and B. N. Parlett. Accurate computation of divided differences of the exponential function. Mathematics of Computation, 43(168): 501–528, October 1984. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Massias:1989:EBM

[MNR89] Jean-Pierre Massias, Jean-Louis Nicolas, and Guy Robin. Effective bounds for the maximal order of an element in the symmetric group. Mathematics of Computation, 53(188):665–678, October 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Meinardus:1989:APP

[MNSS89] G. Meinardus, G. Nürnberger, M. Sommer, and H. Strauss. Algorithms for piecewise polynomials and splines with free knots. Mathematics of Computation, 53(187):235–247, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mollin:1987:CNQ

[Mol87] R. A. Mollin. Class numbers of quadratic fields determined by solvability of Diophantine equations. Mathematics of Computation, 48(177):233–242, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mond:1981:AEL

[Mon81] B. Mond. On algorithmic equivalence in linear fractional programming. Mathematics of Computation, 37(155):185– 187, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Monahan:1985:ARN

[Mon85a] John F. Monahan. Accuracy in random number generation. Mathematics of Computation, 45(172):559–568, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Montgomery:1985:MMT

[Mon85b] Peter L. Montgomery. Modular multiplication without trial division. Mathematics of Computation, 44(170):519– 521, April 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.org/
                                              journals/mcom/1985-44-170/S0025-
                                              5718-1985-0777282-X/.

Monegato:1986:QFF

[Mon86] Giovanni Monegato. Quadrature formulas for functions with poles near the interval of integration. Mathematics of Computation, 47(175):301–312, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Montgomery:1987:SPE

[Mon87] Peter L. Montgomery. Speeding the Pollard and elliptic curve methods of factorization. Mathematics of Computation, 48(177):243–264, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Monk:1988:IFE

[Mon88] P. B. Monk. An iterative finite element method for approximating the biharmonic equation. Mathematics of Computation, 51(184):451–476, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Morain:1989:DEP

[Mor89] François Morain. On the lcm of the differences of eight primes. Mathematics of Computation, 52(185):225–229, January 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigenda [Mor90].

Morain:1990:CLD

[Mor90] François Morain. Corrigenda: “On the lcm of the differences of eight primes” [Math. Comput. 52 (1989), no. 185, 225– 229]. Mathematics of Computation, 54 (190):911, April 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [Mor89].

Mercier:1981:EAM

[MORR81] B. Mercier, J. Osborn, J. Rappaz, and P.-A. Raviart. Eigenvalue approximation by mixed and hybrid methods. Mathematics of Computation, 36(154): 427–453, April 1981. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mossige:1981:ACR

[Mos81] Svein Mossige. Algorithms for computing the h-range of the postage stamp problem. Mathematics of Computation, 36(154):575–582, April 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mosier:1986:RNP

[Mos86] Ronald G. Mosier. Root neighborhoods of a polynomial. Mathematics of Computation, 47(175):265–273, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mostrel:1989:SNS

[Mos89] Marco Mosché Mostrel. On some numerical schemes for transonic flow problems. Mathematics of Computation, 52(186): 587–613, April 1989. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McCabe:1985:ASG

[MP85] J. H. McCabe and G. M. Phillips. Aitken sequences and generalized Fibonacci numbers. Mathematics of Computation, 45(172):553–558, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Moody:1987:CCD

[MP87] R. V. Moody and J. Patera. Computation of character decompositions of class functions on compact semisimple Lie groups. Mathematics of Computation, 48(178):799–827, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mathon:1981:CST

[MPR81] Rudolf A. Mathon, Kevin T. Phelps, and Alexander Rosa. A class of Steiner triple systems of order 21 and associated Kirkman systems. Mathematics of Computation, 37(155):209–222, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Markowich:1983:CMB

[MR83a] Peter A. Markowich and Christian A.

Ringhofer. Collocation methods for boundary value problems on “long” intervals. Mathematics of Computation, 40 (161):123–150, January 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McCormick:1983:UMS

[MR83b] S. F. McCormick and J. W. Ruge. Unigrid for multigrid simulation. Mathematics of Computation, 41(163):43–62, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mattheij:1984:OSI

[MS84] R. M. M. Mattheij and G. W. M.

Staarink. On optimal shooting intervals. Mathematics of Computation, 42(165): 25–40, January 1984. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mjolsness:1987:SPC

[MS87] R. C. Mjolsness and Blair Swartz. Some plane curvature approximations. Mathematics of Computation, 49(179):215– 230, July 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Meinardus:1980:OPN

[MT80] Günter Meinardus and G. D. Taylor. Optimal partitioning of Newton’s method for calculating roots. Mathematics of Computation, 35(152):1221– 1230, October 1980. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McCormick:1986:FAC

[MT86] S. F. McCormick and J. W. Thomas.

The fast adaptive composite grid (FAC) method for elliptic equations. Mathematics of Computation, 46(174):439– 456, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Munz:1981:UEC

[Mun81] Harry Munz. Uniform expansions for a class of finite difference schemes for elliptic boundary value problems. Mathematics of Computation, 36(153):155– 170, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Murray:1982:FF

[Mur82] Francis J. Murray. Formulas for factorial N. Mathematics of Computation, 39 (160):655–662, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Manteuffel:1986:NSS

[MW86a] Thomas A. Manteuffel and Andrew B. White, Jr. The numerical solution of second-order boundary value problems on nonuniform meshes. Mathematics of Computation, 47(176):511–535, S53– S55, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Manteuffel:1986:SNS

[MW86b] Thomas A. Manteuffel and Andrew B. White, Jr. Supplement to the numerical solution of second-order boundary value problems on nonuniform meshes. Mathematics of Computation, 47(176): S53–S55, October 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Milovanovic:1986:LSA

[MW86c] Gradimir V. Milovanović and Staffan Wrigge. Least squares approximation with constraints. Mathematics of Computation, 46(174):551–565, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Markowich:1988:IAT

[MZ88] Peter A. Markowich and Miloš A.

Zlámal. Inverse-average-type finite element discretizations of selfadjoint second-order elliptic problems. Mathematics of Computation, 51(184):431– 449, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Marsaglia:1989:NSS

[MZM89] George Marsaglia, Arif Zaman, and John C. W. Marsaglia. Numerical solution of some classical differentialdifference equations. Mathematics of Computation, 53(187):191–201, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nakao:1984:CGM

[Nak84] Mitsuhiro Nakao. A collocation-H^{−1}- Galerkin method for some elliptic equations. Mathematics of Computation, 42(166):417–426, April 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nassif:1987:EFD

[Nas87] Nabil R. Nassif. Eigenvalue finite difference approximations for regular and singular Sturm–Liouville problems. Mathematics of Computation, 49(180):561– 580, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Naur:1983:NIF

[Nau83] Thorkil Naur. New integer factorizations. Mathematics of Computation, 41 (164):687–695, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Noll:1989:C

[NB89] Landon Curt Noll and David I. Bell. nclusters for 1 < n < 7. Mathematics of Computation, 53(187):439–444, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Neta:1980:DBP

[Net80] Beny Neta. On determination of bestpossible constants in integral inequalities involving derivatives. Mathematics of Computation, 35(152):1191–1193, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Neubauer:1987:FDA

[Neu87] A. Neubauer. Finite-dimensional approximation of constrained Tikhonovregularized solutions of ill-posed linear operator equations. Mathematics of Computation, 48(178):565–583, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Newman:1984:ABS

[New84] J. N. Newman. Approximations for the Bessel and Struve functions. Mathematics of Computation, 43(168):551– 556, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Newman:1985:SVF

[New85] D. J. Newman. A simplified version of the fast algorithms of Brent and Salamin. Mathematics of Computation, 44(169):207–210, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Newman:1986:CWM

[New86] D. J. Newman. Computing when multiplications cost nothing. Mathematics of Computation, 46(173):255–257, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nielson:1983:MIS

[Nie83] Gregory M. Nielson. A method for interpolating scattered data based upon a minimum norm network. Mathematics of Computation, 40(161):253–271, January 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Niederreiter:1989:STC

[Nie89] Harald Niederreiter. The serial test for congruential pseudorandom numbers generated by inversions. Mathematics of Computation, 52(185):135–144, January 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Noll:1980:MP

[NN80] Curt Noll and Laura Nickel. The 25 th and 26 th Mersenne primes. Mathematics of Computation, 35(152):1387– 1390, October 1980. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nocedal:1980:UQN

[Noc80] Jorge Nocedal. Updating quasi-Newton matrices with limited storage. Mathematics of Computation, 35(151):773– 782, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nochetto:1989:SEM

[Noc89] Ricardo H. Nochetto. A stable extrapolation method for multidimensional degenerate parabolic problems. Mathematics of Computation, 53(188):455– 470, October 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nour-Omid:1987:HIS

[NOPEJ87] Bahram Nour-Omid, Beresford N. Parlett, Thomas Ericsson, and Paul S. Jensen. How to implement the spectral transformation. Mathematics of Computation, 48(178):663–673, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Newman:1984:SAS

[NS84] Morris Newman and Daniel Shanks. On a sequence arising in series for π. Mathematics of Computation, 42(165):199– 217, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Neumaier:1985:DDS

[NS85] A. Neumaier and A. Schäfer. Divided differences, shift transformations and Larkin’s root finding method. Mathematics of Computation, 45(171):181– 196, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Newman:1987:NVG

[NT87] Morris Newman and Robert C. Thompson. Numerical values of Goldberg’s coefficients in the series for log(e^{x}e^{y}). with microfiche supplement. Mathematics of Computation, 48(177):265–271, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nagaraja:1982:EI

[NV82] K. S. Nagaraja and G. R. Verma. Evaluation of the integral ∫_{0}^{p} u^{n}e^{−u^{2}}(u + x)^{−1}du. Mathematics of Computation, 39(159):179–194, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nochetto:1988:ELS

[NV88] Ricardo H. Nochetto and Claudio Verdi.

An efficient linear scheme to approximate parabolic free boundary problems: error estimates and implementation. Mathematics of Computation, 51 (183):27–53, July 1988. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nemeth:1982:PTP

[NZ82] Géza Németh and Magda Zimányi. Polynomial type Padé approximants. Mathematics of Computation, 38(158):553– 565, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Odlyzko:1987:DSB

[Odl87] A. M. Odlyzko. On the distribution of spacings between zeros of the zeta function. Mathematics of Computation, 48 (177):273–308, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Osher:1985:ECS

[OHW85] Stanley Osher, Mohamed Hafez, and Woodrow Whitlow, Jr. Entropy condition satisfying approximations for the full potential equation of transonic flow. Mathematics of Computation, 44(169): 1–29, January 1985. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Okecha:1987:QFC

[Oke87] G. E. Okecha. Quadrature formulae for Cauchy principal value integrals of oscillatory kind. Mathematics of Computation, 49(179):259–268, July 1987.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Okui:1983:SIR

[Oku83] Shigehiko Okui. Some integrals relating to the I_{e}-function. Mathematics of Computation, 41(164):613–622, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Olmsted:1986:TFG

[Olm86] Coert Olmsted. Two formulas for the general multivariate polynomial which interpolates a regular grid on a simplex. Mathematics of Computation, 47(175): 275–284, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Olver:1988:EBL

[Olv88] F. W. J. Olver. Error bounds for linear recurrence relations. Mathematics of Computation, 50(182):481–499, April 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Onumanyi:1984:NSS

[OO84] P. Onumanyi and E. L. Ortiz. Numerical solution of stiff and singularly perturbed boundary value problems with a segmented-adaptive formulation of the tau method. Mathematics of Computation, 43(167):189–203, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Osher:1982:UDS

[OS82] Stanley Osher and Fred Solomon. Upwind difference schemes for hyperbolic systems of conservation laws. Mathematics of Computation, 38(158):339– 374, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Osher:1983:NAN

[OS83] Stanley Osher and Richard Sanders.

Numerical approximations to nonlinear conservation laws with locally varying time and space grids. Mathematics of Computation, 41(164):321–336, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

ORiordan:1986:ASR

[OS86a] Eugene O’Riordan and Martin Stynes.

An analysis of a superconvergence result for a singularly perturbed boundary value problem. Mathematics of Computation, 46(173):81–92, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

ORiordan:1986:UAF

[OS86b] Eugene O’Riordan and Martin Stynes.

A uniformly accurate finite-element method for a singularly perturbed onedimensional reaction-diffusion problem. Mathematics of Computation, 47(176): 555–570, October 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Osher:1988:CDA

[OT88] Stanley Osher and Eitan Tadmor. On the convergence of difference approximations to scalar conservation laws. Mathematics of Computation, 50(181): 19–51, January 1988. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Paige:1980:EAS

[Pai80] C. C. Paige. Error analysis of some techniques for updating orthogonal decompositions. Mathematics of Computation, 34(150):465–471, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Paine:1982:CSL

[Pai82] J. Paine. Correction of Sturm–Liouville eigenvalue estimates. Mathematics of Computation, 39(160):415–420, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Papatheodorou:1983:BAI

[Pap83] Theodore S. Papatheodorou. Block AOR iteration for nonsymmetric matrices. Mathematics of Computation, 41 (164):511–525, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Parter:1980:RSC

[Par80] Seymour V. Parter. On the roles of “stability” and “convergence” in semidiscrete projection methods for initial-value problems. Mathematics of Computation, 34(149):127–154, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Parson:1983:LBN

[Par83] L. Alayne Parson. A lower bound for the norm of the theta operator. Mathematics of Computation, 41(164):683– 685, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pasciak:1980:SPS

[Pas80]  Joseph E. Pasciak.  Spectral and

pseudo spectral methods for advection equations. Mathematics of Computation, 35(152):1081–1092, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pasciak:1984:NSP

[Pas84] Joseph E. Pasciak. A new scalar potential formulation of the magnetostatic field problem. Mathematics of Computation, 43(168):433–445, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Petho:1986:PPP

[PdW86] A. Pethö and B. M. M. de Weger. Products of prime powers in binary recurrence sequences. part I: The hyperbolic case, with an application to the generalized Ramanujan–Nagell equation. Mathematics of Computation, 47(176):713– 727, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pennline:1981:ICR

[Pen81] James A. Pennline. Improving convergence rate in the method of successive approximations. Mathematics of Computation, 37(155):127–134, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pereira:1985:ECF

[Per85] N. Costa Pereira. Estimates for the Chebyshev function ψ(x) − θ(x). Mathematics of Computation, 44(169):211– 221, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigendum [Per87].

Pereira:1987:CEC

[Per87] N. Costa Pereira. Corrigendum: “Estimates for the Chebyshev function ψ(x)− θ(x)” [Math. Comp. 44 (1985), no. 169, 211–221, MR 86k:11005]. Mathematics of Computation, 48(177):447, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [Per85].

Pexton:1981:TEF

[Pex81] Robert L. Pexton. Table errata: Formulas and theorems for the special functions of mathematical physics [third edition, Springer, New York, 1966 and F. Oberhettinger MR 38 #1291] by W. Magnus and R. P. Soni. Mathematics of Computation, 36(153):316–317, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Plesken:1984:LSD

[PH84] W. Plesken and W. Hanrath. The lattices of six-dimensional Euclidean space. Mathematics of Computation, 43(168): 573–587, October 1984. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Philip:1980:CPC

[Phi80] J. R. Philip. The convergence and partial convergence of alternating series. Mathematics of Computation, 35(151): 907–916, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Philip:1987:TES

[Phi87] J. R. Maclaurin Philip. Table errata: “The symmetrical Euler-summation formula” [Math. Sci. 6 (1981), no. 1, 35– 41, MR 82g:40002]. Mathematics of Computation, 48(178):851, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Piazza:1981:SRK

[Pia81] Giuseppe Piazza. On the BN stability of the Runge–Kutta methods. Mathematics of Computation, 37(156):399– 401, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Piessens:1984:SEF

[Pie84] R. Piessens. A series expansion for the first positive zero of the Bessel functions. Mathematics of Computation, 42(165):

195–197, January 1984. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pierre:1988:CPN

[Pie88] Roger Pierre. Convergence properties and numerical approximation of the solution of the Mindlin plate bending problem. Mathematics of Computation, 51 (183):15–25, July 1988. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pitkaranta:1980:LSC

[Pit80] Juhani Pitkäranta. Local stability conditions for the Babuška method of Lagrange multipliers. Mathematics of Computation, 35(152):1113–1129, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pitkaranta:1981:FEM

[Pit81] Juhani Pitkäranta. The finite element method with Lagrange multipliers for domains with corners. Mathematics of Computation, 37(155):13–30, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pohst:1987:CIE

[Poh87] M. Pohst. On computing isomorphisms of equation orders. Mathematics of Computation, 48(177):309–314, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pomerance:1981:DP

[Pom81] Carl Pomerance. On the distribution of pseudoprimes. Mathematics of Computation, 37(156):587–593, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pomerance:1987:VSP

[Pom87] Carl Pomerance. Very short primality proofs. Mathematics of Computation, 48 (177):315–322, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Porsching:1985:EER

[Por85] T. A. Porsching. Estimation of the error in the reduced basis method solution of nonlinear equations. Mathematics of Computation, 45(172):487–496, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Plesken:1980:MFIa

[PP80a] Wilhelm Plesken and Michael Pohst. On maximal finite irreducible subgroups of GL(n, Z): III. the nine dimensional case. Mathematics of Computation, 34(149): 245–258, January 1980. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Plesken:1980:MFIb

[PP80b] Wilhelm Plesken and Michael Pohst. On maximal finite irreducible subgroups of GL(n, Z): IV. remarks on even dimensions with applications to n = 8. Mathematics of Computation, 34(149):259– 275, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Plesken:1980:MFIc

[PP80c] Wilhelm Plesken and Michael Pohst. On maximal finite irreducible subgroups of GL(n, Z): V. the eight dimensional case and a complete description of dimensions less than ten. Mathematics of Computation, 34(149):277–301, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Plesken:1985:CIL

[PP85a] W. Plesken and M. Pohst. Constructing integral lattices with prescribed minimum. I. Mathematics of Computation, 45(171):209–221, S5–S16, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Plesken:1985:SCI

[PP85b] W. Plesken and M. Pohst. Supplement to constructing integral lattices with prescribed minimum. I. Mathematics of Computation, 45(171):S5–S16, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pritchard:1983:EPA

[Pri83] Paul A. Pritchard. Eighteen primes in arithmetic progression. Mathematics of Computation, 41(164):697, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pritchard:1985:LAP

[Pri85] Paul A. Pritchard. Long arithmetic progressions of primes: some old, some new. Mathematics of Computation, 45(171): 263–267, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pitkaranta:1983:ASM

[PS83] J. Pitkäranta and R. Stenberg. Analysis of some mixed finite element methods for plane elasticity equations. Mathematics of Computation, 41(164):399– 423, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pitkaranta:1985:MVS

[PS85] Juhani Pitkäranta and Tuomo Saarinen.

A multigrid version of a simple finite element method for the Stokes problem.

Mathematics of Computation, 45(171): 1–14, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Parlett:1982:ELE

[PSS82] B. N. Parlett, H. Simon, and L. M. Stringer. On estimating the largest eigenvalue with the Lánczos algorithm. Mathematics of Computation, 38(157): 153–165, January 1982. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pintz:1989:ISP

[PSS89] János Pintz, William L. Steiger, and Endre Szemerédi. Infinite sets of primes with fast primality tests and quick generation of large primes. Mathematics of Computation, 53(187):399–406, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pomerance:1980:P

[PSW80] Carl Pomerance, J. L. Selfridge, and Samuel S. Wagstaff, Jr. The pseudoprimes to 25 · 10^{9}. Mathematics of Computation, 35(151):1003–1026, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pasquini:1985:GCM

[PT85] L. Pasquini and D. Trigiante. A globally convergent method for simultaneously finding polynomial roots. Mathematics of Computation, 44(169):135– 149, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Parlett:1985:LAL

[PTL85] Beresford N. Parlett, Derek R. Taylor, and Zhishun A. Liu. A look-ahead Lánczos algorithm for unsymmetric matrices. Mathematics of Computation, 44 (169):105–124, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Puckett:1989:CRP

[Puc89] Elbridge Gerry Puckett. Convergence of a random particle method to solutions of the Kolmogorov equation u_{t} = νu_{xx} + u(1 − u). Mathematics of Computation, 52(186):615–645, April 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Patterson:1985:SPC

[PW85] C. D. Patterson and H. C. Williams.

Some periodic continued fractions with long periods. Mathematics of Computation, 44(170):523–532, April 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pohst:1982:ECFb

[PWZ82] Michael Pohst, Peter Weiler, and Hans Zassenhaus. On effective computation of fundamental units. II. Mathematics of Computation, 38(157):293–329, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pohst:1982:ECFa

[PZ82] Michael Pohst and Hans Zassenhaus.

On effective computation of fundamental units. I. Mathematics of Computation, 38(157):275–291, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Queen:1981:MFA

[Que81] Larissa Queen. Modular functions arising from some finite groups. Mathematics of Computation, 37(156):547– 580, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rabinowitz:1980:EDP

[Rab80] Philip Rabinowitz. The exact degree of precision of generalized Gauss–Kronrod integration rules. Mathematics of Computation, 35(152):1275–1283, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rabinowitz:1983:GKI

[Rab83] Philip Rabinowitz. Gauss–Kronrod integration rules for Cauchy principal value integrals. Mathematics of Computation, 41(163):63–78, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rabinowitz:1985:CGK

[Rab85] Philip Rabinowitz. Correction: “Gauss– Kronrod integration rules for Cauchy principal value integrals” [Math. Comp. 41 (1983), no. 1, 63–78, MR 84i:65029]. Mathematics of Computation, 45(171): 277, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rabinowitz:1986:DGK

[Rab86a] Philip Rabinowitz. On the definiteness of Gauss–Kronrod integration rules. Mathematics of Computation, 46(173): 225–227, January 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rabinowitz:1986:RCG

[Rab86b] Philip Rabinowitz. Rates of convergence of Gauss, Lobatto, and Radau integration rules for singular integrands. Mathematics of Computation, 47(176):625– 638, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rabinowitz:1988:CRP

[Rab88a] Philip Rabinowitz. Convergence results for piecewise linear quadratures for Cauchy principal value integrals. Mathematics of Computation, 51(184):741– 747, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rabinowitz:1988:CGK

[Rab88b] Philip Rabinowitz. Corrigenda: “Gauss– Kronrod integration rules for Cauchy principal value integrals” [Math. Comp. 41 (1983), no. 163, 63–78, MR 84i:65029]. Mathematics of Computation, 50(182):655, April 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rall:1983:RIO

[Ral83] L. B. Rall. Representations of intervals and optimal error bounds. Mathematics of Computation, 41(163):219–227, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rayner:1988:WUDa

[Ray88a] Francis J. Rayner. Weak uniform distribution for divisor functions. I. Mathematics of Computation, 50(181):335– 342, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rayner:1988:WUDb

[Ray88b] Francis J. Rayner. Weak uniform distribution for divisor functions. II. Mathematics of Computation, 51(183):331– 337, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Reichel:1985:PAC

[Rei85] Lothar Reichel. On polynomial approximation in the complex plane with application to conformal mapping. Mathematics of Computation, 44(170):425– 433, April 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Reichert:1986:EDN

[Rei86] Markus A. Reichert. Explicit determination of nontrivial torsion structures of elliptic curves over quadratic number fields. Mathematics of Computation, 46(174):637–658, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rice:1981:PLS

[RHD81] John R. Rice, Elias N. Houstis, and Wayne R. Dyksen. A population of linear, second order, elliptic partial differential equations on rectangular domains. part I. Mathematics of Computation, 36(154):475–484, April 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rheinboldt:1984:DAS

[Rhe84] Werner C. Rheinboldt. Differentialalgebraic systems as differential equations on manifolds. Mathematics of Computation, 43(168):473–482, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Richter:1981:NIS

[Ric81] Gerard R. Richter. Numerical identification of a spatially varying diffusion coefficient. Mathematics of Computation, 36(154):375–386, April 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Richter:1988:OOE

[Ric88] Gerard R. Richter. An optimal-order error estimate for the discontinuous Galerkin method. Mathematics of Computation, 50(181):75–88, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Riley:1983:ACI

[Ril83] Robert Riley. Applications of a computer implementation of Poincaré’s theorem on fundamental polyhedra. Mathematics of Computation, 40(162):607– 632, April 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rabinowitz:1989:NIR

[RL89] P. Rabinowitz and D. S. Lubinsky.

Noninterpolatory integration rules for Cauchy principal value integrals. Mathematics of Computation, 53(187):279– 295, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Roberts:1989:CRW

[Rob89] Stephen Roberts. Convergence of a random walk method for the Burgers equation. Mathematics of Computation, 52 (186):647–673, April 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rose:1983:NMF

[Ros83] Michael E. Rose. Numerical methods for flows through porous media. I. Mathematics of Computation, 40(162):435– 467, April 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rotkiewicz:1982:ELP

[Rot82] A. Rotkiewicz. On Euler Lehmer pseudoprimes and strong Lehmer pseudoprimes with parameters L, Q in arithmetic progressions. Mathematics of Computation, 39(159):239–247, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rotkiewicz:1984:C

[Rot84] A. Rotkiewicz. On the congruence 2^{n−2} ≡ 1(mod n). Mathematics of Computation, 43(167):271–272, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rannacher:1982:SOE

[RS82] Rolf Rannacher and Ridgway Scott. Some optimal error estimates for piecewise linear finite element approximations. Mathematics of Computation, 38(158):437–445, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ruck:1987:NEC

[Rüc87] Hans-Georg Rück. A note on elliptic curves over finite fields. Mathematics of Computation, 49(179):301–304, July 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rusin:1989:CGO

[Rus89] David J. Rusin. The cohomology of the groups of order 32. Mathematics of Computation, 53(187):359–385, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rodrigue:1984:PIB

[RW84] Garry Rodrigue and Donald Wolitzer. Preconditioning by incomplete block cyclic reduction. Mathematics of Computation, 42(166):549–565, April 1984.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Saad:1981:KSM

[Saa81] Y. Saad. Krylov subspace methods for solving large unsymmetric linear systems. Mathematics of Computation, 37(155):105–126, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Saad:1984:CAT

[Saa84] Youcef Saad. Chebyshev acceleration techniques for solving nonsymmetric eigenvalue problems. Mathematics of Computation, 42(166):567–588, April 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Saad:1987:LMS

[Saa87] Youcef Saad. On the Lánczos method for solving symmetric systems with several right-hand sides. Mathematics of Computation, 48(178):651–662, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Salzer:1980:MCF

[Sal80] Herbert E. Salzer. Misstatements in Milne-Thomson, Calculus of Finite Differences, Macmillan, London, 1933. Mathematics of Computation, 34(149): 323–324, January 1980. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Salzer:1981:EHN

[Sal81] Herbert E. Salzer. Errata: A history of numerical analysis from the 16th through the 19th century [Springer, New York, 1977]. MR 58 #4774 by H. H. Goldstine. Mathematics of Computation, 36(153): 319–320, January 1981. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Salzer:1984:TET

[Sal84] Herbert E. Salzer. Table errata: Table of sines and cosines to ten decimal places at thousandths of a degree [Pergamon, New York, 1962 and MR 26 #361] by Salzer and N. Levine. Mathematics of Computation, 43(167):346, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sanders:1983:CMF

[San83] Richard Sanders. On convergence of monotone finite difference schemes with variable spatial differencing. Mathematics of Computation, 40(161):91–106, January 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sanders:1988:TOA

[San88] Richard Sanders. A third-order accurate variation nonexpansive difference scheme for single nonlinear conservation laws. Mathematics of Computation, 51 (184):535–558, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Saranen:1987:LEE

[Sar87] Jukka Saranen. Local error estimates for some Petrov–Galerkin methods applied to strongly elliptic equations on curves. Mathematics of Computation, 48(178): 485–502, April 1987. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Scarowsky:1984:NDE

[SB84] Manny Scarowsky and Abraham Boyarsky. A note on the Diophantine equation x^{n} + y^{n} + z^{n} = 3. Mathematics of Computation, 42(165):235–237, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schatzman:1989:NAW

[SB89] Michelle Schatzman and Michel Bercovier. Numerical approximation of a wave equation with unilateral constraints. Mathematics of Computation, 53(187):55–79, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schatz:1980:WDM

[Sch80a] Alfred H. Schatz. A weak discrete maximum principle and stability of the finite element method in L_{∞} on plane polygonal domains. I. Mathematics of Computation, 34(149):77–91, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schonfelder:1980:VHA

[Sch80b] J. L. Schonfelder. Very high accuracy Chebyshev expansions for the basic trigonometric functions. Mathematics of Computation, 34(149):237–244, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schneider:1981:PIW

[Sch81] Claus Schneider. Product integration for weakly singular integral equations. Mathematics of Computation, 36(153): 207–213, January 1981. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Scheid:1983:ANS

[Sch83a] Robert E. Scheid, Jr. The accurate numerical solution of highly oscillatory ordinary differential equations. Mathematics of Computation, 41(164):487– 509, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schoof:1983:CGC

[Sch83b] R. J. Schoof. Class groups of complex quadratic fields. Mathematics of Computation, 41(163):295–302, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schwindt:1983:TSC

[Sch83c] H. Schwindt. Three summation criteria for Fermat’s Last Theorem. Mathematics of Computation, 40(162):715– 716, April 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schaffer:1984:HOM

[Sch84a] Steve Schaffer. Higher order multi-grid methods. Mathematics of Computation, 43(167):89–115, S1–19, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schaffer:1984:SHO

[Sch84b] Steve Schaffer. Supplement to higher order multi-grid methods. Mathematics of Computation, 43(167):S1–S19, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schechter:1984:SCR

[Sch84c] Eric Schechter. Sharp convergence rates for nonlinear product formulas. Mathematics of Computation, 43(167):135– 155, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Scheid:1985:DMP

[Sch85a] Robert E. Scheid, Jr. Difference methods for problems with different time scales. Mathematics of Computation, 44(169):81–92, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schonbek:1985:SOC

[Sch85b] Maria E. Schonbek. Second-order conservative schemes and the entropy condition. Mathematics of Computation, 44(169):31–38, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schoof:1985:ECF

[Sch85c] René Schoof. Elliptic curves over finite fields and the computation of square roots mod p. Mathematics of Computation, 44(170):483–494, April 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schmeiser:1986:ASB

[Sch86] Christian Schmeiser. Approximate solution of boundary value problems on infinite intervals by collocation methods. Mathematics of Computation, 46(174): 479–490, April 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Scraton:1988:CST

[Scr88] R. E. Scraton. A comparison of some Taylor and Chebyshev series. Mathematics of Computation, 50(181):207– 213, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Serbin:1980:FCC

[Ser80] Steven M. Serbin. On factoring a class of complex symmetric matrices without pivoting. Mathematics of Computation, 35(152):1231–1234, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Setzer:1980:DAI

[Set80] Bennett Setzer. The determination of all imaginary, quartic, Abelian number fields with class number 1. Mathematics of Computation, 35(152):1383– 1386, October 1980. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Seysen:1987:PFA

[Sey87] Martin Seysen. A probabilistic factorization algorithm with quadratic forms of negative discriminant. Mathematics of Computation, 48(178):757–780, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Smith:1982:NCN

[SF82] David A. Smith and William F. Ford. Numerical comparisons of nonlinear convergence accelerators. Mathematics of Computation, 38(158):481–499, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stinson:1985:STS

[SF85] D. R. Stinson and H. Ferch. 2 000 000 Steiner triple systems of order 19. Mathematics of Computation, 44(170):533– 535, April 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stoyanov:1987:AE

[SF87] Basil J. Stoyanov and Richard A. Farrell. On the asymptotic evaluation of ∫_{0}^{π/2} J_{0}^{2}(γ sin x)dx. Mathematics of Computation, 49(179):275–279, July 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shrager:1980:NCF

[SH80] Richard I. Shrager and Edward Hill, Jr. Nonlinear curve-fitting in the L_{1} and L_{∞} norms. Mathematics of Computation, 34(150):529–541, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanno:1980:VMM

[Sha80] D. F. Shanno. On variable-metric methods for sparse Hessians. Mathematics of Computation, 34(150):499–514, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shampine:1981:TIO

[Sha81] L. F. Shampine. Type-insensitive ODE codes based on implicit A-stable formulas. Mathematics of Computation, 36(154):499–510, April 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shampine:1982:TIO

[Sha82a] L. F. Shampine. Type-insensitive ODE codes based on implicit A(α)-stable formulas. Mathematics of Computation, 39(159):109–123, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1982:CHC

[Sha82b] Daniel Shanks. Corrigendum: H. C.

Williams, “Some primes with interesting digit patterns” [Math. Comp. 32 (1978), no. 144, 1306–1310, MR 58 #484]. Mathematics of Computation, 39(160): 759, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shawyer:1984:BAP

[Sha84] B. L. R. Shawyer. Best approximation of positive power series. Mathematics of Computation, 43(168):529–534, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shampine:1986:SPR

[Sha86] Lawrence F. Shampine. Some practical Runge–Kutta formulas. Mathematics of Computation, 46(173):135–150, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shen:1986:C

[She86] Mok-Kong Shen. On the congruence 2^{n−k} ≡ 1(mod n). Mathematics of Computation, 46(174):715–716, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shi:1984:CPQ

[Shi84] Zhong Ci Shi. On the convergence properties of the quadrilateral elements of Sander and Beckers. Mathematics of Computation, 42(166):493–504, April 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shiu:1986:CST

[Shi86] Peter Shiu. Counting sums of two squares: the Meissel–Lehmer method. Mathematics of Computation, 47(175): 351–360, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigendum [Shi19].

Shi:1987:FMT

[Shi87] Zhong Ci Shi. The F–E–M test for convergence of nonconforming finite elements. Mathematics of Computation, 49 (180):391–405, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shiu:2019:CCS

[Shi19] Peter Shiu. Corrigendum to “Counting sums of two squares: the Meissel– Lehmer method”. Mathematics of Computation, 88(320):2935–2938, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/AMSMathViewer;   https:/
/www.ams.org/journals/mcom/2019-
88-320/S0025-5718-2019-03477-4;

https://www.ams.org/journals/mcom/
2019-88-320/S0025-5718-2019-03477-
4/S0025-5718-2019-03477-4.pdf;
https://www.ams.org/mathscinet/search/
authors.html?mrauthid=160965.

Shrager:1985:RRR

[Shr85] Richard I. Shrager. A rapid robust rootfinder. Mathematics of Computation, 44(169):151–165, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shu:1987:TBT

[Shu87a] Chi-Wang Shu. TVB boundary treatment for numerical solutions of conservation laws. Mathematics of Computation, 49(179):123–134, July 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shu:1987:TUH

[Shu87b] Chi-Wang Shu. TVB uniformly highorder schemes for conservation laws. Mathematics of Computation, 49(179): 105–121, July 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sidi:1980:ACT

[Sid80a] Avram Sidi. Analysis of convergence of the T -transformation for power series. Mathematics of Computation, 35(151): 833–850, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sidi:1980:NQN

[Sid80b] Avram Sidi. Numerical quadrature and nonlinear sequence transformations; unified rules for efficient computation of integrals with algebraic and logarithmic endpoint singularities. Mathematics of Computation, 35(151):851–874, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sidi:1982:NEV

[Sid82a] Avram Sidi. The numerical evaluation of very oscillatory infinite integrals by extrapolation. Mathematics of Computation, 38(158):517–529, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sidi:1982:NQR

[Sid82b] Avram Sidi. Numerical quadrature rules for some infinite range integrals. Mathematics of Computation, 38(157):127– 142, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sidi:1988:UFE

[Sid88] Avram Sidi. A user-friendly extrapolation method for oscillatory infinite integrals. Mathematics of Computation, 51(183):249–266, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Silverman:1987:MPQ

[Sil87] Robert D. Silverman. The multiple polynomial quadratic sieve. Mathematics of Computation, 48(177):329–339, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Silverman:1988:CHE

[Sil88]  Joseph H. Silverman.  Computing

heights on elliptic curves. Mathematics of Computation, 51(183):339–358, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Simon:1984:LAP

[Sim84] Horst D. Simon. The Lánczos algorithm with partial reorthogonalization. Mathematics of Computation, 42(165):115– 142, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Suli:1985:FDA

[SJI85] Endre Süli, Boško Jovanović, and Lav Ivanović. Finite difference approximations of generalized solutions. Mathematics of Computation, 45(172):319– 327, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Skeel:1980:IRI

[Ske80] Robert D. Skeel. Iterative refinement implies numerical stability for Gaussian elimination. Mathematics of Computation, 35(151):817–832, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Skeel:1986:CEF

[Ske86a] R. D. Skeel. Corrigendum: “Equivalent forms of multistep formulas” [Math. Comp. 33 (1979), no. 148, 1229–1250, MR 80j:65027]. Mathematics of Computation, 47(176):769, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Skeel:1986:CVS

[Ske86b] Robert D. Skeel. Construction of variable-stepsize multistep formulas. Mathematics of Computation, 47(176): 503–510, S45–S52, October 1986. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Skeel:1986:SCV

[Ske86c] Robert D. Skeel. Supplement to construction of variable-stepsize multistep formulas. Mathematics of Computation, 47(176):S45–S52, October 1986.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shepherd:1981:CA

[SL81] M. M. Shepherd and J. G. Laframboise. Chebyshev approximation of (1 + 2x) exp(x^{2}) erfc x in 0 ≤ x < ∞. Mathematics of Computation, 36(153):249– 253, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schnorr:1984:MCF

[SL84] C. P. Schnorr and H. W. Lenstra, Jr. A Monte Carlo factoring algorithm with linear storage. Mathematics of Computation, 43(167):289–311, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Salzer:1985:TET

[SL85] Herbert E. Salzer and Norman Levine. Table errata: Tables of sines and cosines to ten decimal places at thousandths of a degree [Pergamon Press, New York, 1962 and MR 26 #361]. Mathematics of Computation, 44(169):281, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sloan:1989:RLQ

[SL89] Ian H. Sloan and James N. Lyness. The representation of lattice quadrature rules as multiple sums. Mathematics of Computation, 52(185):81–94, January 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sloan:1981:QMI

[Slo81] Ian H. Sloan. Quadrature methods for integral equations of the second kind over infinite intervals. Mathematics of Computation, 36(154):511–523, April 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sloan:1983:BCF

[Slo83] D. M. Sloan. Boundary conditions for a fourth order hyperbolic difference scheme. Mathematics of Computation, 41(163):1–11, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sugihara:1982:NHM

[SM82] Masaaki Sugihara and Kazuo Murota. A note on Haselgrove’s method for numerical integration. Mathematics of Computation, 39(160):549–554, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Smith:1983:TEH

[Smi83] John M. Smith. Table errata: Handbook of mathematical functions with formulas, graphs and mathematical tables [Nat. Bur. Standards, Washington, D.C., 1964 and MR 29 #4914]. Mathematics of Computation, 40(162):723– 724, April 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Smith:1989:EMP

[Smi89] David M. Smith. Efficient multipleprecision evaluation of elementary functions. Mathematics of Computation, 52 (185):131–134, January 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Smyth:1981:MTR

[Smy81] C. J. Smyth. On the measure of totally real algebraic integers. II. Mathematics of Computation, 37(155):205–208, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Smyth:1984:MVT

[Smy84] C. J. Smyth. The mean values of totally real algebraic integers. Mathematics of Computation, 42(166):663–681, April 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Streit:1983:NSI

[SN83] Roy L. Streit and Albert H. Nuttall. A note on the semi-infinite programming approach to complex approximation. Mathematics of Computation, 40 (162):599–605, April 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Socolovsky:1986:DSD

[Soc86] E. A. Socolovsky. Difference schemes for degenerate parabolic equations. Mathematics of Computation, 47(176):411– 420, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Solt:1986:TET

[Sol86] G. Solt. Table erratum: Table of integrals, series and products [Academic Press, New York, 1980 and MR 81g:33001] by I. S. Gradshteĭn and I. M. Ryzhik. Mathematics of Computation, 47(176):768, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Speevak:1986:EAO

[Spe86] Ted Speevak. An efficient algorithm for obtaining the volume of a special kind of pyramid and application to convex polyhedra. Mathematics of Computation, 46(174):531–536, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Spekreijse:1987:MSM

[Spe87] Stefan P. Spekreijse. Multigrid solution of monotone second-order discretizations of hyperbolic conservation laws. Mathematics of Computation, 49(179):135–155, July 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Spijker:1985:SRS

[Spi85] M. N. Spijker. Stepsize restrictions for stability of one-step methods in the numerical solution of initial value problems. Mathematics of Computation, 45 (172):377–392, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Spoletini:1984:GPF

[Spo84] Enrico Spoletini. Generation of permutations following Lehmer and Howell. Mathematics of Computation, 43(168): 565–572, October 1984. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sanz-Serna:1984:MNS

[SS84] J. M. Sanz-Serna. Methods for the numerical solution of the nonlinear Schrödinger equation. Mathematics of Computation, 43(167):21–27, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Saad:1985:CGL

[SS85] Youcef Saad and Martin H. Schultz.

Conjugate gradient-like algorithms for solving nonsymmetric linear systems. Mathematics of Computation, 44(170): 417–424, April 1985. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sloan:1986:IEH

[SS86a] I. H. Sloan and A. Spence. Integral equations on the half-line: a modified finitesection approximation. Mathematics of Computation, 47(176):589–595, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stinson:1986:STS

[SS86b] D. R. Stinson and E. Seah. 284 457 Steiner triple systems of order 19 contain a subsystem of order 9. Mathematics of Computation, 46(174):717–729, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Seah:1988:EOF

[SS88] E. Seah and D. R. Stinson. On the enumeration of one-factorizations of complete graphs containing prescribed automorphism groups. Mathematics of Computation, 50(182):607–618, April 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stephan:1989:CVB

[SS89] E. P. Stephan and M. Suri. On the convergence of the p-version of the boundary element Galerkin method. Mathematics of Computation, 52(185):31–48, January 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Smith:1983:PII

[SSO83] William E. Smith, Ian H. Sloan, and Alex H. Opie. Product integration over infinite intervals. I. rules based on the zeros of Hermite polynomials. Mathematics of Computation, 40(162):519– 535, April 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sanz-Serna:1985:GET

[SSP85] J. M. Sanz-Serna and C. Palencia. A general equivalence theorem in the theory of discretization methods. Mathematics of Computation, 45(171):143– 152, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stoer:1987:CSC

[ST87] J. Stoer and R. A. Tapia. On the characterization of q-superlinear convergence of quasi-Newton methods for constrained optimization. Mathematics of Computation, 49(180):581–584, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Starius:1981:AEC

[Sta81] Göran Starius. Asymptotic expansions for a class of elliptic difference schemes. Mathematics of Computation, 37(156): 321–326, October 1981. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Steihaug:1984:SSL

[Ste84a] Trond Steihaug. On the sparse and symmetric least-change secant update. Mathematics of Computation, 42(166): 521–533, April 1984. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stenberg:1984:AMF

[Ste84b] Rolf Stenberg. Analysis of mixed finite element methods for the Stokes problem: a unified approach. Mathematics of Computation, 42(165):9–23, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stewart:1984:ABS

[Ste84c] G. W. Stewart. On the asymptotic behavior of scaled singular value and QR decompositions. Mathematics of Computation, 43(168):483–489, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Steiner:1986:MEP

[Ste86a] Ray P. Steiner. On Mordell’s equation y^{2} − k = x^{3}: a problem of Stolarsky. Mathematics of Computation, 46(174): 703–714, April 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stenger:1986:ENO

[Ste86b] Frank Stenger. Explicit, nearly optimal, linear rational approximation with preassigned poles. Mathematics of Computation, 47(175):225–252, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stewart:1988:CCR

[Ste88] G. W. Stewart. A curiosity concerning the representation of integers in noninteger bases. Mathematics of Computation, 51(184):755–756, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://zbmath.org/?q=
                                              an:0699.10020.

Stokes:1980:SQD

[Sto80] A. N. Stokes. A stable quotientdifference algorithm. Mathematics of Computation, 34(150):515–519, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stothers:1981:SFI

[Sto81] W. W. Stothers. Subgroups of finite index in a free product with amalgamated subgroup. Mathematics of Computation, 36(154):653–662, April 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stroeker:1989:QTE

[Str89] R. J. Stroeker. On quartic Thue equations with trivial solutions. Mathematics of Computation, 52(185):175– 187, January 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stuart:1981:TEC

[Stu81a] S. N. Stuart. Table errata: A course of modern analysis [fourth edition, Cambridge Univ. Press, Cambridge 1927 and 180] by E. T. Whittaker Jbuch 53 and G. N. Watson. Mathematics of Computation, 36(153):319, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stummel:1981:PTE

[Stu81b] F. Stummel. Perturbation theory for evaluation algorithms of arithmetic expressions. Mathematics of Computation, 37(156):435–473, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stynes:1980:FCB

[Sty80] Martin Stynes. On faster convergence of the bisection method for all triangles. Mathematics of Computation, 35(152): 1195–1201, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1981:GFF

[SW81] Daniel Shanks and H. C. Williams. Gunderson’s function in Fermat’s Last Theorem. Mathematics of Computation, 36 (153):291–295, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schatz:1982:QOP

[SW82a] A. H. Schatz and L. B. Wahlbin. On

the quasi-optimality in L_{∞} of the H^{∘}^{1}- projection into finite element spaces. Mathematics of Computation, 38(157): 1–22, January 1982. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shivakumar:1982:AEL

[SW82b] P. N. Shivakumar and R. Wong. Asymptotic expansion of the Lebesgue constants associated with polynomial interpolation. Mathematics of Computation, 39(159):195–200, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schatz:1983:FEM

[SW83] A. H. Schatz and L. B. Wahlbin. On the finite element method for singularly perturbed reaction-diffusion problems in two and one dimensions. Mathematics of Computation, 40(161):47–89, January 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Saranen:1985:ACC

[SW85] J. Saranen and W. L. Wendland. On the asymptotic convergence of collocation methods with spline functions of even degree. Mathematics of Computation, 45(171):91–108, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schneider:1986:SNA

[SW86a] Claus Schneider and Wilhelm Werner. Some new aspects of rational interpolation. Mathematics of Computation, 47(175):285–299, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Segal:1986:WDS

[SW86b] Robert Segal and Robert L. Ward.

Weight distributions of some irreducible cyclic codes. Mathematics of Computation, 46(173):341–354, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schoof:1988:QPR

[SW88a] René Schoof and Lawrence C. Washington. Quintic polynomials and real cyclotomic fields with large class number. Mathematics of Computation, 50(182): 543–556, April 1988. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stephens:1988:CRQ

[SW88b] A. J. Stephens and H. C. Williams.

Computation of real quadratic fields with class number one. Mathematics of Computation, 51(184):809–824, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stephens:1988:SCR

[SW88c] A. J. Stephens and H. C. Williams. Some computational results on a problem concerning powerful numbers. Mathematics of Computation, 50(182):619–632, April 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Swartz:1980:CID

[Swa80] Blair Swartz. Compact, implicit difference schemes for a differential equation’s side conditions. Mathematics of Computation, 35(151):733–746, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Swarztrauber:1986:SF

[Swa86] Paul N. Swarztrauber. Symmetric FFTs. Mathematics of Computation, 47(175): 323–346, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Swartz:1989:SOS

[Swa89a] Blair Swartz. The second-order sharpening of blurred smooth borders. Mathematics of Computation, 52(186):675– 714, S35–S36, April 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Swartz:1989:SSO

[Swa89b] Blair Swartz. Supplement to the secondorder sharpening of blurred smooth borders. Mathematics of Computation, 52 (186):S35–S36, April 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Seah:1983:CLC

[SWW83] Eric Seah, Lawrence C. Washington, and Hugh C. Williams. The calculation of a large cubic class number with an application to real cyclotomic fields. Mathematics of Computation, 41(163):303– 305, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Szepessy:1989:CSC

[Sze89] Anders Szepessy. Convergence of a shock-capturing streamline diffusion finite element method for a scalar conservation law in two space dimensions. Mathematics of Computation, 53(188): 527–545, October 1989. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tadmor:1983:UII

[Tad83] Eitan Tadmor. The unconditional instability of inflow-dependent boundary conditions in difference approximations to hyperbolic systems. Mathematics of Computation, 41(164):309–319, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tadmor:1984:LTB

[Tad84a] Eitan Tadmor. The large-time behavior of the scalar, genuinely nonlinear Lax–Friedrichs scheme. Mathematics of Computation, 43(168):353–368, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tadmor:1984:NVE

[Tad84b] Eitan Tadmor. Numerical viscosity and the entropy condition for conservative difference schemes. Mathematics of Computation, 43(168):369–381, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tadmor:1987:NVE

[Tad87] Eitan Tadmor. The numerical viscosity of entropy stable schemes for systems of conservation laws. I. Mathematics of Computation, 49(179):91–103, July 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tang:1988:FAL

[Tan88] Ping Tak Peter Tang. A fast algorithm for linear complex Chebyshev approximations. Mathematics of Computation, 51(184):721–739, October 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tapia:1988:SUU

[Tap88] Richard Tapia. On secant updates for use in general constrained optimization.

Mathematics of Computation, 51(183): 181–202, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

teRiele:1981:HNT

[te 81] Herman J. J. te Riele. Hyperperfect numbers with three different prime factors. Mathematics of Computation, 36 (153):297–298, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

teRiele:1984:GNA

[te 84] Herman J. J. te Riele. On generating new amicable pairs from given amicable pairs. Mathematics of Computation, 42(165):219–223, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

teRiele:1986:CAA

[te 86a] H. J. J. te Riele. Computation of all the amicable pairs below 10^{10}. Mathematics of Computation, 47(175):361–368, S9– S40, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

teRiele:1986:SCA

[te 86b] H. J. J. te Riele. Supplement to computation of all the amicable pairs below 10^{1}0. Mathematics of Computation, 47(175):S9–S40, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

teRiele:1986:CZRa

[te 86c] Herman J. J. te Riele. Corrigenda: “On the zeros of the Riemann zeta function in the critical strip. II” [Math. Comp. 39 (1982), no. 160, 681–688, MR 83m:10067] by R. P. Brent, J. van de Lune, te Riele and D. T. Winter. Mathematics of Computation, 46(174):

771, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

teRiele:1986:CZRb

[te 86d] Herman J. J. te Riele. Corrigenda: “On the zeros of the Riemann zeta function in the critical strip. III” [Math. Comp. 41 (1983), no. 164, 759–767, MR 85e:11062] by J. van de Lune and te Riele. Mathematics of Computation, 46 (174):771, April 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

teRiele:1987:SD

[te 87] Herman J. J. te Riele. On the sign of the difference π(x) − li(x). Mathematics of Computation, 48(177):323–328, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Templer:1980:P

[Tem80] Mark Templer. On the primality of k!+1 and 2 ∗ 3 ∗ 5 ∗ ⋯ ∗ p + 1. Mathematics of Computation, 34(149):303–304, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Temme:1981:TEF

[Tem81] N. M. Temme. Table errata: Formulas and theorems for the special functions of mathematical physics [third edition, Springer, New York, 1966 and F. Oberhettinger MR 38 #1291] by W. Magnus and R. P. Soni. Mathematics of Computation, 36(153):315–316, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Temme:1986:DIC

[Tem86] N. M. Temme. A double integral containing the modified Bessel function: asymptotics and computation. Mathematics of Computation, 47(176):683– 691, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tezuka:1988:OGP

[Tez88] Shu Tezuka. On optimal GFSR pseudorandom number generators. Mathematics of Computation, 50(182):531– 533, April 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Trefethen:1986:WPO

[TH86] Lloyd N. Trefethen and Laurence Halpern. Well-posedness of one-way wave equations and absorbing boundary conditions. Mathematics of Computation, 47(176):421–435, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Thacher:1981:CAH

[Tha81]  Henry C. Thacher.  Book review:

Computational Analysis with the HP-25 Pocket Calculator, by Peter Henrici. Mathematics of Computation, 37(156): 595–596, October 1981. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842

                                              (electronic). URL https://www.jstor.
                                              org/stable/2007449.

Thomee:1980:NNE

[Tho80] Vidar Thomée. Negative norm estimates and superconvergence in Galerkin methods for parabolic problems. Mathematics of Computation, 34(149):93–113, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Thomas:1981:GMS

[Tho81] K. S. Thomas. Galerkin methods for singular integral equations. Mathematics of Computation, 36(153):193–205, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ting:1981:CIO

[TL81] Bing Yuan Ting and Yudell L. Luke.

Computation of integrals with oscillatory and singular integrands. Mathematics of Computation, 37(155):169– 183, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Traub:1985:OIF

[TL85] J. F. Traub and D. Lee. Optimal integration for functions of bounded variation. Mathematics of Computation, 45(172): 505–512, October 1985. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Toint:1981:SQN

[Toi81] Ph. L. Toint. A sparse quasi-Newton update derived variationally with a nondiagonally weighted Frobenius norm. Mathematics of Computation, 37(156): 425–433, October 1981. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Toint:1986:NSL

[Toi86] Ph. L. Toint. Numerical solution of large sets of algebraic nonlinear equations. Mathematics of Computation, 46(173): 175–189, January 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Trefethen:1985:SFD

[Tre85] Lloyd N. Trefethen. Stability of finitedifference models containing two boundaries or interfaces. Mathematics of Computation, 45(172):279–300, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Troesch:1980:SMA

[Tro80] B. A. Troesch. The shooting method applied to a cyclic inequality. Mathematics of Computation, 34(149):175– 184, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Troesch:1985:SCI

[Tro85] B. A. Troesch. On Shapiro’s cyclic inequality for N = 13. Mathematics of Computation, 45(171):199–207, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Troesch:1989:VSC

[Tro89]  B. A. Troesch.  The validity of

Shapiro’s cyclic inequality. Mathematics of Computation, 53(188):657–664, October 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tsuchiya:1986:TMA

[Tsu86] Takuya Tsuchiya. On two methods for approximating minimal surfaces in parametric form. Mathematics of Computation, 46(174):517–529, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tsuchiya:1987:DSP

[Tsu87] Takuya Tsuchiya. Discrete solution of the Plateau problem and its convergence. Mathematics of Computation, 49(179):157–165, July 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tretter:1980:FCC

[TW80] Marietta J. Tretter and G. W. Walster. Further comments on the computation of modified Bessel function ratios. Mathematics of Computation, 35(151):937– 939, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Thron:1983:TEB

[TW83] W. J. Thron and Haakon Waadeland.

Truncation error bounds for limit periodic continued fractions. Mathematics of Computation, 40(162):589–597, April 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tennenhouse:1986:NCN

[TW86] M. Tennenhouse and H. C. Williams. A note on class-number one in certain real quadratic and pure cubic fields. Mathematics of Computation, 46(173):333– 336, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tanner:1987:NCB

[TW87] Jonathan W. Tanner and Samuel S.

Wagstaff, Jr. New congruences for the Bernoulli numbers. Mathematics of Computation, 48(177):341–350, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tanner:1989:NBF

[TW89] Jonathan W. Tanner and Samuel S.

Wagstaff, Jr. New bound for the first case of Fermat’s last theorem. Mathematics of Computation, 53(188):743– 750, October 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tschope:1987:CNT

[TZ87] Heinz M. Tschöpe and Horst G. Zimmer. Computation of the Néron–Tate height on elliptic curves. Mathematics of Computation, 48(177):351–370, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Thomee:1989:EES

[TZ89] Vidar Thomée and Nai Ying Zhang. Error estimates for semidiscrete finite element methods for parabolic integrodifferential equations. Mathematics of Computation, 53(187):121–139, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tzanakis:1986:RTW

[Tza86] Nicholas Tzanakis. A remark on a theorem of W. E. H. Berwick. Mathematics of Computation, 46(174):623–625, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Underhill:1986:SAP

[Und86] C. Underhill. Some asymptotic properties of Padé approximants to e^{−x}. Mathematics of Computation, 47(175):253– 263, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vanderLinden:1982:CNC

[van82a] F. J. van der Linden. Class number computations of real Abelian number fields. Mathematics of Computation, 39(160): 693–707, October 1982. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

VanderVorst:1982:GLS

[Van82b] H. A. Van der Vorst. A generalized Lánczos scheme. Mathematics of Computation, 39(160):559–561, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vanVeldhuizen:1988:CRI

[van88] M. van Veldhuizen. Convergence results for invariant curve algorithms. Mathematics of Computation, 51(184):677– 697, October 1988. CODEN MCMPAF.

ISSN 0025-5718 (print), 1088-6842 (electronic).

Vaughan:1983:DQE

[Vau83] Theresa P. Vaughan. The discriminant of a quadratic extension of an algebraic field. Mathematics of Computation, 40 (162):685–707, April 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Vaughan:1984:CDQ

[Vau84]  Theresa P. Vaughan.  Corrigenda:

“The discriminant of a quadratic extension of an algebraic field” [Math. Comp. 40 (1983), no. 162, 685–707, MR 84e:12006]. Mathematics of Computation, 43(168):621, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Vaughan:1985:CUC

[Vau85a] Theresa P. Vaughan. The construction of unramified cyclic quartic extensions of Q(\sqrt{m}). Mathematics of Computation, 45(171):233–242, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Vaughan:1985:CDA

[Vau85b] Theresa P. Vaughan. On computing the discriminant of an algebraic number field. Mathematics of Computation, 45 (172):569–584, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Vogelius:1981:DRMa

[VB81a] M. Vogelius and I. Babuška. On a dimensional reduction method. I. the optimal selection of basis functions. Mathematics of Computation, 37(155):31–46, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Vogelius:1981:DRMb

[VB81b] M. Vogelius and I. Babuška. On a dimensional reduction method. II. some approximation-theoretic results. Mathematics of Computation, 37(155):47– 68, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Vogelius:1981:DRMc

[VB81c] M. Vogelius and I. Babuška. On a dimensional reduction method. III. A posteriori error estimation and an adaptive approach. Mathematics of Computation, 37(156):361–384, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Venturino:1986:SSI

[Ven86] Ezio Venturino. On solving singular integral equations via a hyperbolic tangent quadrature rule. Mathematics of Computation, 47(175):159–167, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Veron:1982:SRC

[Vér82] Laurent Véron. Some remarks on the convergence of approximate solutions of nonlinear evolution equations in Hilbert spaces. Mathematics of Computation, 39 (160):325–337, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Vila:1988:HOS

[Vil88] J. P. Vila. High-order schemes and entropy condition for nonlinear hyperbolic systems of conservation laws. Mathematics of Computation, 50(181):53–73, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vanHaeringen:1982:TET

[vK82] H. van Haeringen and L. P. Kok. Table errata: Table of integrals, series, and products [corrected and 1980 enlarged edition, Academic Press, New York and MR 81g:33001] by I. S. Gradshteyn [I. S. Gradshteĭn] and I. M. Ryzhik. Mathematics of Computation, 39(160):747– 757, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vanHaeringen:1983:TEF

[vK83a] H. van Haeringen and L. P. Kok. Table errata: Formulas and theorems for the special functions of mathematical physics [Springer, New York, 1966 and F. Oberhettinger MR 38 #1291] by W. Magnus and R. P. Soni. Mathematics of Computation, 41(164):775–778, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vanHaeringen:1983:TEHd

[vK83b] H. van Haeringen and L. P. Kok. Table errata: Handbook of mathematical functions, with formulas, graphs, and mathematical tables [Dover, New York, 1966 and MR 34 #8606]. Mathematics of Computation, 41(164):780, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vanHaeringen:1983:TEHa

[vK83c] H. van Haeringen and L. P. Kok. Table errata: Higher transcendental functions, Vol. I [McGraw-Hill, New York, 1955, F. Oberhettinger, MR 16, 586] by A. Erdélyi, W. Magnus and F. G. Tricomi. Mathematics of Computation, 41(164): 778–779, October 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vanHaeringen:1983:TEHb

[vK83d] H. van Haeringen and L. P. Kok. Table errata: Higher transcendental functions, Vol. II [McGraw-Hill, New York, 1955, F. Oberhettinger, MR 16, 586] by A. Erdélyi, W. Magnus and F. G. Tricomi. Mathematics of Computation, 41(164): 778, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vanHaeringen:1983:TEHc

[vK83e] H. van Haeringen and L. P. Kok. Table errata: Higher transcendental functions, Vol. III [McGraw-Hill, New York, 1955, F. Oberhettinger, MR 16, 586] by A. Erdélyi, W. Magnus and F. G. Tricomi. Mathematics of Computation, 41(164): 778, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vanHaeringen:1983:TETa

[vK83f] H. van Haeringen and L. P. Kok. Table errata: Tables of integral transforms, Vol. I [McGraw-Hill, New York 1954, F. Oberhettinger MR 15, 868] by A. Erdélyi, W. Magnus and F. G. Tricomi. Mathematics of Computation, 41(164): 778–779, October 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vanHaeringen:1983:TETb

[vK83g] H. van Haeringen and L. P. Kok. Table errata: Tables of integral transforms, Vol. II [McGraw-Hill, New York, 1954, F. Oberhettinger MR 16, 468] by A. Erdélyi, W. Magnus and F. G. Tricomi. Mathematics of Computation, 41(164): 779–780, October 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vonzurGathen:1985:FMP

[vK85] J. von zur Gathen and E. Kaltofen. Factorization of multivariate polynomials over finite fields. Mathematics of Computation, 45(171):251–261, July 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Haeringen:1986:TET

[vK86] H. van Haeringen and L. P. Kok. Table erratum: A table of series and products [Prentice-Hall, Englewood Cliffs, N.J., 1975] by E. R. Hansen. Mathematics of Computation, 47(176):767, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Vigneron:1980:GQI

[VL80] J. P. Vigneron and Ph. Lambin. Gaussian quadrature of integrands involving the error function. Mathematics of Computation, 35(152):1299–1307, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vonzurGathen:1984:HNM

[von84] Joachim von zur Gathen. Hensel and Newton methods in valuation rings. Mathematics of Computation, 42(166): 637–661, April 1984. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Vrscay:1986:JSM

[Vrs86] Edward R. Vrscay. Julia sets and Mandelbrot-like sets associated with higher order Schröder rational iteration functions: a computer assisted study. Mathematics of Computation, 46(173): 151–169, January 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vandeLune:1983:ZRZ

[vt83] J. van de Lune and H. J. J. te Riele. On the zeros of the Riemann zeta function in the critical strip. III. Mathematics of Computation, 41(164):759–767, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vanderHouwen:1985:LMM

[vt85a] P. J. van der Houwen and H. J. J. te Riele. Linear multistep methods for Volterra integral and integro-differential equations. Mathematics of Computation, 45(172):439–461, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vanderHouwen:1985:SLM

[vt85b] P. J. van der Houwen and H. J. J. te Riele. Supplement to: “Linear multistep methods for Volterra integral and integro-differential equations”. Mathematics of Computation, 45(172):S21– S28, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vandeLune:1986:ZRZ

[vtW86] J. van de Lune, H. J. J. te Riele, and D. T. Winter. On the zeros of the Riemann zeta function in the critical strip. IV. Mathematics of Computation, 46(174):667–681, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

JW:1980:C

[W.80] J. W. Corrigendum. Mathematics of Computation, 34(149):332, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). Original journal has only ‘J. W.’ as author.

Wagstaff:1980:DCS

[Wag80] Samuel S. Wagstaff, Jr. p-divisibility of certain sets of Bernoulli numbers. Mathematics of Computation, 34(150):647– 649, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wagstaff:1983:DMN

[Wag83] Samuel S. Wagstaff, Jr. Divisors of Mersenne numbers. Mathematics of Computation, 40(161):385–397, January 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wahlbin:1984:SCL

[Wah84] Lars B. Wahlbin. On the sharpness of

certain local estimates for H^{∘}^{1} projections into finite element spaces: Influence of a reentrant corner. Mathematics of Computation, 42(165):1–8, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Washington:1986:SRC

[Was86] Lawrence C. Washington. Some remarks on Cohen–Lenstra heuristics. Mathematics of Computation, 47(176):741– 747, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Washington:1987:CNS

[Was87] Lawrence C. Washington. Class numbers of the simplest cubic fields. Mathematics of Computation, 48(177):371– 384, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wills:1982:RCA

[WBR82] C. A. Wills, J. M. Blair, and P. L.

Ragde. Rational Chebyshev approximations for the Bessel functions J_{0}(x), J_{1}(x), Y_{0}(x), Y_{1}(x). Mathematics of Computation, 39(160):617–623, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1980:CRP

[WCS80] H. C. Williams, G. Cormack, and E. Seah. Calculation of the regulator of a pure cubic field. Mathematics of Computation, 34(150):567–611, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1984:ANI

[WD84] H. C. Williams and G. W. Dueck. An analogue of the nearest integer continued fraction for certain cubic irrationalities. Mathematics of Computation, 42(166): 683–705, April 1984. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1986:P

[WD86] H. C. Williams and Harvey Dubner. The primality of R1031. Mathematics of Computation, 47(176):703–711, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1983:RME

[WDS83] H. C. Williams, G. W. Dueck, and B. K. Schmid. A rapid method of evaluating the regulator and class number of a pure cubic field. Mathematics of Computation, 41(163):235–286, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Weintraub:1981:LPG

[Wei81] Sol Weintraub. A large prime gap. Mathematics of Computation, 36(153): 279, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Weinmuller:1984:DMS

[Wei84a] Ewa Weinmüller. A difference method for a singular boundary value problem of second order. Mathematics of Computation, 42(166):441–464, April 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Weiss:1984:ABT

[Wei84b] Richard Weiss. An analysis of the box and trapezoidal schemes for linear singularly perturbed boundary value problems. Mathematics of Computation, 42 (165):41–67, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Weinmuller:1986:NSS

[Wei86] Ewa Weinmüller. On the numerical solution of singular boundary value problems of second order by a difference method. Mathematics of Computation, 46(173): 93–117, January 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Werschulz:1980:CCO

[Wer80] Arthur G. Werschulz. Computational complexity of one-step methods for systems of differential equations. Mathematics of Computation, 34(149):155– 174, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Werschulz:1982:OEP

[Wer82] Arthur G. Werschulz. Optimal error properties of finite element methods for second order elliptic Dirichlet problems. Mathematics of Computation, 38(158): 401–413, April 1982. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Werner:1984:PIL

[Wer84a] Wilhelm Werner. Polynomial interpolation: Lagrange versus Newton. Mathematics of Computation, 43(167):205– 217, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Werschulz:1984:DIR

[Wer84b] Arthur G. Werschulz. Does increased regularity lower complexity? Mathematics of Computation, 42(165):69–93, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Werschulz:1986:CIE

[Wer86a] Arthur G. Werschulz. Complexity of indefinite elliptic problems. Mathematics of Computation, 46(174):457–477, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Werschulz:1986:WCR

[Wer86b] Arthur G. Werschulz. What is the complexity of related elliptic, parabolic, and hyperbolic problems? Mathematics of Computation, 47(176):461–472, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wrigge:1982:GMA

[WF82] Staffan Wrigge and Arne Fransén. A general method of approximation. part I. Mathematics of Computation, 38(158): 567–588, April 1982. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wilkins:1983:MQS

[WH83] J. Ernest Wilkins, Jr. and Theodore R. Hatcher. The maximum of a quasismooth function. Mathematics of Computation, 41(164):573–589, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

White:1983:MFD

[Whi83] R. E. White. A modified finite difference scheme for the Stefan problem. Mathematics of Computation, 41(164):337– 347, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wilker:1980:EAS

[Wil80a] Peter Wilker. An efficient algorithmic solution of the Diophantine equation u^{2} +5v^{2} = m. Mathematics of Computation, 35(152):1347–1352, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Willett:1980:AFF

[Wil80b] Michael Willett. Arithmetic in a finite field. Mathematics of Computation, 35(152):1353–1359, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1980:ISC

[Wil80c] H. C. Williams. Improving the speed of calculating the regulator of certain pure cubic fields. Mathematics of Computation, 35(152):1423–1434, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1981:NIL

[Wil81a] H. C. Williams. A numerical investigation into the length of the period of the continued fraction expansion of \sqrt{D}. Mathematics of Computation, 36(154): 593–601, April 1981. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1981:SRC

[Wil81b] H. C. Williams. Some results concerning Voronoĭ’s continued fraction over Q(\sqrt[3]{D}). Mathematics of Computation, 36(154):631–652, April 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1982:MF

[Wil82a] H. C. Williams. A p + 1 method of factoring. Mathematics of Computation, 39(159):225–234, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1982:DPF

[Wil82b] H. C. Williams. Determination of principal factors in Q(\sqrt{D}) and Q(\sqrt[3]{D}). Mathematics of Computation, 38(157): 261–274, January 1982. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1987:EPT

[Wil87] H. C. Williams. Effective primality tests for some integers of the forms A5^{n} − 1 and A7^{n} − 1. Mathematics of Computation, 48(177):385–403, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1988:CCC

[Wil88] H. C. Williams. Corrigenda: “Computation of the class number and class group of a complex cubic field” [Math. Comp. 45 (1985), no. 171, 223–231, MR 86m:11078] by G. W. Dueck and Williams. Mathematics of Computation, 50(182):655–657, April 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Winther:1980:CFE

[Win80a] Ragnar Winther. A conservative finite element method for the Korteweg–de Vries equation. Mathematics of Computation, 34(149):23–43, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Winther:1980:IVM

[Win80b] Ragnar Winther. Initial value methods for parabolic control problems. Mathematics of Computation, 34(149):115– 125, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Winther:1981:SFE

[Win81] Ragnar Winther. A stable finite element method for initial-boundary value problems for first-order hyperbolic systems. Mathematics of Computation, 36(153): 65–86, January 1981. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wong:1981:MAE

[WM81] R. Wong and J. P. McClure. On a method of asymptotic evaluation of multiple integrals. Mathematics of Computation, 37(156):509–521, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wolkenfelt:1983:MMM

[Wol83] P. H. M. Wolkenfelt. Modified multilag methods for Volterra functional equations. Mathematics of Computation, 40 (161):301–316, January 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wong:1988:AE

[Won88] R. Wong. Asymptotic expansion of ∫_{0}^{π/2} J_{ν}^{2}(λ cos θ)dθ. Mathematics of Computation, 50(181):229–234, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wrigge:1981:CTS

[Wri81a] Staffan Wrigge. Calculation of the Taylor series expansion coefficients of the Jacobian elliptic function sn(x, k). Mathematics of Computation, 36(154):555– 564, April 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wrigge:1981:NTS

[Wri81b] Staffan Wrigge. A note on the Taylor series expansion coefficients of the Jacobian elliptic function sn(x, k). Mathematics of Computation, 37(156):495– 497, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wrigge:1984:NMG

[Wri84] Staffan Wrigge. A note on the moment generating function for the reciprocal gamma distribution. Mathematics of Computation, 42(166):617–621, April 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wright:1985:FAS

[Wri85] Alden H. Wright. Finding all solutions to a system of polynomial equations. Mathematics of Computation, 44(169):125– 133, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wroblewski:1984:NSI

[Wró84] J. Wróblewski. A nonaveraging set of integers with a large sum of reciprocals. Mathematics of Computation, 43(167): 261–262, July 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Warsi:1982:NMG

[WT82] Z. U. A. Warsi and J. F. Thompson.

A noniterative method for the generation of orthogonal coordinates in doublyconnected regions. Mathematics of Computation, 38(158):501–516, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wunderlich:1983:PAS

[Wun83] M. C. Wunderlich. A performance analysis of a simple prime-testing algorithm. Mathematics of Computation, 40(162): 709–714, April 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wunderlich:1985:ICF

[Wun85] Marvin C. Wunderlich. Implementing the continued fraction factoring algorithm on parallel machines. Mathematics of Computation, 44(169):251– 260, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1987:PGR

[WW87] H. C. Williams and M. C. Wunderlich. On the parallel generation of the residues for the continued fraction factoring algorithm. Mathematics of Computation, 48(177):405–423, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Winker:1981:SAI

[WWL81] S. K. Winker, L. Wos, and E. L.

Lusk. Semigroups, antiautomorphisms, and involutions: a computer solution to an open problem. I. Mathematics of Computation, 37(156):533–545, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Weiser:1988:NPL

[WZ88] Alan Weiser and Sergio E. Zarantonello. A note on piecewise linear and multilinear table interpolation in many dimensions. Mathematics of Computation, 50 (181):189–196, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Xu:1989:WFC

[Xu89] Yuan Xu. Weight functions for Chebyshev quadrature. Mathematics of Computation, 53(187):297–302, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Yamaki:1986:CFS

[Yam86a] Hiroyoshi Yamaki. A conjecture of Frobenius and the sporadic simple groups. II. Mathematics of Computation, 46(174):609–611, S43–S46, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Yamaki:1986:SCF

[Yam86b] Hiroyoshi Yamaki. Supplement to A conjecture of Frobenius and the sporadic simple groups. II. Mathematics of Computation, 46(174):S43–S46, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Yan:1987:PCC

[Yan87] Zheng Yan. Piecewise cubic curve fitting algorithm. Mathematics of Computation, 49(179):203–213, July 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Young:1988:TFN

[YB88] Jeff Young and Duncan A. Buell. The twentieth Fermat number is composite. Mathematics of Computation, 50(181):

261–263, January 1988. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Young:1989:FOP

[YP89] Jeff Young and Aaron Potler. First occurrence prime gaps. Mathematics of Computation, 52(185):221–224, January 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ypma:1983:LCD

[Ypm83] T. J. Ypma. Local convergence of difference Newton-like methods. Mathematics of Computation, 41(164):527– 536, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Yserentant:1986:CML

[Yse86] Harry Yserentant. The convergence of multi-level methods for solving finiteelement equations in the presence of singularities. Mathematics of Computation, 47(176):399–409, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zagier:1982:NMN

[Zag82] Don Zagier. On the number of Markoff numbers below a given bound. Mathematics of Computation, 39(160):709– 723, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zagier:1987:LIP

[Zag87] Don Zagier. Large integral points on elliptic curves. Mathematics of Computation, 48(177):425–436, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zagier:1988:ALI

[Zag88] Don Zagier. Addendum to: “Large integral points on elliptic curves” [Math. Comp. 48 (1987), no. 177, 425–436, MR 87k:11062]. Mathematics of Computation, 51(183):375, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zajta:1983:SDE

[Zaj83] Aurel J. Zajta. Solutions of the Diophantine equation A^{4} + B^{4} = C^{4} + D^{4}. Mathematics of Computation, 41(164): 635–659, October 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zennaro:1986:NCE

[Zen86] M. Zennaro. Natural continuous extensions of Runge–Kutta methods. Mathematics of Computation, 46(173):119– 133, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ziv:1982:RDE

[Ziv82] Abraham Ziv. Relative distance—an error measure in round-off error analysis. Mathematics of Computation, 39(160): 563–569, October 1982. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zlamal:1983:LSN

[Zlá83] Miloš Zlámal. A linear scheme for the numerical solution of nonlinear quasistationary magnetic fields. Mathematics of Computation, 41(164):425–440, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zlamal:1986:FES

[Zlá86] Miloš Zlámal. Finite element solution of the fundamental equations of semiconductor devices. I. Mathematics of Computation, 46(173):27–43, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zulehner:1988:SHM

[Zul88] Walter Zulehner. A simple homotopy method for determining all isolated solutions to polynomial systems. Mathematics of Computation, 50(181):167– 177, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).